### v2

# Learning large softmax mixtures with warm start EM

Abstract
Softmax mixture models (SMMs) are discrete $K$-mixtures introduced to model the probability of choosing an attribute $x_{j}\in\mathbb{R}^{L}$ from $p$ possible candidates, in heterogeneous populations. They have been known, for several decades, as mixed multinomial logits in the econometrics literature, and are gaining traction in the LLM literature, where single softmax models are routinely used in the final layer of a neural network. The theoretical understanding of this mixture model lags behind its growing popularity, and we aim to narrow this gap in this work.
This paper provides a comprehensive analysis of the Expectation-Maximization (EM) algorithm for SMMs, in high dimensions. It complements and extends existing results currently restricted to Gaussian Mixture Models (GMMs). Its population-level theoretical analysis offers key insights into the model that go beyond the typical parameter estimation EM usage. It forms the basis for proving (i) local identifiability, in SSMs with generic features and, further, via a stochastic argument, (ii) full identifiability in SSMs with random features, when $p$ is large enough. To the best of our knowledge, these are the first results in this direction for SSMs with $L>1$.
The population-level EM analysis includes the characterization of the initialization radius for algorithmic convergence. This also guides the construction of possible warm starts of the sample level EM algorithm. Under any warm start initialization, the EM algorithm is shown to recover the mixture atoms of the SSM at the parametric rate, up to logarithmic factors.
We provide two main directions for warm start construction, both based on a new method for estimating the moments of the mixing measure underlying an SSM with random design. First, we construct a method of moments (MoM) preliminary estimator of the mixture parameters, and provide its first theoretical analysis in SSMs. While MoM can enjoy parametric rates of convergence, and thus can serve as a warm-start, the estimator’s quality degrades exponentially in $K$, a fact already demonstrated for GMMs, even when $L=1$. Our recommendation, especially when $K$ is not small, is to follow common practice and run the EM algorithm several times with random initializations. We again make use of the novel estimation method tailored to latent moments in SSMs to further estimate the $K$-dimensional subspace of $\mathbb{R}^{L}$ spanned by the atoms of the mixture. Sampling from this subspace reduces substantially the number of required draws, from $\exp(L)$ to $\exp(K)$, and is also shown to have empirical success.
Keywords: Softmax, mixture models, method of moments, EM algorithm, parameter estimation, mixed multinomial logits, latent class models.
1 Introduction
1.1 The softmax mixture model
“Softmax mixtures” define a parametric discrete mixture model $\pi\in\Delta^{p}$, the probability simplex in $\mathbb{R}^{p}$, supported on a known set of vectors $x_{1},\dots,x_{p}\in\mathbb{R}^{L}$.
For a given, known and finite $K$, we let ${\bm{\theta}}_{k}\in\mathbb{R}^{L}$, $k\in[K]:=\{1,\ldots,K\}$ be distinct vectors in $\mathbb{R}^{L}$. Each mixture component $A({\bm{\theta}}_{k}):=A(\cdot;{\bm{\theta}}_{k})$ of a softmax mixture is a probability vector in $\Delta^{p}$, supported on $x_{1},\ldots,x_{p}$, parametrized via the softmax function $\mathsf{softmax}:\mathbb{R}^{p}\to\Delta^{p}$,
$A(x_{j};{\bm{\theta}}_{k})=\bigl{[}\mathsf{softmax}(x_{1}^{\top}{\bm{\theta}}_{k},\ldots,x_{p}^{\top}{\bm{\theta}}_{k})\bigr{]}_{j}=\frac{{\exp({x_{j}^{\top}{{\bm{\theta}}_{k}}})}}{\sum_{i=1}^{p}\exp({x_{i}^{\top}{{\bm{\theta}}_{k}}})},$ (1)
for each $j\in[p]$. If we let ${\bm{\alpha}}:=(\alpha_{1},\ldots,\alpha_{K})^{\top}\in\Delta^{K}$ denote the vector of mixing weights, and write ${\bm{\omega}}:=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$, the softmax mixture model is given by
$\pi(y;{\bm{\omega}}):=\sum_{k=1}^{K}\alpha_{k}A(y;{\bm{\theta}}_{k}),\qquad\mbox{for}\ y\in\{x_{1},\ldots,x_{p}\}.$ (2)
Throughout this paper, our focus is on estimating the parameters ${\bm{\omega}}^{*}=({\bm{\alpha}}^{*},{\bm{\theta}}_{1}^{*},\dots,{\bm{\theta}}_{K}^{*})$ from a sample $Y_{1},\ldots,Y_{N}$ from $\pi^{*}(y):=\pi(y;{\bm{\omega}}^{*})$.
When $K=1$, the softmax mixture model reduces to what is known in the classical statistical literature as the conditional logit model (McFadden, 1974). Its usage and properties, when both $p$ and $L$ are fixed, have been thoroughly studied, see McFadden (1974) and the literature review in (Agresti, 1990, Chapter 9). Much less is known about the case $K>1$, which has received very little attention in the mathematical statistics literature. This paper bridges this gap, and also complements and extends the existing literature on parameter estimation via the expectation-maximization (EM) algorithm beyond the well-studied case of Gaussian Mixture Models with $K$ components ($K$-GMM). We highlight the main contributions of this paper below.
  1. 1.
    We develop a hybrid EM algorithm for parameter estimation under softmax mixture models and prove that it converges to the true model parameters at a near-parametric rate after $\mathcal{O}(\log N)$ iterations. Each iteration has computational complexity $\mathcal{O}(pL)$. Our analysis gives conditions on the choice of the algorithm’s initialization, and on the separation between mixture components, under which EM converges. Notably, and improving upon the sharpest known result, albeit developed only for $K$-GMMs, we require that the atom separation depend only logarithmically on the number of components and the smallest mixing weights. As a consequence of the convergence of the population-level EM algorithm, we prove that softmax mixtures are locally identifiable. Section 1.2.1 gives more details and the background for these results, which are formally stated and proved in Sections 2.1, 2.2 and 2.3.
  2. 2.
    We develop a new Method of Moments (MoM), specifically tailored to softmax mixtures, for estimating the latent moments of the mixing measure $\rho:=\sum_{k=1}^{K}\alpha_{k}^{*}\delta_{{\bm{\theta}}_{k}^{*}}$, where $\delta$ denotes the Dirac measure on $\mathbb{R}^{L}$. Under the assumption that the features $x_{1},\ldots,x_{p}$ are independent realizations from a given distribution, we make use of this construction in three related, but different, ways. The background is given in Section 1.2.2.
    • We use a system of equations involving appropriate latent moment approximations, at the population level, to find initial atoms and weights close to the true parameters. Those are then used to initialize a population level EM algorithm to prove that softmax mixtures are globally identifiable, for $p$ large enough. This is the content of Sections 3.1 and 3.2.
    • We develop the sample level analogue of this result. We derive MoM parameter estimates in softmax mixtures, and offer the first rate analysis under this model. The analysis complements that for Gaussian Mixture Models ($K$-GMM), and is valid for any $L\geq 1$. We show, in Section 3.3 that MoM estimators can serve as a warm start for the EM algorithm, but their performance deteriorates fast as $K$ increases.
    • We recommend random initialization when $K$ is not small. For this, in Section 3.4 we develop an estimator for the subspace spanned by ${\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}$, tailored to softmax mixtures, and based only on second-order latent moment estimates. We show how to use this subspace estimator to reduce the number of random draws needed to initialize the EM algorithm.
In addition to bridging the existing theoretical and algorithmic gap in softmax mixture estimation via the EM algorithm, our focus on parameter estimation is also motivated by the model’s applications. The model is widely used in the econometrics literature, and could also play an important role in understanding aspects of an LLM output. We give below instances of such applications.
Basic discrete choice models. Softmax mixtures were introduced in the econometrics literature by Boyd and Mellman (1980) and Cardell and Dunbar (1980) under the name “mixed multinomial logits” to model the preference of a heterogeneous set of consumers for a set of mutually exclusive goods. In this application, each vector $x_{j}\in\mathbb{R}^{L}$ reflects the set of attributes of each of the $p$ different goods, while the vector ${{\bm{\theta}}}$ reflects a customer’s preferences for each attribute. The model posits that customers act via random utility maximization: the customer chooses good $j^{\star}=\operatorname*{argmax}_{j\in[p]}x_{j}^{\top}{{\bm{\theta}}}+\epsilon_{j}$. Here $\epsilon_{1},\ldots,\epsilon_{p}$ are independent stochastic terms that reflect idiosyncratic variations in the consumer’s taste. When $\epsilon_{1},\ldots,\epsilon_{p}$ are chosen to have a Gumbel distribution, then (see, e.g., Yellott Jr, 1977)
$\mathbb{P}\left\{j^{\star}=j\right\}=\frac{\exp(x_{j}^{\top}{\bm{\theta}})}{\sum_{i=1}^{p}\exp(x_{i}^{\top}{\bm{\theta}})}\quad\quad\text{for each $j\in[p]$,}$
so that a customer with preference vector $\bm{{\bm{\theta}}}$ chooses among the observed goods $x_{1},\ldots,x_{p}$ according to the probability vector $A({\bm{\theta}})$. This is an appropriate model for the choices of a single customer (or, more generally, for a group of customers with identical preferences). To model the behavior of a large number of consumers with heterogeneous preferences, Boyd and Mellman (1980) and Cardell and Dunbar (1980) suggested to model the population as consisting of a mixture of consumers with different taste vectors. The aggregate probabilities of individual goods being selected is then given by the softmax mixture model (2). This model has been broadly adopted throughout the management science and econometrics literature due to its flexibility and practicality, see (McFadden and Train, 2000; Johnston et al., 2017; Train, 2009; Cameron and Trivedi, 2005) and references therein.
Next word prediction in LLM. Open ended text continuation via LLM is now routinely obtained in response to a prompt of interest, one word at a time. Formally, the prompt is tokenized to yield $u_{1},\ldots,u_{m}\in\mathbb{R}^{L}$, for some initial values of these vectors. This sequence is run through a transformer-based model, initially introduced by Vaswani et al. (2017), to yield contextually embedded vectors $z_{1},\ldots,z_{m}$, of which one is chosen, say $z\in\mathbb{R}^{L}$. Given a vocabulary $x_{1},\ldots,x_{p}$ of vectors in $\mathbb{R}^{L}$ that are viewed as identifiers of the $p$ possible next words (we use tokens and words interchangeably here, although tokens are typically smaller units), the next predicted word is obtained by drawing from a probability on $p$ words with respective masses given by $A(x_{j}|z):=\exp{(z^{\top}x_{j})}/\sum_{i=1}^{p}\exp{(z^{\top}x_{i})}$, $j\in[p]$. This is the reason behind the well-known fact that re-running the LLM with the same prompt can yield different outcomes. In particular, running this process $N$ times, with the same prompt, will yield a sample $Y_{1},\ldots,Y_{N}$, of potentially different words. This sample can thus be viewed as $N$ independent observations on a discrete random variable $Y$, conditionally on the given $z$. Formally, if $Z$ is a latent, the conditional distribution of $Y$ given $Z=z$ is $A(y|z)$, for $y\in\{x_{1},\ldots,x_{p}\}$. If, further, we seek a summary of the complicated LLM process yielding $z$, we can assume that $Z\sim\rho:=\sum_{k=1}^{K}\alpha_{k}\delta_{{\bm{\theta}}_{k}}$, for ${\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}$ being the main directions in $\mathbb{R}^{L}$ explored in order to generate $z$, in response to the initial prompt. Then, the marginal distribution of $Y$ is a softmax mixture,
$Y\sim\pi(y):=\sum_{k=1}^{K}\alpha_{k}{\exp({y^{\top}{\bm{\theta}}_{k}})\over\sum_{j=1}^{p}\exp({{x_{j}}^{\top}{\bm{\theta}}_{k}})},\quad\quad y\in\{x_{1},\ldots,x_{p}\}.$
Estimation of the directions ${\bm{\theta}}_{k}$ and of their respective proportions can be thus used in any additional building block that attempts a correction of the LLM output towards a particular direction.
Finally, we note that our bounds on the rates of estimation of ${\bm{\omega}}^{*}$ trivially imply corresponding error bounds for estimation of $\pi({\bm{\omega}}^{*}):=(\pi(x_{1};{\bm{\omega}}^{*}),\ldots,\pi(x_{p};{\bm{\omega}}^{*}))^{\top}$ via the inequality
$\|\pi({\bm{\omega}})-\pi({\bm{\omega}}^{*})\|_{1}~\leq~\|\bm{\alpha}-\bm{\alpha}^{*}\|_{1}+\max_{k\in[K]}\max_{j\in[p]}|x_{j}^{\top}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*})|\,.$ (3)
Rates of estimation for $\pi({\bm{\omega}}^{*})$ can also be obtained more directly via maximum likelihood estimation (MLE), including through the nonparametric MLE approach (Kiefer and Wolfowitz, 1956), which is known to achieve minimax-optimal rates in related settings (Vinayak et al., 2019). Crucially, however, unlike the estimators we propose and analyze below, the direct computation of the MLE is generally intractable due to the non-concave nature of the log-likelihood function, and there is no known computationally efficient algorithm with sharp theoretical guarantees.
1.2 Our contributions
1.2.1 An EM algorithm for softmax mixtures with provable guarantees
[…]
In contrast to the practical success and popularity of the EM algorithm, its theoretical justification in a general context is scarce. It is often fairly easy to prove algorithmic convergence to a local optimum, but much harder to guarantee that the limit is a near global optimum of the sample likelihood. If the likelihood is unimodal, Wu (1983) shows that the EM algorithm converges to the global optimum under certain regularity conditions. When the likelihood is multimodal, which is typically the case for mixture models, the theoretical understanding of the EM algorithm is largely limited to the settings of Gaussian Mixture Models with $K$ components ($K$-GMM) and its variants. See, for instance, Xu et al. (2016); Daskalakis et al. (2017); Balakrishnan et al. (2017); Cai et al. (2019); Wu and Zhou (2021) for $K=2$, and Dasgupta and Schulman (2007); Yan et al. (2017); Zhao et al. (2020) for $K\geq 2$.
To the best of our knowledge, a theoretical analysis of the EM algorithm for softmax mixture models has not yet been developed. As we elaborate below in Remark 4 and 1 in Section 2.3, establishing convergence to a global maximum in the context of softmax mixtures presents significantly greater challenges than in the $K$-GMM case.
We begin by analyzing the convergence of the population level EM algorithm in Section 2.1. The convergence guarantees are given in Theorem 1, and discussed in the remarks following it. Corollary 1 is the first result that shows that softmax mixture models are locally identifiable. Our next result, stated in Theorem 2 of Section 2.2, shows that with high probability, once initialized within a $\delta_{0}$-neighborhood of any global optimum ${\bm{\omega}}^{*}$ of $\ell({\bm{\omega}})$, the expected value with respect to $\pi^{*}$ of the log-likelihood (5), the EM estimator $\widehat{\bm{\omega}}^{(t)}$ after $t$ iterations satisfies the following bound for all $t\geq 1$:
$d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq~\phi^{t}~d(\widehat{\bm{\omega}}^{(0)},{\bm{\omega}}^{*})+\delta_{N}$ (4)
for some $\phi\in(0,1)$ and some distance $d(\cdot,\cdot)$ defined later in (18). The first term on the right hand side reflects the algorithmic error while the second term $\delta_{N}$ represents the statistical error. In the former, a key quantity is the contraction rate $\phi$ which quantifies how fast the algorithmic error vanishes as the number of iterations increases. Our analysis reveals that $\phi$ depends on both the separation between the mixture parameters ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ and the condition number of the information matrix associated with each softmax mixture component. Under mild conditions on these quantities, the contraction rate satisfies $\phi<1$, which ensures that the EM algorithm converges linearly. We further show that the statistical error $\delta_{N}$ is of order $\sqrt{(L\log N)/N}$. Finally, our analysis characterizes the initialization conditions under which (4) holds, and shows that the size $\delta_{0}$ of the neighborhood $d(\widehat{\bm{\omega}}^{(0)},{\bm{\omega}}^{*})$ depends solely on certain properties of the feature set $\{x_{1},\ldots,x_{p}\}$. Designing an initialization scheme that satisfies such requirement is a challenging task in general. A common practical heuristic is to perform multiple random initializations and select the EM estimate that yields the highest likelihood (Dasgupta and Schulman, 2007). However, this approach typically requires $\mathcal{O}(\exp(L))$ initializations to succeed, which quickly becomes computationally infeasible as $L$ increases. In Section 3, we show that if we view $x_{1},\ldots,x_{p}$ as independent random draws from some distribution, then a Method-of-Moments (MoM) estimator can be constructed to provably satisfy the initialization requirement of the EM algorithm. Furthermore, estimators of second order latent moments of the mixing measure $\rho=\sum_{k=1}^{K}\alpha_{k}^{*}\delta_{{\bm{\theta}}_{k}^{*}}$ can be used to estimate the $K$-dimensional subspace of $\mathbb{R}^{L}$ spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. This can be combined with the random initialization heuristic: by sampling at random from this $K$-dimensional subspace of $\mathbb{R}^{L}$, the number of random initializations required for the success of EM is reduced to $\mathcal{O}(\exp(K))$; see Lemma 3 of Section 3.4.2.
1.2.2 Approximation and estimation of latent moments of softmax mixtures
In Section 3 we explain how to use and modify the general principles underlying the classical Method of Moments for softmax mixtures.
Lemma 1 of Section 3.1 below gives conditions under which the parameters of the mixture are uniquely determined by moments of the mixing measure $\rho=\sum_{k=1}^{K}\alpha_{k}\delta_{{{\bm{\theta}}}_{k}}$. It is a constructive result, in that the parameters are shown to be solutions of equations involving these moments, henceforth referred to as latent moments. Lemma 1 collects the existing results in Lindsay (1989), for univariate mixtures, and in Lindsay and Basak (1993), for multivariate mixtures.
In one-dimensional mixtures, with mixture components belonging to the so-called quadratic variance exponential families, with the Gaussian distribution as a chief example, the latent moments can be equated with moments of appropriate functionals of the observable data distribution, henceforth called observable moments; see, for instance, Tucker (1963) Brockett (1977), Lindsay (1989) for earlier references, and also  Wu and Yang (2020) for Gaussian mixtures and  Tian et al. (2017), for binomial mixtures. Extensions to the estimation of latent moments and mixed moments of multivariate mixtures are restricted to Gaussian mixtures Lindsay and Basak (1993). These results can be further combined with Lemma 1, to obtain method of moments (MoM) estimators of the mixture parameters, by replacing the latent moments with observable moment estimates.
It is not known how to construct moments of functionals of a softmax mixture $\pi^{*}(y)$ that equal the latent moments prescribed by Lemma 1, for softmax mixtures with generic design. However, in Proposition 1, the main result of Section 3, we show that we can construct functionals of $\pi^{*}(y)$ that lead to estimable accurate approximations of the latent moments, with expressions given in Section 3.2, when $p$ is large enough and the support points of the mixture $x_{1},\ldots,x_{p}$ are treated as a random sample from $\mu$, a continuous distribution on $\mathbb{R}^{L}$.
Solving the (population level) Lemma 1 with latent moments replaced by these approximations, gives solutions that are, using Proposition 2, close to the true mixture parameters. Using them as the initialization of a population level EM algorithm allows us to show, in Corollary 2, that the softmax mixture model is identifiable, for $p$ large enough. To the best of our knowledge this is the only proof, to date, of this fact, for $L>1$. For one-dimensional mixtures ($L=1$), identifiability follows from the (non-stochastic) classical arguments in Lindsay (1995), but the arguments cannot be extended to higher dimensions, as they make use of Chebyshev systems which unfortunately do not exist when $L>1$.
The final estimator of the latent moments required by Lemma 1 is given in Section 3.3 and leads to the construction of a MoM estimator for softmax mixture parameters.
Theorems 3, 4 and 5 give the rates of convergence for MoM, showing that it can indeed serve as a warm start for the EM algorithm. However, implementing the MoM requires knowledge of a direction $v\in\mathbb{S}^{L-1}$ (referred to as the primary axis), along which the projections of the parameters ${\bm{\theta}}_{1}^{*},\dots,{\bm{\theta}}_{K}^{*}$ are well separated. While it is possible to obtain a weak guarantee by selecting $v$ at random, the resulting estimation rates exhibit suboptimal scaling with the ambient dimension $L$ (see Section 3.4.1). Since $L$ is often much larger than $K$, we adapt our procedure in Section 3.4 to estimate the subspace of $\mathbb{R}^{L}$ spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, and show how this subspace can be used to select $v$ (Lemma 2), thereby removing the suboptimal dependence on $L$. Finally, in Lemma 3, we show that the same estimated subspace can be used to reduce the number of random initializations required for the EM algorithm. The latter is particularly relevant as it is common practice to start the EM algorithm with random draws and select the one with the highest likelihood.
This paper is organized as follows. Section 2 proposes a hybrid EM algorithm to estimate ${\bm{\omega}}^{*}$. It establishes local identifiability and near-parametric rates of convergence. Section 3 develops a method of moments estimation of ${\bm{\omega}}^{*}$ when the features $x_{i}$’s are viewed as random draws from a known distribution. The resulting estimator of ${\bm{\omega}}^{*}$ is shown to be consistent and can serve as a warm start for the EM algorithm. Application of the latent moment estimation procedure to the estimation of the subspace spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ is discussed in Section 3.4. The simulation study in Section 4 confirms our theoretical findings.
2 An EM algorithm for softmax mixtures with generic features: local identifiability and rates of convergence
This section is devoted to softmax mixture parameter estimation via the EM algorithm. The population-level EM algorithm and its convergence guarantees are presented in Section 2.1, along with an important implication of these results, the local identifiability of the softmax mixture model. The sample-level EM algorithm for parameter estimation together with its theoretical guarantees is stated in Section 2.2. We prove these results in Section 2.3.
2.1 Local identifiability of softmax mixtures with generic features
In this section we show that the softmax mixture model is locally identifiable, for any given set of support points $\{x_{1},\ldots,x_{p}\}$ of the softmax mixture. For any ${\bm{\omega}}=(\bm{\alpha},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$, let
$\ell({\bm{\omega}})$ $=\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*})\log\left(\pi(x_{j};{\bm{\omega}})\right)=\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*})\log\left(\sum_{k=1}^{K}\alpha_{k}{\exp\left(x_{j}^{\top}{\bm{\theta}}_{k}\right)\over\sum_{\ell=1}^{p}\exp\left(x_{\ell}^{\top}{\bm{\theta}}_{k}\right)}\right)$ (5)
be the negative cross-entropy, which is just the the expected value, under $\pi^{*}=\pi(\cdot;{\bm{\omega}}^{*})$, of the log-likelihood function of a single observation $Y$ from $\pi(y;{\bm{\omega}})$. For future reference, we write
${\bm{\omega}}^{*}\in\Omega^{*},\quad\Omega^{*}:=\Bigl{\{}{\bm{\omega}}:\ell({\bm{\omega}})=\max_{{\bm{\omega}}^{\prime}}~\ell({\bm{\omega}}^{\prime})\Bigr{\}}.$ (6)
The main result of this section is Theorem 1, which gives the population level construction and theoretical guarantees of an optimizer of $\ell({\bm{\omega}})$, via the EM algorithm. Since $\ell({\bm{\omega}})$ is not concave in ${\bm{\omega}}$, the EM algorithm aims to find a maximizer of it via iterative maximization of a so-called $Q$-function which is given below shortly. As an important consequence, Corollary 1 shows that any two optimizers ${\bm{\omega}}_{1}^{*}$ and ${\bm{\omega}}_{2}^{*}$ that are at a small distance of one another must coincide, and we give a precise quantification of this distance. This local identifiability result under softmax mixture models is, to the best of our knowledge, new in the literature.
We need to introduce additional quantities. First, for any ${\bm{\omega}}=(\bm{\alpha},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$, let $Z$ be the random vector taking values in the set $\{{\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{K}\}$ with corresponding probabilities in ${\bm{\alpha}}$. We define the conditional probability of $Z={\bm{\theta}}_{k}$ given $Y=x_{j}$, for any $k\in[K]$ and $j\in[p]$, as
$g({\bm{\theta}}_{k}\mid x_{j};{\bm{\omega}}):={\alpha_{k}A(x_{j};{\bm{\theta}}_{k})\over\pi(x_{j};{\bm{\omega}})}={\alpha_{k}A(x_{j};{\bm{\theta}}_{k})\over\sum_{a=1}^{K}\alpha_{a}A(x_{j};{\bm{\theta}}_{a})}.$ (7)
Second, we define the joint probability of $Z={\bm{\theta}}_{k}$ and $Y=x_{j}$ as
$\log f(x_{j},{\bm{\theta}}_{k};{\bm{\omega}})$ $:=~\log\mathbb{P}_{{\bm{\omega}}}\{Y=x_{j},Z={\bm{\theta}}_{k}\}$
$~=~\log(\alpha_{k})+x_{j}^{\top}{\bm{\theta}}_{k}-\log\left(\sum_{\ell=1}^{p}\exp\left({x_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right).$ (8)
Instead of maximizing $\ell({\bm{\omega}})$, the EM algorithm iteratively maximizes the following $Q$-function
$Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $~=\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*})\sum_{k=1}^{K}g({\bm{\theta}}_{k}^{\prime}\mid x_{j};{\bm{\omega}}^{\prime})\log f(x_{j},{\bm{\theta}}_{k};{\bm{\omega}})$ (9)
over its first argument ${\bm{\omega}}$. After we plug (7) and (2.1) in (9), we get
$Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $=\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*})\sum_{k=1}^{K}{\alpha_{k}^{\prime}A(x_{j};{\bm{\theta}}_{k}^{\prime})\over\pi(x_{j};{\bm{\omega}}^{\prime})}\left[\log(\alpha_{k})+x_{j}^{\top}{\bm{\theta}}_{k}-\log\left(\sum_{\ell=1}^{p}\exp\left({x_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right)\right].$ (10)
In the parlance of the EM algorithm literature, evaluating the $Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ function at a given ${\bm{\omega}}^{\prime}$ corresponds to the “E-step”, while maximizing over ${\bm{\omega}}$ is the “M-step”. Starting at some initial point ${\bm{\omega}}^{(0)}$, the classical population-level EM algorithm iterates as follows:
${\bm{\omega}}^{(t+1)}=\operatorname*{argmax}_{{\bm{\omega}}}~Q({\bm{\omega}}\mid{\bm{\omega}}^{(t)})\qquad\text{for}~t=0,1,2,\ldots$ (11)
until convergence.
For the problem at hand, the maximization in (11) over ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$ is a concave optimization problem. More specifically,
  • maximizing with respect to ${\bm{\alpha}}\in\Delta^{K}$ yields the closed-form solution: for $k\in[K]$,
    $\alpha_{k}^{(t+1)}$ $=\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*}){\alpha_{k}^{(t)}A(x_{j};{\bm{\theta}}_{k}^{(t)})\over\pi(x_{j};{\bm{\omega}}^{(t)})}:=M_{k}({\bm{\omega}}^{(t)}).$ (12)
  • maximization over ${\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}$ does not admit a closed-form solution, and we adopt a gradient-ascent step, which is often used in such circumstances. For all $k\in[K]$, let $\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{(t)})$ be the gradient of $Q({\bm{\omega}}\mid{\bm{\omega}}^{(t)})$ with respect to ${\bm{\theta}}_{k}$ in the first argument ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$. Given a chosen step size $\eta_{k}>0$, the M-step update for maximizing over ${\bm{\theta}}_{k}$ is given by
    ${\bm{\theta}}_{k}^{(t+1)}$ $={\bm{\theta}}_{k}^{(t)}+\eta_{k}~\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{(t)})\mid_{{\bm{\omega}}={\bm{\omega}}^{(t)}}$ (13)
    $={\bm{\theta}}_{k}^{(t)}+\eta_{k}~\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*}){\alpha_{k}^{(t)}A(x_{j};{\bm{\theta}}_{k}^{(t)})\over\pi(x_{j};{\bm{\omega}}^{(t)})}\left(x_{j}-\mathbf{X}^{\top}A({\bm{\theta}}_{k}^{(t)})\right)$ (14)
    where
    $\mathbf{X}^{\top}A({\bm{\theta}}_{k}^{(t)})=\sum_{j=1}^{p}x_{j}A(x_{j};{\bm{\theta}}_{k}^{(t)})={\sum_{j=1}^{p}x_{j}\exp(x_{j}^{\top}{\bm{\theta}}_{k}^{(t)})\over\sum_{\ell=1}^{p}\exp(x_{\ell}^{\top}{\bm{\theta}}_{k}^{(t)})}.$
Since the update in (12) is given in closed form, whereas (13) involves a gradient ascent step, the population-level EM algorithm for softmax mixtures can be viewed as a hybrid procedure.
In the following we show that for any maximizer ${\bm{\omega}}^{*}$ of $\ell({\bm{\omega}})$ that satisfies the separation condition in (19), the above EM-iterates ${\bm{\omega}}^{(t)}$, when initialized within a local neighborhood of ${\bm{\omega}}^{*}$, converge linearly to ${\bm{\omega}}^{*}$ as $t\to\infty$, with respect to a distance defined shortly below.
We begin by stating a condition on the feature matrix $\mathbf{X}=(x_{1}^{\top},\ldots,x_{p}^{\top})^{\top}\in\mathbb{R}^{p\times L}$ upon which the softmax mixture model is defined. For any ${\bm{\theta}}\in\mathbb{R}^{L}$ with $A({\bm{\theta}})\in\Delta^{p}$, we write $H_{{\bm{\theta}}}=\mathbf{X}^{\top}(\mathrm{diag}(A({\bm{\theta}}))-A({\bm{\theta}})A({\bm{\theta}})^{\top})\mathbf{X}\in\mathbb{R}^{L\times L}$ and denote by. We denote by $\lambda_{1}(M)\geq\cdots\geq\lambda_{d}(M)$ the eigenvalues of any symmetric, positive semidefinite matrix $M\in\mathbb{R}^{d\times d}$.
Assumption 1.
There exist some constants $0<\underline{\sigma}^{2}\leq\overline{\sigma}^{2}<\infty$ and $\varsigma^{2}<\infty$ such that for any ${\bm{\omega}}^{*}\in\Omega^{*}$, with $\Omega^{*}$ given by (6), all $a,b\in[K]$ and $u\in[0,1]$ with ${\bm{\theta}}=u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{b}^{*}$,
$\underline{\sigma}^{2}\leq\lambda_{L}(H_{\bm{\theta}})\leq\lambda_{1}(H_{\bm{\theta}})\leq\overline{\sigma}^{2}$ (15)
and
$\lambda_{1}\bigl{(}H_{{\bm{\theta}}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A({\bm{\theta}}))\mathbf{X}H_{{\bm{\theta}}}^{-1/2}\bigr{)}\leq\varsigma^{2}.$ (16)
The matrix $H_{{\bm{\theta}}}$ in 1 denotes the Fisher information matrix under a single softmax parametrization $A({\bm{\theta}})$. The first condition (15) ensures that $H_{{\bm{\theta}}}$ remains well-conditioned along the line segment connecting any pair of mixture components ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. The second condition (16) is technical, but follows from (15) and $\lambda_{1}(\mathbf{X}^{\top}\mathrm{diag}(A({\bm{\theta}}))\mathbf{X})\leq\varsigma^{2}/\underline{\sigma}^{2}$, that is, the $L\times L$ matrix $\mathbf{X}^{\top}\mathrm{diag}(A({\bm{\theta}}))\mathbf{X}$ is well-behaved. In Theorem 4 of Section 3.2, we verify that 1 holds with high probability when the rows of $\mathbf{X}$ are i.i.d. samples from a multivariate Gaussian distribution. A similar conclusion holds when the rows of $\mathbf{X}$ are i.i.d. sub-Gaussian vectors, provided that the population-level Fisher information matrix has its smallest eigenvalue bounded away from zero along the line segment between any two ${\bm{\theta}}_{a}^{*}$ and ${\bm{\theta}}_{b}^{*}$. For future reference, note that $\varsigma\geq 1$ and $\|\mathbf{X}\|_{\infty,2}=\max_{j\in[p]}\|x_{j}\|_{2}\geq\overline{\sigma}$.
We introduce the following quantities $\underline{\alpha},\overline{\alpha}\in(0,1)$ on the mixing probabilities of any ${\bm{\omega}}^{*}$:
$\underline{\alpha}\leq\min_{k\in[K]}\alpha_{k}^{*}\leq\max_{k\in[K]}\alpha_{k}^{*}\leq\overline{\alpha}.$ (17)
For any ${\bm{\omega}}$ and ${\bm{\omega}}^{\prime}$, we define their distance as
$d({\bm{\omega}},{\bm{\omega}}^{\prime})=\max\left\{\overline{\sigma}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2},~{1\over\underline{\alpha}}\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\right\}$ (18)
with $\overline{\sigma}$ defined in Assumption 1 above. The following theorem presents the convergence rate of the population-level EM updates with respect to the above distance.
Theorem 1 (Convergence of the population-level EM).
Grant 1. For any ${\bm{\omega}}^{*}\in\Omega^{*}$ given by (6) that satisfies the separation condition
$\underline{\sigma}^{2}\min_{k\neq k^{\prime}}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2}~\geq~C\left\{\log K+\log{\overline{\sigma}^{2}\over\underline{\sigma}^{2}}+\log{\overline{\alpha}\over\underline{\alpha}}\right\}$ (19)
for some absolute constant $C>0$, assume the initialization ${\bm{\omega}}^{(0)}$ satisfies
$d({\bm{\omega}}^{(0)},{\bm{\omega}}^{*})\leq\delta_{0}\qquad\text{ with }\quad\delta_{0}\leq{c_{0}\over\varsigma^{2}}{\overline{\sigma}\over\|\mathbf{X}\|_{\infty,2}}$ (20)
for some sufficiently small constant $c_{0}\in[0,1/2)$. Then, there exist some $0<\phi<1$ and step-sizes $\eta_{k}>0$, $k\in[K]$, such that the EM iterates ${\bm{\omega}}^{(t)}$ in (12) and (13) satisfy: for all $t\geq 0$,
$d({\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq~\phi^{t}~\delta_{0}.$
We outline the proof of Theorem 1 and discuss its technical challenges in Section 2.3. A few remarks on the results in Theorem 1 are provided below.
Remark 1 (Separation among softmax mixture components).
Convergence of the EM iterates requires the separation condition in (19) between the mixture components. Our analysis explicitly captures the dependence of this requirement on the number of mixture components $K$, the condition number $\overline{\sigma}^{2}/\underline{\sigma}^{2}$ of the information matrix, and the balancing ratio $\underline{\alpha}/\overline{\alpha}$ of the mixing probabilities. When any of these quantities are large, the required separation increases only logarithmically. As illustrated in Section 2.3, deriving such a mild separation requirement under softmax mixtures is highly non-trivial and presents significantly greater challenges than in the case of Gaussian mixture models. Even for Gaussian location mixtures on $\mathbb{R}^{L}$ with $K\geq 3$ components, the weakest known separation in terms of the squared Euclidean distances between mean vectors required for the EM algorithm to succeed is on the order of $L\wedge K$ (Yan et al., 2017; Zhao et al., 2020), whereas for Lloyd’s algorithm, it is of order $K/\underline{\alpha}$ (Lu and Zhou, 2016).
Remark 2 (Initialization).
It is well known that the EM algorithm is very sensitive to its starting value ${\bm{\omega}}^{(0)}$. Our analysis specifies the initialization requirement under softmax mixtures, as given in (20), and quantifies its dependence on the feature matrix. As we will discuss shortly, the bound of $\delta_{0}$ in (20) also characterizes the size of the neighborhood in which local identifiability holds. In Theorem 4 and Remark 7 of Section 3.2, we provide a more explicit bound on $\delta_{0}$ when $x_{1},\ldots,x_{p}$ are treated as i.i.d. realizations from a sub-Gaussian distribution.
Remark 3 (Effect of the step size).
Our theory also reveals that the step size $\eta_{k}$ cannot be chosen to be too large, in order to ensure convergence of the EM updates. On the other hand, choosing a smaller $\eta_{k}$ results in a slower convergence rate (i.e., $\phi$ gets closer to 1), but does not affect the final statistical accuracy of the sample-level EM algorithm, as shown in Section 2.2. The explicit choice of $\eta_{k}$ for our analysis along with the corresponding form of $\phi$ is given in our proof of Appendix B. We found that the choice of $\eta_{k}=1$ yields overall satisfactory results in our numerical experiments.
An important implication of Theorem 1 is the following local identifiability result for the softmax mixture model.
Corollary 1 (Local identifiability).
Grant 1. Suppose there exist two parameter points ${\bm{\omega}}_{1}^{*}$ and ${\bm{\omega}}_{2}^{*}$ such that $\pi^{*}=\pi({\bm{\omega}}_{1}^{*})=\pi({\bm{\omega}}^{*}_{2})$, and both satisfy (19) for their corresponding ${\bm{\theta}}_{k}^{*}$’s. If $d({\bm{\omega}}_{1}^{*},{\bm{\omega}}_{2}^{*})\leq\delta_{0}/2$, for $\delta_{0}$ given by (20), then ${\bm{\omega}}_{1}^{*}={\bm{\omega}}_{2}^{*}$.
Proof.
Fix any ${\bm{\omega}}^{(0)}$ that satisfies $d({\bm{\omega}}^{(0)},{\bm{\omega}}_{1}^{*})\leq\delta_{0}/2$. By triangle inequality, we also have $d({\bm{\omega}}^{(0)},{\bm{\omega}}_{2}^{*})\leq\delta_{0}$. By Theorem 1, $\lim_{t\rightarrow\infty}d({\bm{\omega}}^{(t)},{\bm{\omega}}_{1}^{*})=0=\lim_{t\rightarrow\infty}d({\bm{\omega}}^{(t)},{\bm{\omega}}_{2}^{*})$, and thus ${\bm{\omega}}_{1}^{*}={\bm{\omega}}_{2}^{*}$, by the uniqueness of the limit in metric spaces. ∎
Corollary 1, via Theorem 1, offers sufficient conditions for local identifiability of the softmax mixture model. The proof is constructive, and shows that any global maximizer ${\bm{\omega}}^{*}$ can be identified, via the proposed EM algorithm: any ${\bm{\omega}}^{\prime}$ that is observationally equivalent to ${\bm{\omega}}^{*}$ in the stated $(\delta_{0}/2)$ neighborhood must coincide with ${\bm{\omega}}^{*}$.
It is classically known (Rothenberg, 1971) that under weak regularity conditions local identifiability is equivalent to non-singularity of the information matrix for general parametric families. In the context of this paper, these conditions would therefore be relative to the mixture model. In contrast, our results in Theorem 1 and Corollary 1 on local identifiability under softmax mixtures rely on 1, a more transparent condition that depends only on the information matrix of a single softmax component, rather than that of the entire mixture. Moreover, the bound on $\delta_{0}$ in (20) provides an explicit quantification of the neighborhood within which local identifiability holds.
As mentioned in the introduction, although global identifiability (up to label switching) is more desirable, establishing it for the softmax mixtures with more than two mixture components remains a challenging problem in its own right; see the discussion in Zhao and Xia (2019); Hu (2022); Chierichetti et al. (2018); Tang (2020) for two mixture components. In Section 3.2, we establish such an identifiability result when the features are viewed as independent realizations from an underlying distribution.
2.2 EM parameter estimates of softmax mixtures with generic features: rates of convergence
We first state the hybrid EM algorithm for parameter estimation based on samples $Y_{1},\ldots,Y_{N}$ i.i.d. drawn from the softmax mixtures. Essentially, it follows from its population-level counterpart in Section 2.1 by replacing $\pi({\bm{\omega}}^{*})$ with the empirical frequency $\widehat{\pi}\in\Delta^{p}$ of each $x_{j}$ observed in the sample. Since the sample log-likelihood at any ${\bm{\omega}}$ equals
$\ell_{N}({\bm{\omega}})$ $={1\over N}\sum_{i=1}^{N}\log\left(\sum_{k=1}^{K}\alpha_{k}A(Y_{i};{\bm{\theta}}_{k})\right)=\sum_{i=1}^{p}\widehat{\pi}_{i}\log\left(\sum_{k=1}^{K}\alpha_{k}{\exp(x_{i}^{\top}{\bm{\theta}}_{k})\over\sum_{j=1}^{p}\exp(x_{j}^{\top}{\bm{\theta}}_{k})}\right),$ (21)
which is also not concave, we iteratively maximize the following sample-level $Q$-function
$\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})=\sum_{j=1}^{p}\widehat{\pi}_{j}\sum_{k=1}^{K}{\alpha_{k}^{\prime}A(x_{j};{\bm{\theta}}_{k}^{\prime})\over\pi(x_{j};{\bm{\omega}}^{\prime})}\left[\log(\alpha_{k})+x_{j}^{\top}{\bm{\theta}}_{k}-\log\left(\sum_{\ell=1}^{p}\exp\left({x_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right)\right].$ (22)
Starting at some initial point $\widehat{\bm{\omega}}^{(0)}$, the sample-level EM algorithm proceeds iteratively until convergence. For all $t\geq 0$ and $k\in[K]$, the updates are given by:
$\widehat{\alpha}_{k}^{(t+1)}$ $=\sum_{j=1}^{p}\widehat{\pi}_{j}{\widehat{\alpha}_{k}^{(t)}A(x_{j};\widehat{\bm{\theta}}_{k}^{(t)})\over\pi(x_{j};\widehat{\bm{\omega}}^{(t)})}:=\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)}),$ (23)
$\widehat{\bm{\theta}}_{k}^{(t+1)}$ $=\widehat{\bm{\theta}}_{k}^{(t)}+\eta_{k}~\sum_{j=1}^{p}\widehat{\pi}_{j}{\widehat{\alpha}_{k}^{(t)}A(x_{j};\widehat{\bm{\theta}}_{k}^{(t)})\over\pi(x_{j};\widehat{\bm{\omega}}^{(t)})}\left(x_{j}-\mathbf{X}^{\top}A(\widehat{\bm{\theta}}_{k}^{(t)})\right)$ (24)
with $\widehat{\bm{\omega}}^{(t+1)}=(\widehat{\bm{\alpha}}^{(t+1)},\widehat{\bm{\theta}}_{1}^{(t+1)},\ldots,\widehat{\bm{\theta}}_{K}^{(t+1)})$.
In the following, we state our theoretical guarantees on the convergence rate of the above sample-level EM updates.
Theorem 2.
Under 1, assume there exists some large absolute constant $C>0$ such that
${\underline{\alpha}N\over\log N}\geq C~{\overline{\alpha}\over\underline{\alpha}}{\overline{\sigma}^{2}\over\underline{\sigma}^{2}}{\|\mathbf{X}\|_{\infty,2}^{2}\over\underline{\sigma}^{2}}~KL.$ (25)
For any ${\bm{\omega}}^{*}$ satisfying (19), further assume the initialization $\widehat{\bm{\omega}}^{(0)}$ satisfies (20) with initial bound $\delta_{0}$. Then, there exist some $0<\phi<1$, some absolute constant $C^{\prime}>0$ and step-sizes $\eta_{k}>0$, $k\in[K]$, such that with probability at least $1-\mathcal{O}(N^{-L})$, the following holds for the whole sequence $\widehat{\bm{\omega}}^{(t)}$ in (23) – (24), with $t\geq 0$,
$d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq~\phi^{t}~\delta_{0}+C^{\prime}\sqrt{\overline{\alpha}KL\log N\over\underline{\alpha}^{2}N}.$ (26)
Theorem 2 states that the estimates $\widehat{\bm{\omega}}^{(t)}$, initialized from any $\widehat{\bm{\omega}}^{(0)}$ satisfying (20) and updated according to the steps in (23) and (24), converge at the rate specified in (26), with explicit dependence on $K$, $\underline{\alpha}$, $\overline{\alpha}$, and $L$. In the case of balanced mixing probabilities, where $\overline{\alpha}\asymp\underline{\alpha}$, the convergence rate simplifies to $\sqrt{KL\log(N)/(\underline{\alpha}N)}$, where $\underline{\alpha}N$ represents the smallest effective sample size across all mixture components. In light of this, condition (25) imposes a lower bound on this smallest sample size and is required for the convergence rate to vanish asymptotically. For fixed $K$ as considered in this paper, the rate further simplifies to $\sqrt{L\log(N)/N}$, which differs from the parametric rate for estimating an $L$-dimensional vector from $N$ i.i.d. samples by only a logarithmic factor. Moreover, we emphasize that the convergence rate in (26) holds individually for each quantity: $\max_{k\in[K]}\overline{\sigma}\|\widehat{\bm{\theta}}_{k}^{(t)}-{\bm{\theta}}_{k}^{*}\|_{2}$ and $\|\widehat{\bm{\alpha}}^{(t)}-{\bm{\alpha}}^{*}\|_{\infty}/\underline{\alpha}$, after $\mathcal{O}(\log N)$ iterations. Since the updates of $\widehat{\bm{\alpha}}^{(t)}$ in (23) also depend on $\widehat{\bm{\theta}}_{k}^{(t)}$, the convergence rate of $\|\widehat{\bm{\alpha}}^{(t)}-{\bm{\alpha}}^{*}\|_{\infty}$ is primarily determined by the rate of $\|\widehat{\bm{\theta}}_{k}^{(t)}-{\bm{\theta}}_{k}^{*}\|_{2}$. If one is interested in obtaining refined rates for estimating ${\bm{\alpha}}^{*}$, a natural approach is to refit by maximizing the likelihood in (21) over ${\bm{\alpha}}$, with ${\bm{\theta}}_{k}$ replaced by $\widehat{\bm{\theta}}_{k}^{(t)}$, and then appeal to the analysis in Bing et al. (2022).
2.3 Proofs of Theorems 1 & 2
Proof of Theorem 1.
The proof follows from that of Theorem 2 below, if we replace the quantities $\widehat{M}_{k}$, $\widehat{Q}$ and $\widehat{\bm{\omega}}^{(t)}$ by $M_{k}$, $Q$ and ${\bm{\omega}}^{(t)}$, respectively, and set $\epsilon_{N}=0$. ∎
Proof of Theorem 2.
The problem at hand is non-standard in that we are dealing with a hybrid between the standard EM for ${\bm{\alpha}}$ in step (23) and a first-order EM for ${\bm{\theta}}_{k}$, $k\in[K]$, in step (24). We use induction to prove that with the desired probability,
$d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$ $~\leq~\phi^{t}\delta_{0}+{1-\phi^{t}\over 1-\phi}\delta_{N},\qquad\forall~t\geq 0,$ (27)
for $\delta_{N}=\mathcal{O}(\epsilon_{N}/\underline{\alpha})$ with $\epsilon_{N}$ given in Lemma 6 and for some $\phi\in(0,1)$ with $\delta_{N}\leq(1-\phi)\delta_{0}$.
[…]
$|\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|\leq\sup_{{\bm{\omega}}\in{\mathbb{B}}({\bm{\omega}}^{*},\delta_{0})}|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|+|M_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$
as $d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq\delta_{0}$, Lemmas 6 and 5 imply that, for $\kappa$ given in Lemma 5,
$\|\widehat{\bm{\alpha}}^{(t+1)}-{\bm{\alpha}}^{*}\|_{\infty}\leq C\epsilon_{N}+\kappa~d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$
holds with probability $1-\mathcal{O}(N^{-L})$. We analyze the first-order EM-updates $\widehat{\bm{\theta}}_{k}^{(t+1)}$ in (24) as follows:
$\|\widehat{\bm{\theta}}_{k}^{(t+1)}-{\bm{\theta}}_{k}^{*}\|_{2}$ $=\|\widehat{\bm{\theta}}^{(t)}_{k}+\eta_{k}\nabla_{{\bm{\theta}}_{k}}\widehat{Q}(\widehat{\bm{\omega}}^{(t)}\mid\widehat{\bm{\omega}}^{(t)})-{\bm{\theta}}_{k}^{*}\|_{2}$
$\leq\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid{\bm{\omega}}^{*})\|_{2}+\eta_{k}\|\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid\widehat{\bm{\omega}}^{(t)})-\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid{\bm{\omega}}^{*})\|_{2}$
$\qquad+\eta_{k}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}(\widehat{\bm{\omega}}^{(t)}\mid\widehat{\bm{\omega}}^{(t)})-\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid\widehat{\bm{\omega}}^{(t)})\|_{2}$
Invoking Lemmas 5 and 6 gives that, with probability $1-\mathcal{O}(N^{-L})$,
$\|\widehat{\bm{\theta}}_{k}^{(t+1)}-{\bm{\theta}}_{k}^{*}\|_{2}$ $\leq\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}~q(\widehat{\bm{\omega}}^{(t)})\|_{2}+\eta_{k}\left(\overline{\sigma}\kappa~d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})+C\overline{\sigma}\epsilon_{N}\right).$
Here, we write $q_{k}({\bm{\omega}}):=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ with $q_{k}({\bm{\omega}}^{*})=0$, and its smoothness and strong-concavity properties are stated in Lemma 4. After we square the first term on the right and work out the squares, we find
$\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}q_{k}(\widehat{\bm{\omega}}^{(t)})\|_{2}^{2}$
$=\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}+\eta_{k}^{2}\|q_{k}(\widehat{\bm{\omega}}^{(t)})\|_{2}^{2}+2\eta_{k}(\widehat{\bm{\theta}}_{k}^{(t)}-{\bm{\theta}}_{k}^{*})^{\top}\left(q_{k}(\widehat{\bm{\omega}}^{(t)})-q_{k}({\bm{\omega}}^{*})\right)$
$\leq\left(1-{2\eta_{k}\mu_{k}\gamma_{k}\over\mu_{k}+\gamma_{k}}\right)\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}+\eta_{k}\left(\eta_{k}-{2\over\mu_{k}+\gamma_{k}}\right)\|q_{k}(\widehat{\bm{\omega}}^{(t)})\|_{2}^{2}$ by Lemma 4
$\leq\left({\mu_{k}-\gamma_{k}\over\mu_{k}+\gamma_{k}}\right)^{2}\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$ $\text{by }\eta_{k}={2\over\mu_{k}+\gamma_{k}}.$
Summarizing, we find with probability $1-\mathcal{O}(N^{-L})$ that
$d(\widehat{\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})$ $\leq C\max\left\{{1\over\underline{\alpha}},~\overline{\sigma}^{2}\max_{k}{\eta_{k}}\right\}\epsilon_{N}+\phi~d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$
where
$\phi$ $=\max_{k}{\mu_{k}-\gamma_{k}\over\mu_{k}+\gamma_{k}}+\kappa\max\left({1\over\underline{\alpha}},~\max_{k}{2\overline{\sigma}^{2}\over\mu_{k}+\gamma_{k}}\right)$
$\leq{(1+c_{0})\overline{\sigma}^{2}-(1-c_{0})\underline{\sigma}^{2}\over(1+c_{0})\overline{\sigma}^{2}+(1-c_{0})\underline{\sigma}^{2}}+{2\overline{\sigma}^{2}\over(1+c_{0})\overline{\sigma}^{2}+(1-c_{0})\underline{\sigma}^{2}}{\kappa\over\underline{\alpha}}$
$<1-{2(1-2c_{0})\underline{\sigma}^{2}\over(1+c_{0})\overline{\sigma}^{2}+(1-c_{0})\underline{\sigma}^{2}}$ since $\kappa<c_{0}\underline{\alpha}\,{\underline{\sigma}^{2}\over\overline{\sigma}^{2}}$ by (19)
<1<1< 1 since $c_{0}<1/2$
Now by setting $\delta_{N}=2C\epsilon_{N}/\underline{\alpha}\leq(1-2c_{0})\delta_{0}{\underline{\sigma}^{2}}/\overline{\sigma}^{2}\leq(1-\phi)\delta_{0}$ by (25), we obtain
$d(\widehat{\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})\leq\delta_{N}+\phi~d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq\phi^{t+1}\delta_{0}+{1-\phi^{t+1}\over 1-\phi}\delta_{N}$
so that (27) holds for $t+1$. This proves the induction step and the proof is complete. ∎
Remark 4.
We follow the road-map developed in Balakrishnan et al. (2017) for analyzing the EM algorithm for general mixture models. Specifically, we establish (a) the smoothness and strong concavity of ${\bm{\omega}}\mapsto Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ and (b) the Lipschitz continuity of ${\bm{\omega}}^{\prime}\mapsto\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ for all ${\bm{\omega}}$ in a local neighborhood of ${\bm{\omega}}^{*}$ and (c) the rate of convergence of $\max_{k\in[K]}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}$ uniformly over ${\bm{\omega}}$ within a size $\delta_{0}$-neighborhood of ${\bm{\omega}}^{*}$. Although these are high-level quantities, as the authors noted in Balakrishnan et al. (2017), the real challenge in analyzing EM-type algorithms lies in establishing properties (a), (b) and (c) under specific models. Their work demonstrates this framework for the standard 2-GMM and two of its variants. To the best of our knowledge, a theoretical analysis of the EM algorithm under softmax mixture models has not yet been developed. The establishment of properties (a), (b) and (c) proves to be significantly more challenging under softmax mixture models than in the GMM setting, see Example 1 below. Indeed, for property (a), the fact that $Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ are quadratic in ${\bm{\theta}}_{k}$ under the GMM implies that their gradient $\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ is linear in ${\bm{\theta}}_{k}$. As a result, the strong concavity and smoothness of $Q(\cdot\mid{\bm{\omega}}^{*})$ with respect to ${\bm{\theta}}_{k}$ follows immediately. In stark contrast, $\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ under the softmax mixture model is non-linear in ${\bm{\theta}}_{k}$, and its expression in (14) still involves $\mathsf{softmax}(x_{1}^{\top}{\bm{\theta}}_{k},\ldots,x_{p}^{\top}{\bm{\theta}}_{k})$. The strong concavity and smoothness of $Q(\cdot\mid{\bm{\omega}}^{*})$ in this setting are established in Lemma 4 of Appendix B, and require a careful perturbation analysis of several softmax-related quantities stated in Lemmas 7 and 8 of Section B.2. The difficulty is further elevated when establishing property (b), which concerns the Lipschitz continuity of $\|\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})\|_{2}$, for all ${\bm{\omega}}$ within a $\delta_{0}$-neighborhood of ${\bm{\omega}}^{*}$. This step involves the most technically demanding derivations, even in the simple case of the symmetric and isotropic 2-GMM (Balakrishnan et al., 2017), and extending the analysis to isotropic $K$-GMMs already requires substantial refinements (Yan et al., 2017). In 1 below, we illustrate that verifying property (b) for softmax mixtures – even in the case $K=2$ – is significantly more challenging than for the 2-GMM. For general $K\geq 2$, property (b) is established in Lemma 5 of Appendix B, building on several technical results presented in Lemmas 7, 8, 9 and 10 of Section B.2. Finally, since our EM algorithm employs a hybrid M-step to estimate both ${\bm{\alpha}}^{*}$ and ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, an analogous version of property (b) must also be verified for the closed-form update of ${\bm{\alpha}}^{(t)}$. This result is also stated in Lemma 5.
Existing analyses of property (c) under GMMs typically rely on empirical process techniques such as symmetrization and Ledoux and Talagrand-type contraction results (Balakrishnan et al., 2017; Yan et al., 2017; Cai et al., 2019). However, in the case of softmax mixture models, the Lipschitz conditions required for applying the Ledoux and Talagrand contraction are challenging to verify. Instead, we develop a carefully tailored discretization argument to establish the necessary uniform convergence guarantees in Lemma 6 of Appendix B.
In the following example, we illustrate the difficulty of verifying the Lipschitz continuity of the map ${\bm{\omega}}^{\prime}\mapsto\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ under softmax mixture models (Lemma 5) by comparing it to the Gaussian mixture model case in an even simplified setting.
Example 1.
We focus the discussion on two mixture components with equal weights. Start with a 2-GMM where the observable feature $Y\in\mathbb{R}^{p}$ comes from $\mathcal{N}_{p}((\eta/2){\bm{\theta}}^{*},\mathbf{I}_{p})$, conditioning on $\eta$, with $\mathbb{P}(\eta=1)=\mathbb{P}(\eta=-1)=1/2$. The only parameter is ${\bm{\theta}}^{*}\in\mathbb{R}^{p}$, with the separation between $Y\mid\eta=1$ and $Y\mid\eta=-1$ being $\|{\bm{\theta}}^{*}\|_{2}^{2}$. The M-step of the population-level EM algorithm uses the operator $M$ given by
$M({\bm{\theta}})={\mathbb{E}[\gamma(Y;{\bm{\theta}})Y]\over\mathbb{E}[\gamma(Y;{\bm{\theta}})]}$
while evaluating $\gamma(Y;{\bm{\theta}})=1/(1+\exp(-Y^{\top}{\bm{\theta}}))$ is the E-step. Establishing its contraction requires deriving the Lipschitz continuity of $M$, which in turn hinges on bounding the difference $|\mathbb{E}[\gamma(Y;{\bm{\theta}})-\gamma(Y;{\bm{\theta}}^{*})]|\leq\kappa\|{\bm{\theta}}-{\bm{\theta}}^{*}\|_{2}$ for some small $\kappa$. Derivation of $\kappa$ is intuitively simple as
${d\gamma(Y;{\bm{\theta}})\over d{\bm{\theta}}}={\exp(-Y^{\top}{\bm{\theta}})\over[1+\exp(-Y^{\top}{\bm{\theta}})]^{2}}Y={\exp(-(\eta/2){\bm{\theta}}^{\top}{\bm{\theta}}^{*}-W^{\top}{\bm{\theta}})\over[1+\exp(-(\eta/2){\bm{\theta}}^{\top}{\bm{\theta}}^{*}-W^{\top}{\bm{\theta}})]^{2}}Y$
for some $W\sim\mathcal{N}_{p}(0,\mathbf{I}_{p})$. When ${\bm{\theta}}$ is close to ${\bm{\theta}}^{*}$, by $W^{\top}{\bm{\theta}}\sim\mathcal{N}(0,\|{\bm{\theta}}\|_{2}^{2})$, the fraction in front of $Y$ can be bounded (in expectation) by $\exp(-c\|{\bm{\theta}}^{*}\|_{2}^{2})$, which leads to $\kappa\leq\exp(-c\|{\bm{\theta}}^{*}\|_{2}^{2})$.
Now consider the softmax mixtures with $K=2$, $\alpha_{1}=\alpha_{2}=1/2$ and ${\bm{\theta}}_{1}^{*}=-{\bm{\theta}}_{2}^{*}=:{\bm{\theta}}^{*}$. Recalling (14), bounding $\|\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})\|_{2}$ requires to bound the $\ell_{2}$-norm of
$\sum_{j=1}^{p}\pi(x_{j};{\bm{\omega}}^{*})\left({A(x_{j};{\bm{\theta}}^{\prime})\over A(x_{j};{\bm{\theta}}^{\prime})+A(x_{j};-{\bm{\theta}}^{\prime})}-{A(x_{j};{\bm{\theta}}^{*})\over A(x_{j};{\bm{\theta}}^{*})+A(x_{j};-{\bm{\theta}}^{*})}\right)(x_{j}-\mathbf{X}^{\top}A({\bm{\theta}}_{k}))$
where, explicitly,
${A(x_{j};{\bm{\theta}}^{\prime})\over A(x_{j};{\bm{\theta}}^{\prime})+A(x_{j};-{\bm{\theta}}^{\prime})}={1\over 1+\exp(-2x_{j}^{\top}{\bm{\theta}}^{\prime}){\sum_{\ell}\exp(x_{\ell}^{\top}{\bm{\theta}}^{\prime})\over\sum_{\ell}\exp(-x_{\ell}^{\top}{\bm{\theta}}^{\prime})}}.$
The derivative of the above term with respect to ${\bm{\theta}}^{\prime}$ is notably complex, and even when ignoring the ratio involving the summations over $\ell$ in the denominator, deriving a Lipschitz constant in terms of $\exp(-c\|{\bm{\theta}}^{*}\|_{2}^{2})$ remains highly non-trivial. This difficulty is further exacerbated when the mixing weights are unknown and the number of mixture components exceeds two, a case we address in Lemma 5 of Appendix B.
3 Latent moment estimation in softmax mixtures with random features, with applications to EM initialization
In Section 2, we studied the general setting where the support set $x:=\{x_{1},\ldots,x_{p}\}$ of the softmax mixture is deterministic, and showed in Theorems 1 and 2 that the parameters ${\bm{\omega}}^{*}=({\bm{\alpha}}^{*},{\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*})$ can be recovered by the EM algorithm when initialized within a $\delta_{0}$-neighborhood, as specified in (20). In this section, we treat $x$ as a realization of i.i.d. random vectors $X_{1},\ldots,X_{p}\sim\mu$, where $\mu$ is a distribution on $\mathbb{R}^{L}$. Accordingly, $\pi(y;{\bm{\omega}}^{*})$ is interpreted as a conditional distribution, which we emphasize by writing $\pi(y;{\bm{\omega}}^{*}|x)$. Conform to the parlance in the bootstrap literature, our statements in this section will hold either $\mu$-almost surely or in $\mu$-probability. For example, in Theorem 3 the dimension $L$ is fixed and hence $\mu$ is a fixed measure and its statement holds for $\mu$-almost all realizations $x$. In contrast, we consider the more general case $L=L(p)\to\infty$ in Theorem 4 and now $\mu=\mu_{L}$ is a sequence of measures. Now, we can only state its conclusion in $\mu$-probability, that is, there exist Borel sets $A_{p}$, $p\geq 1$, such that $\mathbb{P}\left[(X_{1},\ldots,X_{p})\in A_{p}\right]\to 1$ (More precisely, we show that $\mathbb{P}\left[(X_{1},\ldots,X_{p})\in A_{p}\right]\geq 1-p^{-s}$ for any $s\geq 1$). Our goal is to show that a method-of-moments (MoM) algorithm can recover ${\bm{\omega}}^{*}$ within a small neighborhood, so that the EM algorithm, when initialized using the MoM, recovers ${\bm{\omega}}^{*}$ at optimal statistical precision. We start with the population-level recovery of ${\bm{\omega}}^{*}$ in Section 3.2, which has model identifiability as a consequence, and then state the sample-level estimation results in Section 3.3.
3.1 Preliminaries
In this section we collect background results on population level parameter recovery from latent moments, in finite mixture models. We begin by recalling a fundamental result in Lemma 1. It shows that the mixture model parameters can be uniquely determined from the moments and mixed moments of the mixing measure defined below. The result is constructive, in that it provides explicit parameter expressions as functions of these moments. In the next section we will make use of these expressions for parameter estimation. Results (32) and (34) below can be found in Lindsay (1989), whereas (33) is implicit in Lindsay and Basak (1993), and we derive its explicit form here.
By the modeling assumption, the true parameters ${\bm{{\bm{\theta}}}}_{1}^{*},\dots,{\bm{{\bm{\theta}}}}_{K}^{*}$ are distinct. The arguments presented below rely on the existence of a unit vector $v\in\mathbb{S}^{L-1}$ such that the inner products ${v}^{\top}{\bm{{\bm{\theta}}}}_{1}^{*}$, $\dots$, ${v}^{\top}{\bm{{\bm{\theta}}}}_{K}^{*}$ are all different from each other; this vector is called the primary axis in Lindsay and Basak (1993). The existence of such a vector is guaranteed, as detailed in Section 3.4.1. For ease of presentation, we assume ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ are distinct in their first coordinates:
$\Delta(\theta_{11}^{*},\ldots,\theta_{1K}^{*}):=\min_{k\neq k^{\prime}}|\theta_{1k}^{*}-\theta_{1k^{\prime}}^{*}|>0.$ (28)
Let $\rho^{*}:=\sum_{k=1}^{K}\alpha_{k}^{*}\delta_{{\bm{\theta}}_{k}^{*}}$ be the $K$-atomic measure associated with ${\bm{\omega}}^{*}$. As explained below, one can first recover $\theta_{11}^{*},\ldots,\theta_{1K}^{*}$ and then use them to recover the remaining coefficients $\theta_{ik}^{*}$ for $2\leq i\leq L$ and $k\in[K]$, based on certain moments of $\rho^{*}$.
[…]
$m_{r}:=\mathbb{E}_{\rho^{*}}[Z_{1}^{r}]=\sum_{k=1}^{K}\alpha_{k}^{*}(\theta_{1k}^{*})^{r}.$ (29)
Similarly, consider the following mixed-moments: for $0\leq r\leq K-1$ and $2\leq i\leq L$
$m_{r1;i}:=\mathbb{E}_{\rho^{*}}\left[Z_{1}^{r}Z_{i}\right]=\sum_{k=1}^{K}\alpha_{k}^{*}(\theta_{1k}^{*})^{r}\theta_{ik}^{*}.$ (30)
The subscripts $r$ and $r1;i$ of $m$ are mnemonic of the fact that we consider either the $r$-th moment of $Z_{1}$, or moments of the product of $Z_{1}^{r}Z_{i}^{1}$. Let
$\bm{m}:=(m_{0},m_{1},\dots,m_{2K-1})^{\top},\quad\bm{m}_{1;i}:=(m_{01;i},\ldots,m_{(K-1)1;i})^{\top},\quad 2\leq i\leq L.$ (31)
The following lemma shows that ${\bm{\omega}}^{*}$ can be uniquely recovered from the moments in (31).
Lemma 1.
For any ${\bm{\omega}}^{*}\in\Omega^{*}$ satisfying (28) and $\min_{k}\alpha_{k}^{*}>0$, the system of equations given by (29) and (30) has a unique solution which equals to ${\bm{\omega}}^{*}$, up to label switching. Moreover, the solution can be found explicitly and is given by the expressions below.
  1. 1.
    The first coordinates $\theta_{11}^{*},\theta_{12}^{*},\ldots,\theta_{1K}^{*}$ are the unique $K$ roots of the degree $K$ polynomial $P(x)$, in one variable, given by
    $P(x):=\mbox{\rm det}\begin{pmatrix}1&m_{1}&\dots&m_{K}\\ m_{1}&m_{2}&\dots&m_{K+1}\\ \vdots&\vdots&&\vdots\\ m_{K-1}&m_{K}&\dots&m_{2K-1}\\ 1&x&\dots&x^{K}\end{pmatrix}.$ (32)
  2. 2.
    For each $k\in[K]$, the remaining $L-1$ coordinates $\{\theta_{ik}^{*}\}_{2\leq i\leq L}$ are uniquely given by
    $\theta_{ik}^{*}=\begin{pmatrix}m_{01;i}\\ \vdots\\ m_{(K-1)1;i}\end{pmatrix}^{\top}\begin{pmatrix}1&m_{1}&\dots&m_{K-1}\\ m_{1}&m_{2}&\dots&m_{K}\\ \vdots&\vdots&&\vdots\\ m_{K-1}&m_{K}&\dots&m_{2K-2}\end{pmatrix}^{-1}\begin{pmatrix}1\\ \theta_{1k}^{*}\\ \vdots\\ (\theta_{1k}^{*})^{K-1}\end{pmatrix}.$ (33)
  3. 3.
    The mixture weight vector ${\bm{\alpha}}^{*}$ is uniquely given by
    ${\bm{\alpha}}^{*}=\begin{pmatrix}1&1&\ldots&1\\ \theta_{11}^{*}&\theta_{12}^{*}&\ldots&\theta_{1K}^{*}\\ \vdots&\vdots&&\vdots\\ (\theta_{11}^{*})^{K-1}&(\theta_{12}^{*})^{K-1}&\ldots&(\theta_{1K}^{*})^{K-1}\end{pmatrix}^{-1}\begin{pmatrix}1\\ m_{1}\\ \vdots\\ m_{K-1}\end{pmatrix}.$ (34)
Remark 5.
Note that $\bm{\alpha}^{*}$ and ${{\bm{\theta}}_{1}^{*}},\dots,{{\bm{\theta}}_{K}^{*}}$ are uniquely determined given all moments of the form
$m_{r_{1},\dots,r_{L}}:=\mathbb{E}_{\rho^{*}}[(\bm{e}_{1}^{\top}Z)^{r_{1}}\cdots(\bm{e}_{L}^{\top}Z)^{r_{L}}]\quad\quad\forall~r_{1},\dots,r_{L}\in\mathbb{N}\,,$
where $\bm{e}_{1},\ldots,\bm{e}_{L}$ denote the canonical basis vectors in $\mathbb{R}^{L}$. Knowledge of these moments is equivalent to knowing $\mathbb{E}_{\rho^{*}}[(v^{\top}Z)^{r}]$ for all $v\in\mathbb{S}^{L-1}$ and $r\in\mathbb{N}$, which uniquely determines the measure $\rho^{*}$. The virtue of Lemma 1 lies in identifying a minimal collection of such moments. Indeed, one can show that any strict subset of $\bm{m}$ and $\{\bm{m}_{i}\}_{2\leq i\leq L}$ fails to uniquely identify $\rho^{*}$.
Lemma 1 in fact gives a constructive procedure of recovering ${\bm{\omega}}^{*}$ from appropriate moments of its $K$-atomic measure $\rho^{*}$. When the atomic measure $\rho^{*}$ is interpreted as the mixing measure inducing $\pi(y;{\bm{\omega}}^{*}|x)$, its moments, now viewed as latent, must be estimated from the observable moments of $\pi(y;{\bm{\omega}}^{*}|x)$. The latter is the main novelty in our MoM procedure, and is explained in the next section.
3.2 Population-level global parameter recovery of softmax mixtures with random features
In the following we show that the latent moments and mixed-moments $\bm{m}$ and $\bm{m}_{1;i}$ used by Lemma 1, can be approximated from the moments of $\pi(\cdot;{\bm{\omega}}^{*}|x)=:\pi^{*}(\cdot|x)$. This construction is one of our main contributions. Moreover, it leads to a MoM algorithm that provably yields a parameter $\bar{\bm{\omega}}$ that is within a small neighborhood of ${\bm{\omega}}^{*}$, thereby enabling the EM algorithm, when initialized with this $\bar{\bm{\omega}}$, to recover ${\bm{\omega}}^{*}$ exactly. To illustrate the general idea of approximating the latent moments, it is enough to consider $m_{r}$, given by (29) above, of the first coordinate of $\rho^{*}$. Pick any $r\in\mathbb{N}$. Since information about ${\bm{\omega}}^{*}$ is in $Y\sim\pi^{*}(\cdot|x)$, we are naturally lead to searching for a function $h_{r}:\mathbb{R}^{L}\to\mathbb{R}$ such that
$\bar{m}_{r}:=$ $~\mathbb{E}_{\pi({\bm{\omega}}^{*}\mid x)}[h_{r}(Y)]=~\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}h_{r}(x_{j})\exp(x_{j}^{\top}{\bm{\theta}}_{k}^{*})\over{1\over p}\sum_{\ell=1}^{p}\exp(x_{\ell}^{\top}{\bm{\theta}}_{k}^{*})}$ (35)
is close to $m_{r}=\sum_{k=1}^{K}\alpha_{k}^{*}(\theta_{1k}^{*})^{r}$. Since $x_{1},\ldots,x_{p}$ are i.i.d. realizations from $\mu$, it is therefore enough to construct a function $h_{r}$ such that, for $X\sim\mu$ and a generic ${\bm{\theta}}\in\mathbb{R}^{L}$, we have
${\mathbb{E}_{\mu}\left[h_{r}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]}=\theta_{1}^{r},$ (36)
implying that the right hand side of (35) will be the $\mu$-a.s. limit of $m_{r}$. Similarly, we also need to construct appropriate functions $h_{r1;i}$ for $2\leq i\leq L$, such that
${\mathbb{E}_{\mu}\left[h_{r1;i}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]}=\theta_{1}^{r}\theta_{i}^{r}.$ (37)
This will ensure $\bar{m}_{r1;i}:=\mathbb{E}_{\pi({\bm{\omega}}^{*}|x)}[h_{r1;i}(Y)]$ is close to the mixed-moments $m_{r1;i}$.
It is not clear whether functions $h_{r}$ and $h_{r1;i}$ satisfying (36) and (37) exist for all $\mu$. However, under the following assumption on $\mu$, Proposition 1 below establishes their existence and provides explicit expressions for these functions. Its proof is given in Section C.2. To the best of our knowledge, this is a novel result.
Assumption 2.
$\mu$ is a strictly positive $C^{\infty}$ density on $\mathbb{R}^{L}$ whose moment generating function is finite everywhere. The mixed partial derivatives of $\mu$ of all orders decay super-exponentially at infinity.
Proposition 1.
Let $X\sim\mu$ with $\mu$ satisfying 2. For any $r\in\mathbb{N}$ and $i\in\{2,\ldots,L\}$, define, for all $x\in\mathbb{R}^{L}$,
$h_{r}(x)$ $~:=~(-1)^{r}{1\over\mu(x)}~\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\mu(x+t\mathbf{e}_{1})\big{|}_{t=0},$ (38)
$h_{r1;i}(x)$ $~:=~(-1)^{r+1}{1\over\mu(x)}~\frac{\mathrm{d}^{r+1}}{\mathrm{d}t^{r}\mathrm{d}s}\mu(x+t\mathbf{e}_{1}+s\mathbf{e}_{i})\big{|}_{t,s=0}.$ (39)
Then for any given ${\bm{\theta}}\in\mathbb{R}^{L}$, both (36) and (37) hold.
Example 2 (Explicit choices of $h_{r}$).
When $\mu=\mathcal{N}_{L}(a,\Sigma)$, the functions given by (38) or (39) take more familiar forms, and can be expressed in terms of the classical probabilist’s Hermite polynomials $\{H_{r}\}_{r\geq 0}$, defined by
$H_{r}(x)=r!\sum_{b=0}^{\lfloor r/2\rfloor}\frac{(-1)^{b}}{b!(r-2b)!2^{b}}x^{r-2b},\qquad\forall~x\in\mathbb{R}.$ (40)
Then,
$h_{r}(x):=h_{r}(x;a,\Sigma)=\|\Sigma^{-1/2}\mathbf{e}_{1}\|_{2}^{r}H_{r}\left((x-a)^{\top}\Sigma^{-1}\mathbf{e}_{1}/\|\Sigma^{-1/2}\mathbf{e}_{1}\|_{2}\right).$
When $\mu$ is a finite Gaussian mixture $\sum_{j=1}^{J}\lambda_{j}\mathcal{N}_{L}(a_{j},\Sigma_{j})$, then
$h_{r}(x)=\frac{\sum_{j=1}^{J}\lambda_{j}\mu^{(j)}(x)~h_{r}(x;a_{j},\Sigma_{j})}{\sum_{j=1}^{J}\lambda_{j}\mu^{(j)}(x)}.$
Here $\mu^{(j)}$ is the density of $\mathcal{N}_{L}(a_{j},\Sigma_{j})$.
Proposition 1 readily implies that the observable moments $\bar{m}_{r}$ in (35) are close to the true moments $m_{r}$ (and similarly, $\bar{m}_{r1;i}$ to $m_{r1;i}$) in the following sense:
$\bar{m}_{r}-m_{r}$ $=\sum_{k=1}^{K}\alpha_{k}^{*}\left\{{{1\over p}\sum_{i=1}^{p}[h_{r}(x_{i})\exp(x_{i}^{\top}{\bm{\theta}}_{k}^{*})]\over{1\over p}\sum_{j=1}^{p}[\exp(x_{j}^{\top}{\bm{\theta}}_{k}^{*})]}-{\mathbb{E}_{\mu}[h_{r}(X)\exp(X^{\top}{\bm{\theta}}_{k}^{*})]\over\mathbb{E}_{\mu}[\exp(X^{\top}{\bm{\theta}}_{k}^{*})]}\right\}$ (41)
The population-level MoM algorithm thus recovers ${\bm{\omega}}^{*}$ based on applying a variant version of Lemma 1 to $\bar{\bm{m}}=(\bar{m}_{0},\ldots,\bar{m}_{2K-1})^{\top}$ and $\bar{\bm{m}}_{1;i}=(\bar{m}_{01;i},\ldots,\bar{m}_{(K-1)1;i})^{\top}$, as detailed below.
To recover the first coordinates $\theta_{11}^{*},\ldots,\theta_{1K}^{*}$, Lemma 1 requires solving for the $K$ roots of a polynomial that uses $\bar{m}_{r}$ in place of $m_{r}$. This in turn requires the entries of $\bar{\bm{m}}$ to be bona fide moments of a distribution, a condition that is not guaranteed in general. This is discussed in detail, for general mixtures, in Lindsay (1989), together with potential corrections that may be difficult to implement. An alternative approach, in the univariate case, was developed by Wu and Yang (2020), who proposed to project the moments $\bar{\bm{m}}$ onto the set $\mathcal{M}$ of valid moments.
We adopt a similar strategy below, and begin by making the following assumption, that will be used for the remaining of the paper.
Assumption 3.
There exists some known constant $B<\infty$ such that $\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}\|_{2}\leq B.$
3 in conjunction with 2 ensures that $\mathbb{E}_{\mu}[X_{j}^{\top}{\bm{\theta}}_{k}^{*}]=\mathcal{O}(1)$ for all $j\in[p]$ and $k\in[K]$ so that the probabilities in $A({\bm{\theta}}_{k}^{*})$ are not too spiky. This is crucial in order to have the softmax parametrization be meaningful, as pointed out in Arora et al. (2016).
[…]
:=$\mathcal{M}:=\left\{(M_{1}(\nu),\dots,M_{2K-1}(\nu)):{\rm supp}(\nu)\in[-B,B]\right\}.$ (42)
The projection of $\bar{\bm{m}}$ onto this space is obtained by solving
$\widetilde{\bm{m}}=\operatorname*{argmin}_{{\bm{u}}\in\mathcal{M}}\|{\bm{u}}-\bar{\bm{m}}\|_{2}.$ (43)
Crucially, as Wu and Yang (2020) observed, the optimization problem in (43) can be written as a semi-definite program, which can be solved in polynomial time (Vandenberghe and Boyd, 1996). We remark that only the moments in $\bar{\bm{m}}$ need to be projected onto $\mathcal{M}$, and not the mixed moments $\bar{\bm{m}}_{1;i}$.
Now let $\widetilde{P}(x)$ be the degree $K$ polynomial obtained by replacing $m_{r}$ in (32) by $\widetilde{m}_{r}$, the $r$-th entry of $\widetilde{\bm{m}}$, for each $r\in\{1,\ldots,2K-1\}$. The $K$ roots of $\widetilde{P}$, denoted by $\bar{\theta}_{11},\dots,\bar{\theta}_{1K}$, are the recovered first coordinates by the population-level MoM.
To recover the remaining coordinates, we consider counterparts of (33) and (34) of Lemma 1. First, for all $i\in\{2,\ldots,L\}$ and $k\in[K]$, we define the preliminary parameter $\bar{\theta}_{ik}^{\prime}$ by
$\bar{\theta}_{ik}^{\prime}=\begin{pmatrix}\bar{m}_{01;i}\\ \vdots\\ \bar{m}_{(K-1)1;i}\end{pmatrix}^{\top}\begin{pmatrix}1&\widetilde{m}_{1}&\dots&\widetilde{m}_{K-1}\\ \widetilde{m}_{1}&\widetilde{m}_{2}&\dots&\widetilde{m}_{K}\\ \vdots&\vdots&&\vdots\\ \widetilde{m}_{K-1}&\widetilde{m}_{K}&\dots&\widetilde{m}_{2K-2}\end{pmatrix}^{\dagger}\begin{pmatrix}1\\ \bar{\theta}_{1k}\\ \vdots\\ (\bar{\theta}_{1k})^{K-1}\end{pmatrix}.$ (44)
[…]
$\bar{\bm{\alpha}}=\Pi_{\Delta^{K}}\left\{\begin{pmatrix}1&1&\ldots&1\\ \bar{\theta}_{11}&\bar{\theta}_{12}&\ldots&\bar{\theta}_{1K}\\ \vdots&\vdots&&\vdots\\ (\bar{\theta}_{11})^{K-1}&(\bar{\theta}_{12})^{K-1}&\ldots&(\bar{\theta}_{1K})^{K-1}\end{pmatrix}^{\dagger}\begin{pmatrix}1\\ \widetilde{m}_{1}\\ \vdots\\ \widetilde{m}_{K-1}\end{pmatrix}\right\}\,,$ (45)
where $\Pi_{\Delta^{K}}$ is the projection operator to the simplex $\Delta^{K}$.
We summarize in Algorithm 1 this analogue of Lemma 1 that recovers the parameter based on the approximated moments of its corresponding mixing measure.
Algorithm 1 Parameter Recovery via Approximated Latent Moments
1:The moment vectors $\bar{\bm{m}}\in\mathbb{R}^{2K}$, $\bar{\bm{m}}_{1;2},\ldots,\bar{\bm{m}}_{1;L}\in\mathbb{R}^{K}$ and a positive constant $B>0$.
[…]
3:  Compute the projected moment vector $\widetilde{\bm{m}}$ as in (43).
4:  Solve the $K$ roots $\bar{\theta}_{11},\ldots,\bar{\theta}_{1K}$ from $\widetilde{P}(x)=0$ with $\widetilde{P}(x)$ using $\widetilde{\bm{m}}$ in place of $\bm{m}$.
[…]
6:  Solve for the weights $\bar{\bm{\alpha}}$ from (45).
7:  return The mixing weights $\bar{\bm{\alpha}}\in\Delta^{K}$ and the vectors $\bar{\bm{\theta}}_{1},\ldots,\bar{\bm{\theta}}_{K}\in\mathbb{R}^{L}$.
8:end procedure
Let $\bar{\bm{\omega}}=(\bar{\bm{\alpha}},\bar{\bm{\theta}}_{1},\ldots,\bar{\bm{\theta}}_{K})$ be the output of Algorithm 1. In the following, we quantify its distance to ${\bm{\omega}}^{*}$ in terms of the difference between $\bar{\bm{m}}$ and $\bm{m}$. We need additional separation conditions between mixture components.
Assumption 4.
There exists a constant $\Delta_{1}>0$ such that $\Delta(\theta_{11}^{*},\ldots,\theta_{1K}^{*})>\Delta_{1}$.
Assumption 5.
The quantity $\underline{\alpha}$ in (17) is bounded away from zero.
4 requires the first coordinates in ${\bm{\theta}}_{k}^{*}$ are well-separated while 5 ensures that the mixing probabilities in ${\bm{\alpha}}^{*}$ are non-degenerate. We refer to Remark 10 in Section C.4 for discussion when such conditions are not met.
The following proposition is purely deterministic, and shows that the error of estimating both mixing components and weights is of the same order as that of estimating the moments. The proof of Proposition 2 reveals that, under the stated assumptions, its conclusion is valid for any finite mixture estimated by the classical method proposed by Lindsay (1989) and Lindsay and Basak (1993). Although partial results can be extracted from existing proofs, we are not aware of a complete, deterministic, result valid for high-dimensional mixture models, and we provide it below. We give more comments in Remark 10.
Proposition 2.
Grant Assumptions 2, 3, 4 and 5. There exists some constant $D$, depending on $K,B,\Delta_{1}$ and $\underline{\alpha}$, such that, up to relabeling,
$\|\bar{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{2}$ $~\leq~D\|\bar{\bm{m}}-{\bm{m}}\|_{2},$
$\max_{k\in[K]}\|\bar{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$ $~\leq~D\Bigl{(}L\|\bar{\bm{m}}-{\bm{m}}\|^{2}_{2}+\sum_{i=2}^{L}\|\bar{\bm{m}}_{1;i}-\bm{m}_{1;i}\|^{2}_{2}\Bigr{)}.$
Proof.
Its proof can be found in Section C.3. ∎
The constant $D$ can be shown to scale as $\Delta_{1}^{-cK}$ for some absolute constant $c$ and this scaling is tight; see Remark 10 in Section C.4.
We are now ready to state our global parameter recovery results. Recall the distance $d({\bm{\omega}},{\bm{\omega}}^{\prime})$ in (18) and the quantities $\varsigma$ and $\overline{\sigma}$ defined in 1. If $d(\bar{\bm{\omega}},{\bm{\omega}}^{*})\leq\delta_{0}\leq c_{0}\varsigma^{-2}\overline{\sigma}\|\mathbf{X}\|_{\infty,2}^{-1}$, that is, $\bar{\bm{\omega}}$ meets the initialization requirement (20), Theorem 1 states that the population-level EM iterates ${\bm{\omega}}^{(t)}$ in (12) and (13), initialized by $\bar{\bm{\omega}}$, recover ${\bm{\omega}}^{*}$, that is, $\lim_{t\to\infty}d({\bm{\omega}}^{(t)},{\bm{\omega}}^{*})=0$. In view of Proposition 2, we need to find the rates $\epsilon_{p}$ for
$\max_{r<2K}|\bar{m}_{r}-m_{r}|+\max_{r<K,2\leq i\leq L}|\bar{m}_{r1;i}-m_{r1;i}|\leq\epsilon_{p}$ (46)
and show that $\epsilon_{p}\sqrt{L}\ll\delta_{0}$. We observe that Assumptions 3, 4 & 5 are mild conditions on the parameter space. 2 states that $\mu$ is smooth with super-exponential tails. Since 1 depends on $X_{1},\ldots,X_{p}$, we formulate its population counterpart. The $L\times L$ information matrix is given by
$H_{{\bm{\theta}}}^{(\mu)}={\mathbb{E}_{\mu}[XX^{\top}\exp(X^{\top}{\bm{\theta}})]\over\mathbb{E}_{\mu}[\exp(X^{\top}{\bm{\theta}})]}-{\mathbb{E}_{\mu}[X\exp(X^{\top}{\bm{\theta}})]\mathbb{E}_{\mu}[X\exp(X^{\top}{\bm{\theta}})]^{\top}\over(\mathbb{E}_{\mu}[\exp(X^{\top}{\bm{\theta}})])^{2}}.$ (47)
Assumption 6.
There exist constants $0<\underline{\sigma}^{2}\leq\overline{\sigma}^{2}<\infty$ and $\varsigma^{2}<\infty$ such that $\underline{\sigma}^{2}\leq\lambda_{L}(H_{{\bm{\theta}}}^{(\mu)})\leq\lambda_{1}(H_{{\bm{\theta}}}^{(\mu)})\leq\overline{\sigma}^{2}$ and
$\lambda_{1}\left([H_{\bm{\theta}}^{(\mu)}]^{-1/2}{\mathbb{E}_{\mu}[XX^{\top}\exp(X^{\top}{\bm{\theta}})]\over\mathbb{E}_{\mu}[\exp(X^{\top}{\bm{\theta}})]}[H_{\bm{\theta}}^{(\mu)}]^{-1/2}\right)\leq\varsigma^{2}$ (48)
for all ${\bm{\theta}}=u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{b}^{*}$ with $u\in[0,1]$ and $a,b\in[K]$.
We will distinguish between two cases: (a) $\mu$ is a fixed measure and (b) $\mu$ depends on $p$.
In case (a), the sequence $X_{1},X_{2},\ldots$ are i.i.d. from a fixed distribution $\mu$ on $\mathbb{R}^{L}$. This implies that $L$ is fixed and the rate for $\epsilon_{p}$ in Eq. 46 is of order $\mathcal{O}(\{(\log\log p)/p\}^{1/2})$ by the Law of the Iterated Logarithm. 6 implies that 1 holds, $\mu$-almost surely, with $\underline{\sigma}^{2}/2$ and $2\overline{\sigma}^{2}$ in place of $\underline{\sigma}^{2}$ and $\overline{\sigma}^{2}$, and with $2\varsigma^{2}\leq C(\overline{\sigma}^{2},B)$ in place of $\varsigma^{2}$. Finally, 2 implies that $\|\mathbf{X}\|_{\infty,2}=\mathcal{O}(\log p)$, $\mu$-almost surely.
Theorem 3.
Assume $\mu$ is fixed and satisfies 2. Assume that ${\bm{\omega}}^{*}$ satisfies Assumptions 3, 4, 5, 6 and condition (19). Then, almost surely,
(1) the population-level MoM estimator satisfies $d(\bar{\bm{\omega}},{\bm{\omega}}^{*})=\mathcal{O}\left(\sqrt{L\log\log p/p}\right)$
(2) the EM-iterations ${\bm{\omega}}^{(t)}$, initialized at $\bar{\bm{\omega}}$, satisfy $\lim_{t\to\infty}d({\bm{\omega}}^{(t)},{\bm{\omega}}^{*})=0$, for all but finitely many $p$.
From Theorem 3 and Corollary 1, we can actually conclude that the softmax mixture model is identifiable in the following sense.
Corollary 2.
Assume $\mu$ is fixed and satisfies 2. Suppose ${\bm{\omega}}^{\dagger}$ and ${\bm{\omega}}^{*}$ satisfy Assumptions 3, 4, 5, 6 and condition (19). Then we have ${\bm{\omega}}^{*}={\bm{\omega}}^{\dagger}$ if and only if $\pi(\cdot;{\bm{\omega}}^{*}|x)=\pi(\cdot;{\bm{\omega}}^{\dagger}|x)$ with $\mu$-probability one.
Proof.
If $\pi(\cdot;{\bm{\omega}}^{*}|x)=\pi(\cdot;{\bm{\omega}}^{\dagger}|x)$ with $\mu$-probability one, then the moments (35) are equal. Theorem 3 and the triangle inequality further imply that $d({\bm{\omega}}^{*},{\bm{\omega}}^{\dagger})\leq\delta_{0}/2$, with probability one, for all $p$ large enough, and Corollary 1 forces, with probability one, $d({\bm{\omega}}^{*},{\bm{\omega}}^{\dagger})=0$. ∎
Case (b) is more challenging since $\mu$ changes with $p$ and we can no longer make almost sure statements. Instead, we will state finite sample result. We start with $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. This enables us to give explicit computations of the matrix $H_{{\bm{\theta}}}^{(\mu)}$ in 6 to verify 1. The rate for $\epsilon_{p}=\mathcal{O}(\sqrt{\log p/p})$ and the Gaussian tails of $\mu$ imply that $\|\mathbf{X}\|_{\infty,2}=\mathcal{O}(\sqrt{L}+\sqrt{\log p})$ with overwhelming probability.
Theorem 4.
Assume $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$ and Assumptions 3, 4 & 5 and (19) hold. Then, with probability at least $1-p^{-s}$, for sufficiently large $p\geq p(B,\overline{\sigma},s)$ and any $s>1$,
(1) the population-level MoM estimator satisfies $d(\bar{\bm{\omega}},{\bm{\omega}}^{*})=\mathcal{O}\left(\sqrt{L\log p/p}\right)$
(2) the EM-iterations ${\bm{\omega}}^{(t)}$, initialized at $\bar{\bm{\omega}}$, satisfy $\lim_{t\to\infty}d({\bm{\omega}}^{(t)},{\bm{\omega}}^{*})=0$.
Proof.
The proof of part (1) requires establishing finite-sample deviation inequalities for (46), which depend on random quantities such as $\sum_{j=1}^{p}H_{r}(X_{j})\exp(X_{j}^{\top}{\bm{\theta}})$ with $r<2K$, where $H_{r}$ denotes the Hermite polynomials defined in (40). Such analysis is complicated by the presence of $\exp(X_{j}^{\top}{\bm{\theta}})$, which arises from the softmax parametrization. Proving part (2) requires verifying 6 for $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$, and hence establishing 1. We defer the full proof to Section C.6. ∎
Remark 6 (Extension to $\mathcal{N}_{L}(0,\Sigma)$).
For $\mu=\mathcal{N}_{L}(0,\Sigma)$, suppose there exists constants $0<\underline{\sigma}^{2}\leq\overline{\sigma}^{2}<\infty$ such that $\underline{\sigma}^{2}\leq\lambda_{L}(\Sigma)\leq\lambda_{1}(\Sigma)\leq\overline{\sigma}^{2}$. Note that assuming $\mu$ has mean zero can be made without loss of generality, since subtracting the same constant from $x_{j}^{\top}{\bm{\theta}}_{k}$ for all $j\in[p]$ does not affect the value of $A({\bm{\theta}}_{k})$. In Section C.7 we show that one can continue using Algorithm 1 with $h_{r}$ and $h_{r1;i}$ chosen as (2) with $a=0$ and $\Sigma=\mathbf{I}_{L}$. Consequently, the MoM output $\bar{\bm{\alpha}}$ still approximates ${\bm{\alpha}}^{*}$ whereas $\bar{\bm{\theta}}_{1},\ldots,\bar{\bm{\theta}}_{K}$ approximates $\Sigma{\bm{\theta}}_{1}^{*},\ldots,\Sigma{\bm{\theta}}_{K}^{*}$, so that the rescaled version $\Sigma^{-1}\bar{\bm{\theta}}_{k}$ satisfies
$\max_{k\in[K]}\|\Sigma^{-1}\bar{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq(C^{\prime}/\underline{\sigma}^{2})\sqrt{L\log p/p}.$
On the other hand, the EM guarantees remain valid, as both 1 and 6 can be verified to hold with high probability (see Lemmas 15, 16 and 17). Consequently, Theorem 4 continues to hold with $(\bar{\bm{\alpha}},\Sigma^{-1}\bar{\bm{\theta}}_{1},\ldots,\Sigma^{-1}\bar{\bm{\theta}}_{K})$ in place of $\bar{\bm{\omega}}$.
Remark 7 (Extension to sub-Gaussian distributions).
A careful inspection of the proof reveals that the same conclusion in Theorem 4 holds when $\mu$ is any sub-Gaussian distribution with a finite sub-Gaussian constant, provided that the corresponding $H_{\theta}^{(\mu)}$ satisfies 6 and the functions $h_{r}(x)$ and $h_{r1;i}(x)$ are bounded (in order) by $C_{r}\|x\|_{\infty}^{r}$. This latter condition is satisfied, for example, when $\mu$ is a finite Gaussian mixture in which each component has bounded means and covariance matrices with bounded eigenvalues.
Remark 8 (The importance of random features).
We end this section by highlighting the importance played by the randomness of $X_{1},\ldots,X_{p}\sim\mu$ in our argument. It is enough to consider $m_{r}$ for some $r\in\mathbb{N}$. We did show above that $m_{r}({\bm{\omega}}^{*})\approx\bar{m}_{r}({\bm{\omega}}^{*})$, for $h_{r}$ defined by (38), by using a law of large numbers-type argument. It is natural to ask if we could use a different construction that would, instead, give exact equality. Specifically, we ask the following question: Given generic, non-random $x_{1},\ldots,x_{p}$, does there exist a function $s_{r}:\mathbb{R}^{L}\to\mathbb{R}$ such that $m_{r}=\bar{m}_{r}$? We show in Section C.5 that, unfortunately, no such function can exist, even for $r=1$.
3.3 Sample-level estimation of softmax mixtures with random features
We state the MoM based estimator of the mixture parameters. Its rate of convergence is derived in Theorem 5 below, and is shown to satisfy the warm start requirement under which the EM estimator converges to ${\bm{\omega}}^{*}$ at near-parametric rates.
Let $Y_{1},\ldots,Y_{N}$ be i.i.d. from $\pi(\cdot;{\bm{\omega}}^{*}|x)$. Given functions $h_{r}$ and $h_{1r;i}$ defined by (38) and (39), it is natural to estimate $\bar{m}_{r}$ and $\bar{m}_{r1;i}$ by
$\widehat{m}_{r}:=\frac{1}{N}\sum_{\ell=1}^{N}h_{r}(Y_{\ell}),\quad\mbox{and}\quad\widehat{m}_{r1;i}:=\frac{1}{N}\sum_{\ell=1}^{N}h_{r1;i}(Y_{\ell}).$ (49)
By forming the vectors
$\widehat{\bm{m}}=\left(\widehat{m}_{1},\ldots,\widehat{m}_{2K-1}\right)^{\top}\quad\text{and}\quad\widehat{\bm{m}}_{1;i}:=\left(\widehat{m}_{01;i},\ldots,\widehat{m}_{(K-1)1;i}\right)^{\top},$ (50)
for $i\in\{2,\ldots,L\}$, the sample level MoM estimator $\widehat{\bm{\omega}}=(\widehat{\bm{\alpha}},\widehat{\bm{\theta}}_{1},\ldots,\widehat{\bm{\theta}}_{K})$ is given by Algorithm 1 with $\bar{\bm{m}}$ and $\bar{\bm{m}}_{1;i}$ replaced by $\widehat{\bm{m}}$ and $\widehat{\bm{m}}_{1;i}$, respectively.
The following theorem gives the rate of convergence of $d(\widehat{\bm{\omega}},{\bm{\omega}}^{*})$ for the two cases discussed in Theorem 3 and Theorem 4. For both cases, Theorem 5 shows that the sample level MoM estimator $\widehat{\bm{\omega}}$ is also an excellent warm start candidate for the EM algorithm in Section 2.2: it trivially meets the initialization requirement of the EM in (20) for any $p$ that satisfies $p\geq(L\log p)^{2}$ and $N$ satisfying (25).
Theorem 5.
Under the conditions of Theorem 3, we have almost surely,
$d(\widehat{\bm{\omega}},{\bm{\omega}}^{*})=\mathcal{O}_{\mathbb{P}}\left(\sqrt{L\log(L)/N}+\epsilon_{p}\sqrt{L}\right)$ (51)
with $\epsilon_{p}=\sqrt{\log\log p/p}$ for all but finitely many $p$. Under the conditions of Theorem 4, (51) holds, with $\epsilon_{p}=\sqrt{\log p/p}$ and with probability at least $1-p^{-s}$, for sufficiently large $p\geq p(B,\overline{\sigma},s)$ and any $s>1$.
Proof.
The proof is given in Section C.8. ∎
3.4 Subspace estimation via MoM under softmax mixtures
Since in practice the feature dimension $L$ could be (much) larger than the number of mixture components $K$, we focus on the case $L\geq K$ in this section and show that the MoM procedure in previous sections can be adapted to estimate the subspace of $\mathbb{R}^{L}$ spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, which hasc dimension at most $K$. As important applications, the estimated subspace can be used in two ways: (1) to select a basis in which the primary axis condition in 4 holds (see Section 3.4.1); and (2) to reduce the number of required random initializations for the EM algorithm, when such initializations are employed (see Section 3.4.2).
It suffices to consider estimating the $K$-dimensional subspace spanned by the columns of the following $L\times L$ matrix
$\Gamma:=\sum_{k=1}^{K}\alpha_{k}^{*}{\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}.$ (52)
Recall Proposition 1 and the choice of $h_{r}$ in (38). For the choice (with $r=2$)
$h_{2}(x,\mathbf{e}_{1})={1\over\mu(x)}~\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mu(x+t\mathbf{e}_{1})\Big{|}_{t=0}={1\over\mu(x)}\mathbf{e}_{1}^{\top}\nabla^{2}\mu(x)\mathbf{e}_{1}$
with $\nabla^{2}\mu(x)$ being the Hessian matrix of $\mu$ at $x$, we have, for any generic ${\bm{\theta}}\in\mathbb{R}^{L}$,
${\mathbb{E}_{\mu}\left[h_{2}(X,\mathbf{e}_{1})\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]}=\mathbf{e}_{1}^{\top}{\bm{\theta}}{\bm{\theta}}^{\top}\mathbf{e}_{1}.$
As the above holds for all $\mathbf{e}_{1},\ldots,\mathbf{e}_{L}$ and for any ${\bm{\theta}}$, it suggests to consider the moment matrix
$\bar{\Gamma}:=\mathbb{E}_{\pi({\bm{\omega}}^{*}|x)}\left[(\mu(Y))^{-1}\nabla^{2}\mu(Y)\right]=\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}(\mu(X_{j}))^{-1}\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})~\nabla^{2}\mu(X_{j})\over{1\over p}\sum_{j=1}^{p}\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})}$ (53)
and its population version
$\sum_{k=1}^{K}\alpha_{k}^{*}{\mathbb{E}_{\mu}\left[(\mu(X))^{-1}\exp(X^{\top}{\bm{\theta}}_{k}^{*})~\nabla^{2}\mu(X)\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}}_{k}^{*})\right]}=\sum_{k=1}^{K}\alpha_{k}^{*}{\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}$
which equals $\Gamma$. Therefore, the sample analogue of $\bar{\Gamma}$
$\widehat{\Gamma}:={1\over N}\sum_{i=1}^{N}{1\over\mu(Y_{i})}\nabla^{2}\mu(Y_{i})$ (54)
should estimate $\Gamma$ well, so that its first $K$ eigenvectors can be used to estimate the span of ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. The following proposition provides the justification and its proof, stated in Section C.9, reasons similarly as in Propositions 1 and 5.
Proposition 3.
Grant Assumptions 2 and 3. Assume
$\|(\mu(x))^{-1}\nabla^{2}\mu(x)\|_{\mathrm{op}}\leq C\|x\|_{2}^{2},\qquad\forall~x\in\mathbb{R}^{L}$ (55)
for some constant $C>0$. Then for any converging sequence $\epsilon_{p}^{\prime}$, on the event $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime}):=\left\{\|\bar{\Gamma}-\Gamma\|_{\mathrm{op}}\leq\epsilon_{p}^{\prime}\right\},$ for sufficiently large $p$, with probability at least $1-N^{-1}$, one has
$\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}~\leq~C^{\prime}\|\mathbf{X}\|_{\infty,2}\sqrt{\log N\over N}+\epsilon_{p}^{\prime}.$
As mentioned in Remark 7, condition (55) is a mild Lipschitz requirement, which holds, for instance, for $\mu$ being Gaussian, or for finite Gaussian mixtures.
In the setting of Theorem 3 for fixed $\mu$, it is easy to see that $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})$ holds $\mu$ almost surely, with $\epsilon_{p}^{\prime}\leq\epsilon_{p}\sqrt{L}=\mathcal{O}(\sqrt{\log\log p/p})$ for sufficiently large $p$.
For $\mu$ allowed to depend on $p$, we give explicit results for Gaussian below. Similar results hold for sub-Gaussian $\mu$ under conditions mentioned in Remark 7.
Example 3.
For $\mu=\mathcal{N}_{L}(0,\Sigma)$, we prove in Section C.10 that the event $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})$ holds for $\epsilon_{p}^{\prime}=\mathcal{O}(\sqrt{L\log(p)/p})$ with probability at least $1-p^{-1}$, provided that $p\geq L^{a}$ for some $a>3$. On the other hand, using (2), the choice of $\widehat{\Gamma}$ in (54) becomes
$\widehat{\Gamma}={1\over N}\sum_{\ell=1}^{N}\Sigma^{-1}Y_{\ell}Y_{\ell}^{\top}\Sigma^{-1}-\Sigma^{-1}.$ (56)
As a result, with probability at least $1-N^{-1}-p^{-1}$, one has
$\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}~\lesssim~\lambda_{L}^{-1}(\Sigma)\sqrt{(L+\log(p))\log(N)\over N}+\lambda_{L}^{-1}(\Sigma)\sqrt{L\log(p)\over p}.$ (57)
When $\Sigma$ is unknown, it can be consistently estimated by the sample $\widehat{\Sigma}=p^{-1}\mathbf{X}^{\top}\mathbf{X}$.
3.4.1 Application to practical choice of the primary axis
Recall in Section 3.1 that we have chosen the primary axis as $\mathbf{e}_{1}$, which leads to $\Delta_{1}$ in 4. As mentioned in Remark 10, finding a good primary axis $v$ relative to which $\Delta_{1}$ is large is crucial for the success of the MoM estimation technique. In rare cases, the statistician may have a priori knowledge of a good direction $v$ for which a lower bound on $\Delta(v^{\top}{\bm{\theta}}_{1}^{*},\dots,v^{\top}{\bm{\theta}}_{K}^{*})$, defined in (28), is sufficiently large. In general, to obtain results that hold uniformly over the parameter space, one could choose $v$ randomly on the sphere $\mathbb{S}^{L-1}$. Let
$\Delta^{2}:=\min_{k\neq k^{\prime}}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2},$ (58)
then a simple probabilistic argument (see, Lemma 32) gives that, for any $t\in(0,1)$ and any $v$ uniformly drawn from $\mathbb{S}^{L-1}$,
$\mathbb{P}\left\{\Delta(v^{\top}{\bm{\theta}}_{1}^{*},\dots,v^{\top}{\bm{\theta}}_{K}^{*})\geq{t\Delta\over K^{2}\sqrt{L}}\right\}\geq 1-t.$ (59)
Fix such $v$, and let $R=(v,w_{2},\ldots,w_{L})$ be an $L\times L$ rotation matrix with orthonormal columns. Following an identical argument to that of the proof of Proposition 2, applied to the re-scaled targets $R^{\top}{\bm{\theta}}_{1}^{*},\ldots,R^{\top}{\bm{\theta}}_{K}^{*}$, the constant $D$ scales as multiples of $(\Delta/(K^{2}\sqrt{L}))^{-cK}$, for some $c>0$. Thus, it scales as $K^{\mathcal{O}(K)}$ if $L<K$, while it scales as $L^{\mathcal{O}(K)}$ otherwise. Fortunately, it is possible to eliminate the dependency of $D$ on $L$ altogether, by choosing a direction $v$ from the lower-dimensional subspace spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, which will allow us to improve upon (59). Recall that the subspace of ${\bm{\theta}}_{k}^{*}$’s are contained in $\Gamma$ given by (52). Denote by $\widehat{V}\in\mathbb{O}_{L\times K}$ the first $K$ eigenvectors of its estimator $\widehat{\Gamma}$ in (54). In view of Proposition 3, we propose to choose the projection vector $v$ as
$v={\widehat{V}\widehat{V}^{\top}u\over\|\widehat{V}\widehat{V}^{\top}u\|_{2}}$ (60)
where the vector $u\in\mathbb{R}^{L}$ contains i.i.d. entries of $\mathcal{N}(0,1)$. The following lemma gives a lower bound on the desired minimum pairwise separation relative to this choice for $v$. It is worth mentioning that our analysis does not require any spectral condition on $\Gamma$.
Lemma 2.
Grant Assumptions 2, 3, 5 and condition (55). Then for any $t\in(0,1)$ and any $v$ drawn as (60), on the event $C^{\prime}\|\mathbf{X}\|_{\infty,2}\sqrt{\log N/N}+\epsilon_{p}^{\prime}\leq\underline{\alpha}\Delta^{2},$ one has
$\mathbb{P}\left\{\Delta(v^{\top}{\bm{\theta}}_{1}^{*},\ldots,v^{\top}{\bm{\theta}}_{K}^{*})\geq{t\Delta\over 2K^{2}\sqrt{K}}\right\}\geq 1-t.$
Proof.
The proof is given in Section C.11. ∎
Compared to (59), the dimension reduction in (60) eliminates the dependency on $L$ in the constant $D$ of Proposition 2.
Remark 9 (A practical heuristic).
In practice we recommend to take multiple random projection vectors $\{v_{1},\ldots,v_{n}\}$, and select the one that yields the largest separation. However since the separation $\Delta(v_{i}^{\top}{\bm{\theta}}_{1}^{*},\ldots,v_{i}^{\top}{\bm{\theta}}_{K}^{*})$ is unknown, we propose to use the following criterion. For any $i\in[n]$, compute the moment vector $\widehat{\bm{m}}(v_{i})$ as in (49) with $h_{r}$ given by (124), for each $v_{i}$, and its denoised version $\widetilde{\bm{m}}(v_{i})$ as in (43), form the moment matrix of $\widetilde{\bm{m}}(v_{i})$ as
$\widetilde{\mathbf{M}}(v_{i}):=\begin{pmatrix}1&\widetilde{m}_{1}(v_{i})&\dots&\widetilde{m}_{K-1}(v_{i})\\ \widetilde{m}_{1}(v_{i})&\widetilde{m}_{2}(v_{i})&\dots&\widetilde{m}_{K}(v_{i})\\ \vdots&\vdots&&\vdots\\ \widetilde{m}_{K-1}(v_{i})&\widetilde{m}_{K}(v_{i})&\dots&\widetilde{m}_{2K-2}(v_{i})\end{pmatrix},$
and choose $v_{i^{*}}$ with $i_{*}$ selected as
$i_{*}=\operatorname*{argmax}_{i\in[n]}~\det\left(\widetilde{\mathbf{M}}(v_{i})\right).$
The intuition lies in the important result in (Lindsay, 1989, Theorem A2) that
$\det({\mathbf{M}}(v_{i}))={1\over K!}\prod_{1\leq k<k^{\prime}\leq K}\left(v_{i}^{\top}{\bm{\theta}}_{k}^{*}-v_{i}^{\top}{\bm{\theta}}_{k^{\prime}}^{*}\right)^{2}$
so that the selected $v_{i^{*}}$ approximately maximizes $\det({\mathbf{M}}(v_{i}))$, thereby leading to the largest separation among $v_{i}^{\top}{\bm{\theta}}_{1}^{*},\ldots,v_{i}^{\top}{\bm{\theta}}_{K}^{*}.$
3.4.2 Application to random initialization of the EM
In Theorem 2 of Section 2.2 we show that the EM algorithm has provable guarantees once its initialization meets (20). In addition to using the MoM estimator developed in Section 3.3, it is common in practice to simply deploy random initializations, that is, by simply drawing ${\bm{\theta}}_{1}^{(0)},\ldots,{\bm{\theta}}_{K}^{(0)}$ uniformly from a chosen sphere multiple times, and selecting the corresponding EM estimate that yields the highest likelihood.
The intuition is the following: Let ${\bm{\theta}}^{*}\in\mathbb{S}^{L-1}$ be a given target vector and let $\delta>0$ be the desired accuracy. Then for any $\varepsilon>0$, with probability at least $1-\epsilon$, there exists at least one vector $\bar{v}$ in independent draws $\{v_{1},\ldots,v_{m}\}$ from $\mathbb{S}^{L-1}$ such that
$\|\bar{v}-{\bm{\theta}}^{*}\|_{2}\leq\delta$ (61)
provided that
$m\geq\exp\left(L(1-{\delta^{2}/2})\right)\log(1/\varepsilon).$
The above result follows from a simple union bound argument together with the spherical cap probability bound in Tkocz (2012). For completeness, we include its proof in Section C.12. As a result, in the worst case one needs to use $\exp(\mathcal{O}(L))$ random initializations and run the EM algorithm this many times, which is computationally expensive when $L$ is not small.
However, if ${\bm{\theta}}^{*}$ is known to lie within a subspace of dimension at most $K\ll L$, then one only needs $\exp(\mathcal{O}(K))$ random draws from the unit sphere in this subspace to achieve the desired $\delta_{0}$ accuracy. We formalize this in the following lemma in our context. Recall that ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ lie in the column space of $\Gamma$ given in (52). Further recall that $\widehat{V}\in\mathbb{O}_{L\times K}$ contains the first $K$ leading eigenvectors of $\widehat{\Gamma}$, the estimator of $\Gamma$ given in (54).
Lemma 3.
Fix any $k\in[K]$ and ${\bm{\theta}}_{k}^{*}\in\mathbb{S}^{L-1}$. Let $v_{1},\ldots,v_{m}$ be independently sampled as (60). For arbitrary $\varepsilon>0$ and $\delta_{0}>0$, on the event $\{\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}\leq(\underline{\alpha}/2)\delta_{\Gamma}\}$ for some $\delta_{\Gamma}<\delta_{0}^{2}/2$, with probability at least $1-\varepsilon$, there exists at least one vector $\bar{v}\in\{v_{1},\ldots,v_{m}\}$ such that $\|\bar{v}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}$ provided that
$m\geq\exp\left\{K(1-\delta_{0}^{2}/2+\delta_{\Gamma})\right\}\log(1/\varepsilon).$
Proof.
The proof is stated in Section C.12. ∎
By plugging into the bound of $\delta_{0}$ in (20) as well as the bound of $\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}$ in Proposition 3, the requirement $\delta_{\Gamma}<\delta_{0}^{2}/2$ becomes $\|\mathbf{X}\|_{\infty,2}^{3}\sqrt{\log N/N}+\|\mathbf{X}\|_{\infty,2}^{2}\epsilon_{p}^{\prime}\leq c(\underline{\alpha},\varsigma,\overline{\sigma})$ where we further recall that $\epsilon_{p}^{\prime}=\mathcal{O}(\sqrt{\log\log p/p})$ and $\epsilon_{p}^{\prime}=\mathcal{O}(\sqrt{L\log p/p})$ in the settings of Theorem 3 and 4, respectively.
4 Simulations
In this section we conduct numerical experiments to corroborate our theoretical findings in Sections 3 and 2. In Section 4.1 we first examine how the performance of the EM and MoM estimators depends on $N$, $p$, $L$ and $K$. In Section 4.2 we demonstrate the benefit of using the dimension reduction technique from Section 3.4.2 to initialize the EM algorithm.
To generate the data, we first generate $X_{1},\ldots,X_{p}$ i.i.d. from $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. The mixing weights are set to ${\bm{\alpha}}^{*}=(1/K)1_{K}$ for any given integer $K$. To generate the parameters ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, we first draw a $L\times K$ matrix with entries i.i.d. from $\mathcal{N}(0,1)$. We then set its $K$ left-singular vectors as ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. Finally, we resample $Y_{1},\ldots,Y_{N}$ according to model (2).
We consider the following estimation methods:
  • (1)
    MoM, the Method of Moments estimator in Section 3.3 with $n=200$ for choosing the projection direction as discussed in Remark 9;
  • (2)
    EM-MoM, the EM estimator in Section 2 that uses the MoM estimator as initialization;
  • (3)
[…]
  • (4)
    EM-oracle, the EM estimator that uses the true parameter values as the initialization.
For the EM algorithm in Section 2, we choose the step size $\eta_{k}=0.2$ and use the stopping rule that the relative change of the log-likelihood is smaller than $10^{-6}$.
To evaluate each method, for generic estimators $\widehat{\bm{\alpha}}$ and $\widehat{\bm{\theta}}_{1},\ldots,\widehat{\bm{\theta}}_{K}$, we choose
$\text{Err}_{\bm{\theta}}=\left({1\over{K}}\sum_{k=1}^{K}\bigl{\|}{\bm{\theta}}_{k}^{*}-\widehat{\bm{\theta}}_{\varrho(k)}\bigr{\|}_{2}^{2}\right)^{1/2},\quad\text{Err}_{\bm{\alpha}}=\sum_{k=1}^{K}\bigl{|}\alpha_{k}^{*}-\widehat{\alpha}_{\varrho(k)}\bigr{|}$
where $\varrho:[K]\to[K]$ is the best permutation that minimizes $\text{Err}_{{\bm{\theta}}}$.
4.1 Dependence of estimation error on $N,p,L$ and $K$
We vary $N\in\{2,4,6,8,10\}\times 10^{3}$, $p\in\{1,3,5,7,10\}\times 10^{3}$, $L\in\{20,40,60,80,100\}$ and $K\in\{2,4,6,8,10\}$ one at a time to examine their effects on the estimation errors. The baseline setting uses $L=50$ and $K=3$ when these parameters are not varied. When $N$ is varied, we set $p=5000$ and when $p$ is varied, we set $N=7000$. When varying either $L$ or $K$, we chose $N=10000$ and $p=7000$. For each setting, we report the averaged errors over 200200200 repetitions in Fig. 1 for $\text{Err}_{\bm{\theta}}$ (and in Fig. 3 of Section A.1 for $\text{Err}_{\bm{\alpha}}$).
Regarding $\text{Err}_{\bm{\theta}}$, all methods perform better as $N$ increases and $L$ or $K$ decreases. For EM-oracle, since it has no algorithmic error, our Theorem 2 shows that its estimation error is purely the statistical error which is of order $\sqrt{L\log(N)/N}$. The MoM estimator is outperformed by the EM estimators in all settings. Once $N\geq p$, further increasing $N$ does not improve the performance of MoM. When $p$ increases, the performance of MoM improves, whereas the EM estimators remain unchanged. We also note that the figures in which we vary $N$ and $p$ suggest the rate for MoM is slower than the parametric rate, confirming the observation made in Remark 11 above. In Section A.2, we conduct a separate simulation study below to verify that MoM can indeed enjoy a parametric rate.
Overall, for $K=3$, EM-MoM and EM-dr-rand-10 have overall comparable performance, with the former performing slightly better for large $N$. One drawback of EM-dr-rand-10 is its higher computational cost due to sampling multiple initializations and evaluating their likelihoods (the computational complexity scales linearly with the number of initializations multiplied by the ambient dimension $p$).
As $K$ increases, the performance of all methods deteriorates, with MoM and EM-MoM degrading more rapidly than the others. For $K=10$, MoM (so does EM-MoM) fails to recover all $K$ mixture components, as the root-finding step in Algorithm 1 fails in this case.
These findings are all aligned with our theory in Sections 3 and 2.
Figure 1: The averaged $\text{Err}_{{\bm{\theta}}}$ in different settings
4.2 Benefits of multiple random initializations with dimension reduction
[…]
Figure 2: The averaged $\text{Err}_{{\bm{\theta}}}$ in different settings
Appendix A Additional simulations
A.1 Results of estimating $\alpha^{*}$ in the setting of Section 4.1
Fig. 3 shows the errors $\text{Err}_{\bm{\alpha}}$ of all methods in Section 4.1. For estimating the mixing weights ${\bm{\alpha}}^{*}$, both MoM and EM-MoM exhibit greater fluctuations in their errors due to the method’s sensitivity to the choice of random projection in Section 3.4.1. For large $K$, the errors in estimating ${\bm{\alpha}}^{*}$ are substantially larger for these methods compared to other EM estimators, and are more sensitive than the corresponding errors in estimating ${\bm{\theta}}_{k}^{*}$. EM-dr-rand-10 and EM-oracle perform better for larger $N$ and smaller $K$, whereas their performance remains similar as $p$ and $L$ vary.
Figure 3: The averaged $\text{Err}_{{\bm{\alpha}}}$ in different settings
A.2 Parametric rate of MoM
We conduct a separate simulation study below to verify that MoM can indeed enjoy a parametric rate, by taking $K=2$ and ensuring that the atoms have the theoretically prescribed separation. We let $L=50$ and vary $p=N\in\{1,3,5,7,9,12,15\}\times 10^{3}$. Fig. 4 depicts the estimation errors of MoM, EM-MoM and EM-oracle. We observe the same phenomenon as above except that $\text{Err}_{{\bm{\theta}}}$ of MoM decays in the faster parametric rate as $N$ and $p$ increase. The large fluctuation of $\text{Err}_{{\bm{\alpha}}}$ for MoM can be explained by the sensitivity of the method to the selection of the random projection in Section 3.4.1.
Figure 4: $K=2$:  The averaged $\text{Err}_{{\bm{\theta}}}$ (left) and $\text{Err}_{{\bm{\alpha}}}$ (right) in different settings
Appendix B Proofs of Section 2: theory of the hybrid EM algorithm for generic features
Notation. For any ${\bm{\omega}}^{*}=({\bm{\alpha}}^{*},{\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*})$, recall that $\Delta^{2}=\min_{k\neq k^{\prime}}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2}.$ For any ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$, we write for each $j\in[p]$,
$A_{{\bm{\theta}}_{k}}(x_{j})=A(x_{j};{\bm{\theta}}_{k}),\qquad\pi_{{\bm{\omega}}}(x_{j})=\pi(x_{j};{\bm{\omega}})=\sum_{k=1}^{K}\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j}).$
For any ${\bm{\theta}}\in\mathbb{R}^{L}$ with $A_{\bm{\theta}}=(A_{\bm{\theta}}(x_{1}),\ldots,A_{{\bm{\theta}}}(x_{p}))^{\top}\in\Delta^{p}$, write
$\Sigma_{A_{\bm{\theta}}}:=\mathrm{diag}(A_{\bm{\theta}})-A_{\bm{\theta}}A_{\bm{\theta}}^{\top}.$
Further let
$\begin{split}&N_{\bm{\theta}}=\sum_{j=1}^{p}e^{x_{j}^{\top}{\bm{\theta}}}\in\mathbb{R},\quad\mathrm{I}_{\bm{\theta}}=\sum_{j=1}^{p}e^{x_{j}^{\top}{\bm{\theta}}}x_{j}x_{j}^{\top}\in\mathbb{R}^{L\times L},\quad\mathrm{II}_{\bm{\theta}}=\sum_{j=1}^{p}e^{x_{j}^{\top}{\bm{\theta}}}x_{j}\in\mathbb{R}^{L}\end{split}$ (62)
and note that
$H_{\bm{\theta}}:=\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}}}\mathbf{X}={\mathrm{I}_{\bm{\theta}}\over N_{\bm{\theta}}}-{\mathrm{II}_{\bm{\theta}}\mathrm{II}_{\bm{\theta}}^{\top}\over N_{\bm{\theta}}^{2}}.$ (63)
B.1 Key lemmas for the proof of Theorem 2
The following are non-trivial results that establish strong concavity and local smoothness of the function $q_{k}({\bm{\omega}})=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$, smoothness of ${\bm{\omega}}^{\prime}\mapsto\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$, Lipschitz continuity of $M_{k}({\bm{\omega}})$ and maximal inequalities for $|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|$ and $\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}$, uniformly over ${\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$. The proofs are rather involved and can be found in separate sections below.
Lemma 4.
Under 1 and (20), for all $k\in[K]$, we set
$\gamma_{k}={(1-c_{0})}\alpha_{k}^{*}\underline{\sigma}^{2},\qquad\mu_{k}={(1+c_{0})}\alpha_{k}^{*}\overline{\sigma}^{2}$ (64)
with $c_{0}$ specified in (20). Then for any ${\bm{\omega}},{\bm{\omega}}^{\prime}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ and $k\in[K]$,
$({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}\bigl{(}q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\bigr{)}$ $\leq-\gamma_{k}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}^{2}$ (65)
$\|q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\|_{2}$ $\leq\mu_{k}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}$ (66)
and
$({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}\left(q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\right)\leq-{\mu_{k}\gamma_{k}\over\mu_{k}+\gamma_{k}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}^{2}-{1\over\mu_{k}+\gamma_{k}}\|q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\|_{2}^{2}.$ (67)
Proof.
See Section B.2.2. ∎
Lemma 5.
Under 1 and (20), we have, for any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ ,
$\max_{k\in[K]}|M_{k}({\bm{\omega}})-M_{k}({\bm{\omega}}^{*})|$ $\leq\kappa~d({\bm{\omega}},{\bm{\omega}}^{*})$ (68)
$\max_{k\in[K]}\|\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})\|_{2}$ $\leq\overline{\sigma}~\kappa~d({\bm{\omega}},{\bm{\omega}}^{*})$ (69)
where for some large absolute constant $C>0$,
$\kappa=CK\overline{\alpha}(1+\overline{\sigma}^{2}\Delta^{2})\exp(-\underline{\sigma}^{2}\Delta^{2}/8).$ (70)
Proof.
See Section B.2.3. ∎
Lemma 6.
Grant 1 and conditions (20) & (25). Set
$\epsilon_{N}=\sqrt{\overline{\alpha}KL\log(N)\over N}.$
There exists some absolute constant $C>0$ such that with probability at least $1-\mathcal{O}(N^{-L})$,
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})\right|$ $\leq C~\epsilon_{N},$
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}\left\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\right\|_{2}$ $\leq C~\overline{\sigma}~\epsilon_{N}.$
Proof.
See Section B.2.4. ∎
B.2 Proofs of Lemmas 4, 5 and 6
To prove Lemmas 4, 5 and 6, we first state and prove a few technical lemmas.
B.2.1 Technical lemmas used to prove Lemmas 4, 5 and 6
The following lemma proves certain Lipschitz continuity of the function $A_{\bm{\theta}}$ and $\pi_{\bm{\omega}}$ relative to changes in ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$.
Lemma 7.
For any ${\bm{\omega}},{\bm{\omega}}^{\prime}\in\Omega$ such that
$\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\leq{1\over 3}\min_{k\in[K]}\alpha_{k},\qquad\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}\leq{1\over 2\|\mathbf{X}\|_{\infty,2}},$ (71)
we have, for any $j\in[p]$ and $k\in[K]$,
${|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over A_{{\bm{\theta}}_{k}}(x_{j})\wedge A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})}\leq 3\|\mathbf{X}\|_{\infty,2}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2},$ (72)
${\left|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})\right|\over\pi_{{\bm{\omega}}}(x_{j})}\leq\max_{k\in[K]}{\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\over\alpha_{k}}+4\|\mathbf{X}\|_{\infty,2}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}.$ (73)
Proof.
We first prove (72). Pick any $k\in[K]$ and $j\in[p]$. By definition, we have
$\begin{split}A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})&={1\over\sum_{\ell=1}^{p}e^{(x_{\ell}-x_{j})^{\top}{\bm{\theta}}_{k}}}-{1\over\sum_{\ell=1}^{p}e^{(x_{\ell}-x_{j})^{\top}{\bm{\theta}}_{k}^{\prime}}}\\ &={1\over\sum_{\ell=1}^{p}e^{(x_{\ell}-x_{j})^{\top}{\bm{\theta}}_{k}}}{\sum_{\ell=1}^{p}e^{(x_{\ell}-x_{j})^{\top}{\bm{\theta}}_{k}^{\prime}}\left[1-e^{(x_{\ell}-x_{j})^{\top}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})}\right]\over\sum_{\ell=1}^{p}e^{(x_{\ell}-x_{j})^{\top}{\bm{\theta}}_{k}^{\prime}}}\\ &\leq A_{{\bm{\theta}}_{k}}(x_{j})~\max_{\ell\in[p]}\left|1-e^{(x_{\ell}-x_{j})^{\top}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})}\right|\end{split}$ (74)
This bound, the inequality
[…]
and the basic inequality $|1-e^{t}|\leq 3|t|$ for any $|t|\leq 1$ combined give that
$A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\leq 3A_{{\bm{\theta}}_{k}}(x_{j})~\|\mathbf{X}\|_{\infty,2}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}.$
The same arguments hold after we swap ${\bm{\theta}}_{k}$ and ${\bm{\theta}}_{k}^{\prime}$, and (72) follows.
To prove (73), we have
${|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}$
$\leq{1\over\pi_{{\bm{\omega}}}(x_{j})}\sum_{k=1}^{K}\left(|\alpha_{k}-\alpha_{k^{\prime}}|A_{{\bm{\theta}}_{k}}(x_{j})+\alpha_{k}^{\prime}\left|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\right|\right)$
$\leq\sum_{k=1}^{K}\left({\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\over\alpha_{k}}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha_{k}^{\prime}\over\alpha_{k}}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}3\|\mathbf{X}\|_{\infty,2}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}\right)$ by (72)
$\leq\max_{k\in[K]}\left({\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\over\alpha_{k}}+{3\alpha_{k}^{\prime}\over\alpha_{k}}\|\mathbf{X}\|_{\infty,2}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}\right)$
$\leq\max_{k\in[K]}{\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\over\alpha_{k}}+4\|\mathbf{X}\|_{\infty,2}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}$ by (71)
The proof is complete. ∎
The following lemma controls the eigenvalues of $H_{{\bm{\theta}}_{k}}$ defined in (63) for all ${\bm{\theta}}_{k}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*};\delta_{0}/\overline{\sigma})$ with $\delta_{0}$ satisfying (20) and $k\in[K]$.
Lemma 8.
Fix any ${\bm{\theta}}^{*}\in\mathbb{R}^{L}$ and any $\delta_{0}$ satisfying (20). Under (15) and (16), there exists some constant $c=c(c_{0})\in(0,1)$ such that for all $k\in[K]$,
$(1-c)\underline{\sigma}^{2}\leq\lambda_{L}(H_{{\bm{\theta}}_{k}})\leq\lambda_{1}(H_{{\bm{\theta}}_{k}})\leq(1+c)\overline{\sigma}^{2},\qquad\forall~{\bm{\theta}}_{k}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*};\delta_{0}/\overline{\sigma})$ (75)
Proof.
Fix ${\bm{\theta}}^{*}\in\{{\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{k}^{*}\}$. Write $H_{{\bm{\theta}}^{*}}^{1/2}$ as the matrix square root of $H_{{\bm{\theta}}^{*}}=\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}^{*}}}\mathbf{X}$. Let ${\bm{\theta}}\in{\mathbb{B}}({\bm{\theta}}^{*},\delta_{0}/\overline{\sigma})$ be arbitrary. We first bound from above
$\|H_{{\bm{\theta}}^{*}}^{-1/2}(H_{\bm{\theta}}-H_{{\bm{\theta}}^{*}})H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}$ $=\sup_{v\in\mathbb{S}^{L-1}}v^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}(H_{\bm{\theta}}-H_{{\bm{\theta}}^{*}})H_{{\bm{\theta}}^{*}}^{-1/2}v$
$=\sup_{v\in\mathbb{S}^{L-1}}\left(R_{1}(v)+R_{2}(v)\right)$
with
$R_{1}(v)$ $:=\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}\left(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j})\right),$
$R_{2}(v)$ $:=\left|\sum_{j=1}^{p}x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\left(A_{{\bm{\theta}}^{*}}(x_{j})+A_{{\bm{\theta}}}(x_{j})\right)\right|~\left|\sum_{j=1}^{p}x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\left(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j})\right)\right|.$
We observe that ${\bm{\theta}}\in{\mathbb{B}}({\bm{\theta}}^{*};\delta_{0}/\overline{\sigma})$ implies $\|{\bm{\theta}}^{*}-{\bm{\theta}}\|_{2}\leq\delta_{0}/\overline{\sigma}$. After we invoke (72) in Lemma 7 with $K=1$ and ${\bm{\theta}}^{*}$ and ${\bm{\theta}}$ in lieu of ${\bm{\theta}}_{k}$ and ${\bm{\theta}}_{k}^{\prime}$, respectively, we find
$\sup_{v\in\mathbb{S}^{L-1}}R_{1}(v)$ $\leq\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}A_{{\bm{\theta}}^{*}}(x_{j})~3(\delta_{0}/\overline{\sigma})\|\mathbf{X}\|_{\infty,2}$ (76)
$\leq{3c_{0}\over\varsigma^{2}}\|H_{{\bm{\theta}}^{*}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A_{{\bm{\theta}}^{*}})\mathbf{X}H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}$ by (20)
[…]
Next, we observe that
$R_{2}(v)$ $\leq~2\left|\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)A_{{\bm{\theta}}^{*}}(x_{j})\right|\left|\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j}))\right|$
$\qquad+\left|\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j}))\right|^{2}.$
By repeating the arguments in (76), we find that
$\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)(A_{{\bm{\theta}}^{*}}(x_{j})-A_{{\bm{\theta}}}(x_{j}))$
$\leq 3(\delta_{0}/\overline{\sigma})\|\mathbf{X}\|_{\infty,2}\sum_{j=1}^{p}\left|x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\right|A_{{\bm{\theta}}^{*}}(x_{j})$ (77)
$\leq{3c_{0}\over\varsigma^{2}}\Bigl{(}\sum_{j=1}^{p}(x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v)^{2}A_{{\bm{\theta}}^{*}}(x_{j})\Bigr{)}^{1/2}\Bigl{(}\sum_{j=1}^{p}A_{{\bm{\theta}}^{*}}(x_{j})\Bigr{)}^{1/2}$
$\leq{3c_{0}\over\varsigma}\sqrt{\|H_{{\bm{\theta}}^{*}}^{-1/2}\mathbf{X}^{\top}\mathrm{diag}(A_{{\bm{\theta}}^{*}})\mathbf{X}H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}}$
$\leq 3c_{0}$ (78)
so that
$\sup_{v\in\mathbb{S}^{L-1}}R_{2}(v)$ $\leq{6c_{0}\over\varsigma}\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}\left|x_{j}^{\top}H_{{\bm{\theta}}^{*}}^{-1/2}v\right|A_{{\bm{\theta}}^{*}}(x_{j})+9c_{0}^{2}\leq 3(2c_{0}+3c_{0}^{2}).$
Combination of the bounds for $R_{1}(v)$ and $R_{2}(v)$, uniformly over $v\in\mathbb{S}^{L-1}$ yields
$\|H_{{\bm{\theta}}^{*}}^{-1/2}(H_{\bm{\theta}}-H_{{\bm{\theta}}^{*}})H_{{\bm{\theta}}^{*}}^{-1/2}\|_{\mathrm{op}}\leq 9c_{0}(1+c_{0}).$
Together with Weyl’s inequality, the eigenvalues of $H_{{\bm{\theta}}^{*}}^{-1/2}H_{\bm{\theta}}H_{{\bm{\theta}}^{*}}^{-1/2}$ satisfy
$\left|1-\lambda_{\ell}(H_{{\bm{\theta}}^{*}}^{-1/2}H_{\bm{\theta}}H_{{\bm{\theta}}^{*}}^{-1/2})\right|\leq 9c_{0}(1+c_{0}),\qquad\forall~1\leq\ell\leq L.$
In particular,
$\left(1-9c_{0}-9c_{0}^{2}\right)\lambda_{L}(H_{{\bm{\theta}}^{*}})\leq\lambda_{L}(H_{\bm{\theta}})\leq\lambda_{1}(H_{\bm{\theta}})\leq\lambda_{1}(H_{{\bm{\theta}}^{*}})\left(1+9c_{0}+9c_{0}^{2}\right)$
which completes the proof. ∎
The following two lemmas are crucial to the proof of Lemma 5. For any $a,k\in[K]$, let $\bar{\bm{\theta}}_{ak}^{*}$ be the midpoint of ${\bm{\theta}}_{a}^{*}$ and ${\bm{\theta}}_{k}^{*}$
$\bar{\bm{\theta}}_{ak}^{*}:={1\over 2}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*}).$ (79)
Lemma 9.
For any $a,k\in[K]$ and any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (20), under conditions (15) and (16), we have
$\left\|\mathbf{X}^{\top}\Bigl{(}\mathrm{diag}(A_{\bar{\bm{\theta}}_{ak}^{*}})-A_{{\bm{\theta}}_{k}}A_{{\bm{\theta}}_{k}}^{\top}\Bigr{)}\mathbf{X}\right\|_{\mathrm{op}}\lesssim~\overline{\sigma}^{2}+\overline{\sigma}^{4}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}.$
Proof.
For simplicity, let us write $\bar{\bm{\theta}}^{*}=\bar{\bm{\theta}}_{ak}^{*}$. Fix any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$. Using the notation in (62), it suffices to analyze
${1\over\sum_{\ell=1}^{p}e^{x_{\ell}^{\top}\bar{\bm{\theta}}^{*}}}\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}e^{x_{j}^{\top}\bar{\bm{\theta}}^{*}}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}v$
$={1\over N_{\bar{\bm{\theta}}^{*}}}\sup_{v\in\mathbb{S}^{L-1}}\left\{\sum_{j=1}^{p}e^{x_{j}^{\top}\bar{\bm{\theta}}^{*}}(x_{j}^{\top}v)^{2}+\sum_{j=1}^{p}e^{x_{j}^{\top}\bar{\bm{\theta}}^{*}}(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})^{2}-2\sum_{j=1}^{p}e^{x_{j}^{\top}\bar{\bm{\theta}}^{*}}x_{j}^{\top}v(v^{\top}\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})\right\}$
$=\sup_{v\in\mathbb{S}^{L-1}}\left\{v^{\top}H_{\bar{\bm{\theta}}^{*}}v+v^{\top}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{\bar{\bm{\theta}}^{*}})(A_{{\bm{\theta}}_{k}}-A_{\bar{\bm{\theta}}^{*}})^{\top}\mathbf{X}v\right\}$
$\leq\lambda_{1}(H_{\bar{\bm{\theta}}^{*}})+2\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{\bar{\bm{\theta}}^{*}})\|_{2}^{2}+2\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}^{*}})\|_{2}^{2}$
By repeating the arguments in the proof of (B.2.2), we obtain
$\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{\bar{\bm{\theta}}^{*}})\|_{2}$ $~\leq~\|{\bm{\theta}}_{k}^{*}-\bar{\bm{\theta}}^{*}\|_{2}\sup_{u\in[0,1]}\lambda_{1}(H_{\bar{\bm{\theta}}_{u}^{*}})$
$\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}^{*}})\|_{2}$ $~\leq~\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\sup_{u\in[0,1]}\lambda_{1}(H_{{\bm{\theta}}_{k,u}})$
where we write $\bar{\bm{\theta}}_{u}^{*}=u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{k}^{*}$ and ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}^{*}+(1-u){\bm{\theta}}_{k}$. The proof is completed by invoking (15) and (75). ∎
Recall that ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}+(1-u){\bm{\theta}}_{k}^{*}$ for any $u\in[0,1]$ and $k\in[K]$.
Lemma 10.
Let ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (20). Under condition (15), the following holds for any $a,k\in[K]$ with $a\neq k$,
$\max_{j\in[p]}\sup_{u\in[0,1]}{\alpha_{a}^{*}\alpha_{k}^{*}\over A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}{A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k,u}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\lesssim(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right).$
Proof.
Recall that ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ ensures $\|{\bm{\theta}}_{k,u}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma}$ and $\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\leq\delta_{0}\underline{\alpha}$. Under condition (20), after invoking Lemma 7 twice, we obtain
[…]
Since $A_{{\bm{\theta}}_{k}^{*}}(x_{j})=e^{x_{j}^{\top}{\bm{\theta}}_{k}^{*}}/N_{{\bm{\theta}}_{k}^{*}}$, we further obtain
$\alpha_{a}^{*}\alpha_{k}^{*}{A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}^{*}}(x_{j})}$ $={\alpha_{a}^{*}\alpha_{k}^{*}e^{x_{j}^{\top}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*})}/(N_{{\bm{\theta}}_{k}^{*}}N_{{\bm{\theta}}_{a}^{*}})\over\alpha_{a}^{*}e^{x_{j}^{\top}{\bm{\theta}}_{a}^{*}}/N_{{\bm{\theta}}_{a}^{*}}+\alpha_{k}^{*}e^{x_{j}^{\top}{\bm{\theta}}_{k}^{*}}/N_{{\bm{\theta}}_{k}^{*}}+\sum_{b\neq a,k}\alpha_{b}^{*}e^{x_{j}^{\top}{\bm{\theta}}_{b}^{*}}/N_{{\bm{\theta}}_{b}^{*}}}$
$\leq{\alpha_{a}^{*}\alpha_{k}^{*}e^{x_{j}^{\top}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*})}/(N_{{\bm{\theta}}_{k}^{*}}N_{{\bm{\theta}}_{a}^{*}})\over\alpha_{a}^{*}e^{x_{j}^{\top}{\bm{\theta}}_{a}^{*}}/N_{{\bm{\theta}}_{a}^{*}}+\alpha_{k}^{*}e^{x_{j}^{\top}{\bm{\theta}}_{k}^{*}}/N_{{\bm{\theta}}_{k}^{*}}}$
$\leq{(\alpha_{a}^{*}\vee\alpha_{k}^{*})~\exp\left(x_{j}^{\top}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*})/2\right)\over N_{{\bm{\theta}}_{k}^{*}}\exp\left(x_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)+N_{{\bm{\theta}}_{a}^{*}}\exp\left(-x_{j}^{\top}({\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*})/2\right)}$
$\leq(\alpha_{a}^{*}\vee\alpha_{k}^{*}){\exp(x_{j}^{\top}\bar{\bm{\theta}}_{ak}^{*})\over 2\sqrt{N_{{\bm{\theta}}_{k}^{*}}N_{{\bm{\theta}}_{a}^{*}}}}$
so that
${\alpha_{a}^{*}\alpha_{k}^{*}\over A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}{A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k,u}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\leq{\alpha_{a}^{*}\vee\alpha_{k}^{*}\over 2}{N_{\bar{\bm{\theta}}_{ak}^{*}}\over\sqrt{N_{{\bm{\theta}}_{k}^{*}}N_{{\bm{\theta}}_{a}^{*}}}}.$ (80)
It remains to bound from above
$\log N_{\bar{\bm{\theta}}_{ak}^{*}}-{1\over 2}\left(\log N_{{\bm{\theta}}_{k}^{*}}+\log N_{{\bm{\theta}}_{a}^{*}}\right).$
By letting $g({\bm{\theta}})=\log N_{{\bm{\theta}}}=\log(\sum_{j=1}^{p}e^{x_{j}^{\top}{\bm{\theta}}})$ for any ${\bm{\theta}}\in\mathbb{R}^{L}$, if there exists some $\nu>0$ such that $g({\bm{\theta}})$ is strongly $\nu$-convex over $u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{k}^{*}$ for all $u\in[0,1]$, then
$\log N_{\bar{\bm{\theta}}_{ak}^{*}}-{1\over 2}\left(\log N_{{\bm{\theta}}_{k}^{*}}+\log N_{{\bm{\theta}}_{a}^{*}}\right)=g(\bar{\bm{\theta}}_{ak}^{*})-{1\over 2}\left[g({\bm{\theta}}_{a}^{*})+g({\bm{\theta}}_{k}^{*})\right]\leq-{\nu\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$
which yields the desired result. To verify the strongly $\nu$-convexity of $g({\bm{\theta}})$, taking the derivative with respect to ${\bm{\theta}}$ twice and interchanging the expectation with derivatives give
$\nabla^{2}g({\bm{\theta}})={\sum_{j=1}^{p}x_{j}x_{j}^{\top}e^{x_{j}^{\top}{\bm{\theta}}}\over\sum_{\ell=1}^{p}e^{x_{\ell}^{\top}{\bm{\theta}}}}-{(\sum_{j=1}^{p}x_{j}e^{x_{j}^{\top}{\bm{\theta}}})(\sum_{j=1}^{p}x_{j}e^{x_{j}^{\top}{\bm{\theta}}})^{\top}\over(\sum_{\ell=1}^{p}e^{x_{\ell}^{\top}{\bm{\theta}}})^{2}}=H_{{\bm{\theta}}}.$
By condition (15), we know that $\lambda_{L}(H_{{\bm{\theta}}})\geq\underline{\sigma}^{2}$ for all ${\bm{\theta}}=u{\bm{\theta}}_{a}^{*}+(1-u){\bm{\theta}}_{k}^{*}$. This implies the strong $\underline{\sigma}^{2}$-convexity hence completes the proof. ∎
B.2.2 Proof of Lemma 4: strong concavity and smoothness of the gradient function $q(\cdot)=\nabla_{\theta_{k}}Q(\cdot\mid\omega^{*})$
Proof.
First, the fact that the statement in (67) follows from (65) and (66) is a classical result on strongly-convex, Lipschitz functions, see, for instance, Nesterov (2013, Theorem 2.1.12). We prove (65) and (66) below.
Let ${\bm{\omega}},{\bm{\omega}}^{\prime}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ with $\delta_{0}$ satisfying (20). Pick any $k\in[K]$. From (14), we find
$q_{k}({\bm{\omega}})=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$ $=\sum_{j=1}^{p}{\alpha^{*}_{k}A_{{\bm{\theta}}_{k}^{*}}(x_{j})}(x_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}})=\alpha_{k}^{*}~\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}^{*}}-A_{{\bm{\theta}}_{k}})$
so that we obtain
$({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}\left(q_{k}({\bm{\omega}})-q_{k}({\bm{\omega}}^{\prime})\right)$
$=-\alpha_{k}^{*}\sum_{j=1}^{p}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}x_{j}\left(A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\right)$
$=-\alpha_{k}^{*}\sum_{j=1}^{p}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}x_{j}\int_{0}^{1}A_{{\bm{\theta}}_{k,u}}(x_{j})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k,u}})^{\top}\mathbf{X}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})\mathrm{d}u$
$=-\alpha_{k}^{*}\int_{0}^{1}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})^{\top}\left[\sum_{j=1}^{p}A_{{\bm{\theta}}_{k,u}}(x_{j})x_{j}x_{j}^{\top}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k,u}}A_{{\bm{\theta}}_{k,u}}^{\top}\mathbf{X}\right]({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})\mathrm{d}u$
$\leq-\alpha_{k}^{*}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}^{2}\inf_{u\in[0,1]}\lambda_{L}\left(\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}_{k,u}}}\mathbf{X}\right)$
$=-\alpha_{k}^{*}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}^{2}\inf_{u\in[0,1]}\lambda_{L}(H_{{\bm{\theta}}_{k,u}})$ (81)
The second equality uses an Taylor expansion of $A_{{\bm{\theta}}_{k}}(x_{j})$ around ${\bm{\theta}}_{k}^{\prime}$ and we use the notation ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}+(1-u){\bm{\theta}}_{k}^{\prime}$ for any $u\in[0,1]$. The last step uses (63).
Similarly, we have
$\|\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{*})\|_{2}$
$=\alpha_{k}^{*}\left\|\sum_{j=1}^{p}\int_{0}^{1}A_{{\bm{\theta}}_{k,u}}(x_{j})x_{j}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k,u}})^{\top}\mathbf{X}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})\mathrm{d}u\ \right\|_{2}$
$\leq\alpha_{k}^{*}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}\sup_{u\in[0,1]}\lambda_{1}(H_{{\bm{\theta}}_{k,u}}).$ (82)
In view of (B.2.2) and (B.2.2) and the fact that ${\bm{\omega}},{\bm{\omega}}^{\prime}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ implies ${\bm{\theta}}_{k},{\bm{\theta}}_{k}^{\prime}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0}/\overline{\sigma})$, the Euclidean ball around ${\bm{\theta}}_{k}^{*}$ with radius $\delta_{0}/\overline{\sigma}$, so that ${\bm{\theta}}_{k,u}\in{\mathbb{B}}({\bm{\theta}}_{k}^{*},\delta_{0}/\overline{\sigma})$ for any $u\in[0,1]$, (65) and (66) follow by invoking Lemma 8 with ${\bm{\theta}}^{*}={\bm{\theta}}_{k}^{*}$ and ${\bm{\theta}}_{k}={\bm{\theta}}_{k,u}$ for all $k\in[K]$. The proof is complete. ∎
B.2.3 Proof of Lemma 5: the gradient smoothness of the surrogate function $Q$ and the Lipschitz continuity of $M_{\alpha}$
Proof.
Let ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ be arbitrary. We first prove (69). From (14), we argue
$\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$
$=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(x_{j})\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}^{*}}(x_{j})}\right)\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})$
$=\sum_{j=1}^{p}{1\over\pi_{{\bm{\omega}}}(x_{j})}\left[\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\sum_{a\neq k}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})-\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\sum_{a\neq k}\alpha_{a}A_{{\bm{\theta}}_{a}}(x_{j})\right]\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}}).$
By adding and subtracting terms, it now suffices to bound from above
$T_{1}:=\sum_{a\neq k}\left\|\sum_{j=1}^{p}\left(\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})-\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\right){\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}$
$T_{2}:=\sum_{a\neq k}\left\|\sum_{j=1}^{p}(\alpha_{a}A_{{\bm{\theta}}_{a}}(x_{j})-\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})){\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}$
Bounding $T_{1}$.
We start with the inequality $T_{1}\leq T_{11}+T_{12}$ where
$T_{11}$ $=|\alpha_{k}-\alpha_{k}^{*}|\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}A_{{\bm{\theta}}_{k}}(x_{j})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}$
$T_{12}$ $=\alpha_{k}^{*}\sum_{a\neq k}\left\|\sum_{j=1}^{p}\left(A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{*}}(x_{j})\right){\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}$
For the term $T_{12}$, after a Taylor expansion of $A_{{\bm{\theta}}_{k}}(x_{j})$ around ${\bm{\theta}}_{k}^{*}$, we find that
$T_{12}$ $=\alpha_{k}^{*}\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\int_{0}^{1}A_{{\bm{\theta}}_{k,u}}(x_{j})({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*})^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k,u}})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\mathrm{d}u\right\|_{2}$
$\leq\sup_{u\in[0,1]}\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\alpha_{k}^{*}A_{{\bm{\theta}}_{k,u}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k,u}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}\right\|_{\mathrm{op}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}.$
Here, we recall ${\bm{\theta}}_{k,u}=u{\bm{\theta}}_{k}+(1-u){\bm{\theta}}_{k}^{*}$. Further we denote
$\rho_{j}:={\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\alpha_{k}^{*}A_{{\bm{\theta}}_{k,u}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})},\qquad\forall~j\in[p].$
We proceed to bound from above
$T_{121}:=\sup_{u\in[0,1]}\sum_{a\neq k}\left\|\sum_{j=1}^{p}\rho_{j}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}\right\|_{\mathrm{op}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}$
and
$T_{122}:=\sup_{u\in[0,1]}\sum_{a\neq k}\left\|\sum_{j=1}^{p}\rho_{j}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k,u}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k,u}})^{\top}\mathbf{X}\right\|_{\mathrm{op}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}.$
This is indeed sufficient as $T_{12}\leq\sqrt{T_{121}T_{122}}$ follows from the Cauchy-Schwarz inequality. For any $a\neq k$, define the midpoint between ${\bm{\theta}}_{a}^{*}$ and ${\bm{\theta}}_{k}^{*}$ as
$\bar{\bm{\theta}}_{ak}^{*}:={1\over 2}({\bm{\theta}}_{a}^{*}+{\bm{\theta}}_{k}^{*}).$
We have
$T_{121}$ $=\sup_{u\in[0,1]}\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\rho_{j}\over A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}\right\|_{\mathrm{op}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}$
$\leq\sup_{u\in[0,1]}\sum_{a\neq k}\max_{j\in[p]}{\rho_{j}\over A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}\left\|\sum_{j=1}^{p}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}\right\|_{\mathrm{op}}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}.$
After we invoke Lemmas 9 and 10, we find that
$T_{121}\lesssim\sum_{a\neq k}\left(\overline{\sigma}^{2}+{\overline{\sigma}^{4}\over 4}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}.$ (83)
We can repeat the same arguments to prove that the bound in (83) also holds for $T_{122}$, and hence for $T_{12}$.
Now, regarding the term $T_{11}$, by using the midpoint (79), we have
$T_{11}$ $={|\alpha_{k}-\alpha_{k}^{*}|\over\alpha_{k}^{*}}\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}.$
Since
$\left\|\sum_{j=1}^{p}{\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}$
$=\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}{\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})$
$\leq\sup_{v\in\mathbb{S}^{L-1}}\left(\sum_{j=1}^{p}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})[v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})]^{2}\right)^{1/2}\left(\sum_{j=1}^{p}\left({\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}\right)^{2}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})\right)^{1/2}$
$\leq\left\|\sum_{j=1}^{p}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}\right\|_{\mathrm{op}}^{1/2}\max_{j\in[p]}{\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})},$
invoking Lemma 9 and Lemma 10 gives that
$T_{11}\lesssim{|\alpha_{k}-\alpha_{k}^{*}|\over\alpha_{k}^{*}}\sum_{a\neq k}\left(\overline{\sigma}+{\overline{\sigma}^{2}\over 2}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}\right)(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right).$ (84)
Finally, combing (83) and (84) yields that
$\begin{split}T_{1}&\lesssim\sum_{a\neq k}\left(\overline{\sigma}^{2}+{\overline{\sigma}^{4}\over 4}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\\ &\quad+\sum_{a\neq k}\left(\overline{\sigma}+{\overline{\sigma}^{2}\over 2}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}\right)(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right){|\alpha_{k}-\alpha_{k}^{*}|\over\alpha_{k}^{*}}\\ &\lesssim\left(\overline{\sigma}+\overline{\sigma}^{2}\Delta\right)\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}\left({|\alpha_{k}-\alpha_{k}^{*}|\over\alpha_{k}^{*}}+\left(\overline{\sigma}+\overline{\sigma}^{2}\Delta\right)\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\right).\end{split}$ (85)
Bounding $T_{2}$.
Bounding $T_{2}$ essentially follows the same arguments as that of $T_{1}$. Start with $T_{2}\leq T_{21}+T_{22}$ where
$T_{21}:=\sum_{a\neq k}{|\alpha_{a}-\alpha_{a}^{*}|\over\alpha_{a}^{*}}\left\|\sum_{j=1}^{p}{\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\alpha_{a}^{*}A_{{\bm{\theta}}_{a}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2},$
$T_{22}:=\sum_{a\neq k}\left\|\sum_{j=1}^{p}(A_{{\bm{\theta}}_{a}}(x_{j})-A_{{\bm{\theta}}_{a}^{*}}(x_{j})){\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}.$
Note that
$T_{21}\leq{\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\alpha_{a}^{*}A_{{\bm{\theta}}_{a}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}.$
The same arguments of bounding $T_{11}$ above gives
$T_{21}\lesssim\left(\overline{\sigma}+{\overline{\sigma}^{2}\over 2}\Delta\right)\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}{\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}.$ (86)
On the other hand, repeating the arguments of bounding $T_{12}$ gives that
$T_{22}\lesssim\left(\overline{\sigma}^{2}+{\overline{\sigma}^{4}\over 4}\Delta^{2}\right)\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}\max_{a\in[K]}\|{\bm{\theta}}_{a}-{\bm{\theta}}_{a}^{*}\|_{2}.$ (87)
Collecting (86), (87) as well as (85) yields that
$\|\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})\|_{2}$
$\quad\lesssim\left(\overline{\sigma}+\overline{\sigma}^{2}\Delta\right)\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}\left({\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}+\left(\overline{\sigma}+\overline{\sigma}^{2}\Delta\right)\max_{a\in[K]}\|{\bm{\theta}}_{a}-{\bm{\theta}}_{a}^{*}\|_{2}\right).$
Finally, we complete the proof of (69) by observing that both Lemma 9 and Lemma 10 as well as the arguments above are valid uniformly over ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$.
Next we prove (68). By definition in (12), we can split, for any $k\in[K]$,
$M_{k}({{\bm{\omega}}})-M_{k}({{\bm{\omega}}^{*}})$ $=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(x_{j})\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}^{*}}(x_{j})}\right)=S_{1}+S_{2}$
with
$S_{1}$ $=\sum_{a\neq k}\sum_{j=1}^{p}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})-\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})$
$S_{2}$ $=\sum_{a\neq k}\sum_{j=1}^{p}{\alpha_{a}A_{{\bm{\theta}}_{a}}(x_{j})-\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})$
Bounding of $S_{1}$.
We start with the decomposition of $S_{1}$
$S_{1}$ $={\alpha_{k}-\alpha_{k}^{*}\over\alpha_{k}^{*}}\sum_{a\neq k}\sum_{j=1}^{p}{\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+\sum_{a\neq k}\sum_{j=1}^{p}\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j}){A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}$
$:=S_{11}+S_{12}.$
Using the midpoint notation in (79), we find that
$|S_{11}|$ $={|\alpha_{k}-\alpha_{k}^{*}|\over\alpha_{k}^{*}}\sum_{a\neq k}\sum_{j=1}^{p}{\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{j}(\bar{\bm{\theta}}_{ak}^{*})}A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})$
$\leq{|\alpha_{k}-\alpha_{k}^{*}|\over\alpha_{k}^{*}}\sum_{a\neq k}\max_{j\in[p]}{\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})A_{\bar{\bm{\theta}}_{ak}^{*}}(x_{j})}$
[…]
Regarding $S_{12}$, we have that
$|S_{12}|$ $\leq\sum_{a\neq k}\left|\sum_{j=1}^{p}{\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}(A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{*}}(x_{j}))\right|$
$=\sum_{a\neq k}\left|\sum_{j=1}^{p}{\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\int_{0}^{1}A_{{\bm{\theta}}_{k,u}}(x_{j})({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*})^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\mathrm{d}u\right|$
$\leq\sup_{u\in[0,1]}\sum_{a\neq k}\left\|\sum_{j=1}^{p}{\alpha_{k}^{*}\alpha_{a}^{*}A_{{\bm{\theta}}_{a}^{*}}(x_{j})A_{{\bm{\theta}}_{k,u}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}.$
By repeating the same arguments of bounding $T_{21}$ above, we further find
$|S_{12}|\lesssim\sum_{a\neq k}\left(\overline{\sigma}+{\overline{\sigma}^{2}\over 2}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}\right)(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}.$ (89)
Bounding $S_{2}$.
Using the inequality
$|S_{2}|$ $\leq\sum_{a\neq k}{|\alpha_{a}-\alpha_{a}^{*}|\over\alpha_{a}^{*}}\sum_{j=1}^{p}{\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})A_{{\bm{\theta}}_{a}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+\sum_{a\neq k}\left|\sum_{j=1}^{p}{\alpha_{a}^{*}\alpha_{k}^{*}A_{{\bm{\theta}}_{k}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}(A_{{\bm{\theta}}_{a}}(x_{j})-A_{{\bm{\theta}}_{a}^{*}}(x_{j}))\right|$
and after repeating the above arguments, we find that
$|S_{2}|$ $\lesssim{\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}\sum_{a\neq k}(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right)$
$\quad+\sum_{a\neq k}\|{\bm{\theta}}_{a}-{\bm{\theta}}_{a}^{*}\|_{2}\left(\overline{\sigma}+{\overline{\sigma}^{2}\over 2}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}\right)(\alpha_{a}^{*}\vee\alpha_{k}^{*})\exp\left(-{\underline{\sigma}^{2}\over 8}\|{\bm{\theta}}_{a}^{*}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}\right).$ (90)
Combining (B.2.3), (89) and (B.2.3) yields the following bound
$\max_{k\in[K]}|M_{k}({{\bm{\omega}}})-M_{k}({{\bm{\omega}}^{*}})|$ $\lesssim\exp\left(-{\underline{\sigma}^{2}\over 8}\Delta^{2}\right)K\overline{\alpha}\left({\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\over\underline{\alpha}}+\left(1+\overline{\sigma}\Delta\right)\overline{\sigma}\max_{a\in[K]}\|{\bm{\theta}}_{a}-{\bm{\theta}}_{a}^{*}\|_{2}\right)$
for any fixed ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$. Since the arguments hold uniformly over ${\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$, the proof is complete. ∎
B.2.4 Proof of Lemma 6: concentration inequality of the EM-updates within the specified neighborhood
Proof.
Our proof is based on the following discretization of
${\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})=\left\{{\bm{\omega}}:\|{\bm{\alpha}}-{\bm{\alpha}}^{*}\|_{\infty}\leq\delta_{0}\underline{\alpha},~\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma},~\forall~k\in[K]\right\}.$
For any given $\epsilon_{1}\in(0,\underline{\alpha}/4]$ and $\epsilon_{2}\in(0,\delta_{0}/\overline{\sigma}]$, let $\mathcal{N}_{\epsilon_{1}}(\Delta^{K})$ be an $\epsilon_{1}$-covering set (in $\ell_{\infty}$-norm) of $\Delta^{K}$ and $\mathcal{N}_{\epsilon_{2}}$ be the $\epsilon_{2}$-net (in $\ell_{2}$-norm) of $\{{\bm{\theta}}\in\mathbb{R}^{L}:\|{\bm{\theta}}\|_{2}\leq\delta_{0}/\overline{\sigma}\}$. Then, for any $k\in[K]$, $\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{k}^{*}):=\left\{{\bm{\theta}}+{\bm{\theta}}_{k}^{*}:{\bm{\theta}}\in\mathcal{N}_{\epsilon_{2}}\right\}$ is the $\epsilon_{2}$-net of $\{{\bm{\theta}}\in\mathbb{R}^{L}:\|{\bm{\theta}}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}/\overline{\sigma}\}$. Consider the set
$\mathcal{N}_{\epsilon_{1},\epsilon_{2}}=\mathcal{N}_{\epsilon_{1}}(\Delta^{K})\otimes\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{1}^{*})\otimes\cdots\otimes\mathcal{N}_{\epsilon_{2}}({\bm{\theta}}_{K}^{*}).$
We have that for any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$, there exists some ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ such that
$\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\leq\epsilon_{1},\quad\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}\leq\epsilon_{2}$ (91)
as well as
$\|{\bm{\alpha}}^{*}-{\bm{\alpha}}^{\prime}\|_{\infty}\leq\epsilon_{1}+\delta_{0}\underline{\alpha},\quad\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k}^{\prime}\|_{2}\leq\delta_{0}/\overline{\sigma}.$ (92)
Moreover, from Ghosal and van der Vaart (2001, Lemma A.4) and the classical result on the covering number of an Euclidean ball, the cardinality of $\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfies
$|\mathcal{N}_{\epsilon_{1},\epsilon_{2}}|\leq|\mathcal{N}_{\epsilon_{1}}(\Delta^{K})||\mathcal{N}_{\epsilon_{2}}|^{K}\leq\left(5\over\epsilon_{1}\right)^{K-1}\left(3\delta_{0}\over\overline{\sigma}\epsilon_{2}\right)^{KL}.$ (93)
Since for any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$, there exists ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (91) – (92) such that for all $k\in[K]$,
$|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|\leq|\widehat{M}_{k}({\bm{\omega}}^{\prime})-M_{k}({\bm{\omega}}^{\prime})|+|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})|,$ (94)
we first bound the second term:
$\begin{split}&\ \left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})\right|\\ &=\left|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j}))\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right)\right|\\ &\leq\|\widehat{\pi}-\pi_{{\bm{\omega}}^{*}}\|_{1}\max_{j\in[p]}\left|{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right|\\ &\leq 2\max_{j\in[p]}\left(|\alpha_{k}-\alpha_{k}^{\prime}|{A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\\ &\leq 2\max_{j\in[p]}\left({\epsilon_{1}\over\alpha_{k}}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}+{\alpha^{\prime}_{k}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\\ &\leq 2\max_{j\in[p]}\left({\epsilon_{1}\over\alpha_{k}}+{\alpha^{\prime}_{k}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}+{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{\prime}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right).\end{split}$ (95)
In order to invoke our perturbation bounds in Lemma 7, we need to verify that its conditions are satisfied. This follows by noting that $\epsilon_{2}\leq\delta_{0}/\overline{\sigma}\leq c_{0}/\|\mathbf{X}\|_{\infty,2}$ under (20), $\delta_{0}\leq c_{0}<1/2$ and
$\alpha_{k}\geq\alpha_{k}^{*}-\delta_{0}\underline{\alpha}\geq\alpha_{k}^{*}-{1\over 2}\alpha_{k}^{*}\geq{1\over 2}\underline{\alpha}\geq 2\epsilon_{1}.$ (96)
Hence, invoking Lemma 7 gives
$\max_{k\in[K]}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})-\widehat{M}_{k}({\bm{\omega}}^{\prime})+M_{k}({\bm{\omega}}^{\prime})\right|$
$\quad\leq{4\epsilon_{1}\over\underline{\alpha}}+{2\alpha_{k}^{\prime}\over\alpha_{k}}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}~3\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{4\epsilon_{1}\over\underline{\alpha}}+8\epsilon_{2}\|\mathbf{X}\|_{\infty,2}$
$\quad\lesssim\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}.$ (97)
The last step uses (96) and
$\alpha_{k^{\prime}}\leq\alpha_{k}+\epsilon_{1}\leq\alpha_{k}^{*}+\delta_{0}\underline{\alpha}+\underline{\alpha}/4\leq 2\alpha_{k}^{*}.$ (98)
In conjunction with (94), we further obtain
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\max_{k\in[K]}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})\right|\lesssim\max_{{\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\max_{k\in[K]}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})\right|+\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}.$ (99)
We proceed to bound from above the first term on the right. To this end, fix any ${\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (92). We find that
$\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})\right|$ $=\left|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j})){\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\right|={1\over N}\left|\sum_{i=1}^{N}\left(E_{i}-\pi_{{\bm{\omega}}^{*}}\right)^{\top}h_{k}\right|$
[…]
$\operatorname{Var}(E_{i}^{\top}h_{k})$ $\leq\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(x_{j}){\alpha_{k}^{2}A_{{\bm{\theta}}_{k}}(x_{j})^{2}\over\pi_{{\bm{\omega}}}(x_{j})^{2}}$
$\leq\left(1+{|\pi_{{\bm{\omega}}^{*}}(x_{j})-\pi_{{\bm{\omega}}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\sum_{j=1}^{p}{\alpha_{k}^{2}A_{{\bm{\theta}}_{k}}(x_{j})^{2}\over\pi_{{\bm{\omega}}}(x_{j})}$
$\leq\left(1+{\epsilon_{1}\over\underline{\alpha}}+\delta_{0}+4\epsilon_{2}\|\mathbf{X}\|_{\infty,2}\right)\alpha_{k}$ by (73) in Lemma 7 and (92)
$\leq 4\alpha_{k}.$
The last step uses (20), $\delta_{0}\leq c_{0}<1/2$, $\epsilon_{2}\leq\delta_{0}/\overline{\sigma}$ and $\epsilon_{1}\leq\underline{\alpha}/4$. Further note that
$|E_{i}^{\top}h_{k}|\leq\max_{j\in[p]}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\leq 1.$
An application of the Bernstein inequality together with the union bounds argument yields that, for any $t>0$,
$\mathbb{P}\left\{\max_{{\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\max_{k\in[K}\left|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})\right|\gtrsim\sqrt{\alpha_{k}t\over N}+{t\over N}\right\}$
$\qquad\leq 2\exp\left(-t+\log|\mathcal{N}_{\epsilon_{1},\epsilon_{2}}|\right)$
[…]
In view of (99), by invoking the event $\mathcal{E}_{2}$, the proof of the first result is completed by choosing
$\epsilon_{1}=\underline{\alpha}\left({1\over 4}\wedge{KL\over N}\right),\qquad\epsilon_{2}={1\over\overline{\sigma}}\left(\delta_{0}\wedge{\overline{\sigma}\over\|\mathbf{X}\|_{\infty,2}}{KL\over N}\right),\qquad t=CKL\log(N)$
and using (25) to collect terms.
We use similar argument to bound from above
$\sup_{{\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}\leq\max_{{\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})\|_{2}$
[…]
Pick any ${\bm{\omega}}\in{\mathbb{B}}_{d}({\bm{\omega}}^{*},\delta_{0})$ and ${\bm{\omega}}^{\prime}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ satisfying (91). We bound from above
$\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})+\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}^{\prime}\mid{\bm{\omega}}^{\prime})\|_{2}$
$=\left\|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j}))\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}^{\prime}})\right)\right\|_{2}$
$\leq\left\|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j}))\left({\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right)\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\right\|_{2}$
$\qquad+\left\|\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j})){\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}^{\prime}})\right\|_{2}$
$\leq\|\widehat{\pi}-\pi_{{\bm{\omega}}^{*}}\|_{1}\max_{j\in[p]}\left|{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right|\|\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\|_{2}$
$\qquad+\|\widehat{\pi}-\pi_{{\bm{\omega}}^{*}}\|_{1}\max_{j\in[p]}{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\|\mathbf{X}^{\top}(A_{{\bm{\theta}}_{k}}-A_{{\bm{\theta}}_{k}^{\prime}})\|_{2}$
$\leq 4\max_{j\in[p]}\left|{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}-{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})\over\pi_{{\bm{\omega}}^{\prime}}(x_{j})}\right|\|\mathbf{X}\|_{\infty,2}+2\|\mathbf{X}\|_{\infty,2}\max_{j\in[p]}|A_{{\bm{\theta}}_{k}}(x_{j})-A_{{\bm{\theta}}_{k}^{\prime}}(x_{j})|.$
By the argument in (95), (72) and (73), the above is bounded from above by (in order)
$\left(\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}\right)\|\mathbf{X}\|_{\infty,2}.$ (100)
It remains to bound from above
$\max_{{\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}$
$\quad\leq 2\max_{{\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\max_{v\in\mathcal{N}_{L}(1/2)}v^{\top}\left(\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\right)$
$\quad=2\max_{{\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\max_{v\in\mathcal{N}_{L}(1/2)}\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j})){\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})$
where $\mathcal{N}_{L}(1/2)$ is the (1/2)(1/2)( 1 / 2 )-net of $\mathbb{S}^{L-1}$ and satisfies $|\mathcal{N}_{L}(1/2)|\leq 5^{L}$ (see, for instance, Vershynin (2018)). Fix any ${\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ and $v\in\mathcal{N}_{L}(1/2)$. Observe that
$\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j})){\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}}):={1\over N}\sum_{i=1}^{N}(E_{i}-\pi_{{\bm{\omega}}^{*}})^{\top}h_{v}$
with
$[h_{v}]_{j}={\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}}),\qquad\forall j\in[p].$
Also note that
$\operatorname{Var}(E_{i}^{\top}h_{v})$ $\leq\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(x_{j}){\alpha_{k}^{2}A_{{\bm{\theta}}_{k}}(x_{j})^{2}\over\pi_{{\bm{\omega}}}(x_{j})^{2}}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}v$
$\leq\max_{j\in[p]}{\pi_{{\bm{\omega}}^{*}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}{\alpha_{k}^{2}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}\sum_{j=1}^{p}A_{{\bm{\theta}}_{k}}(x_{j})v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})^{\top}\mathbf{X}v$
$\leq\alpha_{k}\max_{j\in[p]}\left(1+{|\pi_{{\bm{\omega}}}(x_{j})-\pi_{{\bm{\omega}}^{*}}(x_{j})|\over\pi_{{\bm{\omega}}}(x_{j})}\right)\lambda_{1}\left(\mathbf{X}^{\top}\Sigma_{A_{{\bm{\theta}}_{k}}}\mathbf{X}\right)$
$\overset{(i)}{\lesssim}\alpha_{k}\left(1+\epsilon_{2}\|\mathbf{X}\|_{\infty,2}+{\epsilon_{1}\over\underline{\alpha}}+\delta_{0}\right)\lambda_{1}\left(H_{{\bm{\theta}}_{k}}\right)$
$\overset{(ii)}{\lesssim}\alpha_{k}~\overline{\sigma}^{2}$
[…]
$E_{i}^{\top}h_{v}\leq\max_{j\in[p]}{\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}|v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})|\leq 2\|\mathbf{X}\|_{\infty,2},$
applying Bernstein’s inequality and the union bound over ${\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}$ and $v\in\mathcal{N}_{L}(1/2)$ yields
$\max_{{\bm{\omega}}\in\mathcal{N}_{\epsilon_{1},\epsilon_{2}}}\max_{v\in\mathcal{N}_{L}(1/2)}\sum_{j=1}^{p}(\widehat{\pi}_{j}-\pi_{{\bm{\omega}}^{*}}(x_{j})){\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j})\over\pi_{{\bm{\omega}}}(x_{j})}v^{\top}\mathbf{X}^{\top}(\mathbf{e}_{j}-A_{{\bm{\theta}}_{k}})\lesssim\overline{\sigma}\sqrt{\alpha_{k}t\over N}+{t\|\mathbf{X}\|_{\infty,2}\over N}$
with probability at least
$1-2\exp\left(-t+\log|\mathcal{N}_{\epsilon_{1},\epsilon_{2}}|+\log|\mathcal{N}_{L}(1/2)|\right)$
[…]
We complete the proof by choosing
$\epsilon_{1}=\underline{\alpha}\left({1\over 4}\wedge{KL\over N}\right),\qquad\epsilon_{2}={\delta_{0}\over\overline{\sigma}}\wedge{KL\over\|\mathbf{X}\|_{\infty,2}N},\qquad t=CKL\log(N),$
taking the union bounds over $k\in[K]$ and using (25) to collect terms. ∎
Appendix C Proofs of Section 3
C.1 Proof of Lemma 1
Proof.
Results (32) and (34) above can be found in Lindsay (1989), whereas (33) is implicit in Lindsay and Basak (1993), and we derive its explicit form here.
The first and third result of Lemma 1 have been known for several decades, in the theory on univariate mixtures. Consider the first coordinate $\theta_{11}^{*},\ldots,\theta_{1K}^{*}$, respectively, of the $K$ parameter vectors ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ in $\mathbb{R}^{L}$. By assumption, they are distinct and, in the notation of Section 3, they are the $K$ support points of the one-dimensional distribution of $Z_{1}$, the first coordinate of the latent vector $Z\sim\rho^{*}$. Recall that $m_{1},\ldots,m_{2K-1}$ are, by definition, moments of $Z_{1}$. Then by Theorem 2C in Lindsay (1989) (population version), the polynomial equation $P(x)=0$ has $K$ distinct roots, and they are equal to $\theta_{11}^{*},\ldots,\theta_{1K}^{*}$.
Next, one forms the system of equations $m_{r}=\sum_{k=1}^{K}\alpha_{k}\theta_{1k}^{*r}$, for $0\leq r\leq K-1$, which for given $m_{r}$, and for $\theta_{1k}^{*}$ found above, is linear in $\alpha_{1},\ldots,\alpha_{K}$. Since its coefficient matrix is a Vandermonde matrix, it is invertible, and the system has the unique solution ${\bm{\alpha}}^{*}$ given by (34).
Lindsay and Basak (1993) gave the road map to extending the univariate result to the multivariate case and we make it explicit here, in our notation. Consider the matrix of moments
$\bm{M}:=\begin{pmatrix}1&m_{1}&\dots&m_{K-1}\\ m_{1}&m_{2}&\dots&m_{K}\\ \vdots&\vdots&&\vdots\\ m_{K-1}&m_{K}&\dots&m_{2K-2}\end{pmatrix}$
By Theorem 2A of Lindsay (1989), this matrix is non-singular. Consider now the following $(K+1)\times(K+1)$ matrix
$U(t)=\begin{pmatrix}\bm{M}&{\bm{a}}\\ {\bm{b}}^{\top}&t\end{pmatrix},$
for some generic vectors $\bm{a},\bm{b}\in\mathbb{R}^{K}$, and a scalar $t\in\mathbb{R}$.
On the one hand, we have the following facts. Using the formula for block matrix determinants, we have
$\det(U(t))=\det({\bm{M}})\det(t-{\bm{b}}^{\top}{\bm{M}}^{-1}{\bm{a}}).$
Since $\det({\bm{M}})>0$, the unique solution to $\det(U(t))=0$ is given by
$t={\bm{b}}^{\top}{\bm{M}}^{-1}{\bm{a}}.$ (101)
On the other hand, we also have the following. Since $\det({\bm{M}})>0$, then $\text{rank}(U(t))\geq K$, with maximal possible rank $K+1$. We now choose $\bm{a},\bm{b}$ and $t$ such that $\text{rank}(U(t))=K$, and thus such that $\det(U(t))=0$.
The choices ${\bm{a}}:=(\theta_{1k}^{*r})_{r=0}^{K-1}$, ${\bm{b}}:=(m_{r,i})_{r=0}^{K-1}$, $t=\theta_{ik}^{*}$ indeed cause this quantity to vanish, since then the $K+1$ columns of $U(\theta_{ik}^{*})$ are spanned by the $K$ vectors $(1,\theta_{1k}^{*},\dots,(\theta_{1k}^{*})^{(K-1)},\theta_{ik}^{*})^{\top}$, $k=1,\dots,K$. Combining this with (101) gives the stated expression (33) ∎
C.2 Proof of Proposition 1
The proof of Proposition 1 follows immediately from the following Lemma.
Lemma 11.
Let $h_{r}$ and $h_{r1,i}$ be defined as in (38) and (39). Let $X\sim\mu$, where $\mu$ satisfies Assumption 2. Then, for any ${\bm{\theta}}\in\mathbb{R}^{L}$,
$\mathbb{E}_{\mu}\left[h_{r}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]$ $=(v^{\top}{\bm{\theta}})^{r}\,$ (102)
$\mathbb{E}_{\mu}\left[h_{r1;i}(X)\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]$ $=(v^{\top}{\bm{\theta}})^{r}(w_{i}^{\top}{\bm{\theta}})\,,$ (103)
for $i\in\{2,\ldots,L\}$.
Proof.
It suffices to prove the first claim, since the second follows by differentiating both sides of Eq. 102 with respect to $v$ and applying dominated convergence.
Write $g_{r}(X;v,t)=(-1)^{r}\mu(X)^{-1}\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\mu(X+tv)\exp(X^{\top}{\bm{\theta}})$. We will show by induction that
$\mathbb{E}_{\mu}\left[g_{r}(X;v,t)\right]=\left(v^{\top}{\bm{\theta}}\right)^{r}\mathbb{E}_{\mu}[\exp((X-tv)^{\top}{\bm{\theta}})]$ (104)
for all $t\in\mathbb{R}$ and $v\in\mathbb{R}^{L}$, and conclude by taking $t=0$.
When $r=0$, we have
$\mathbb{E}_{\mu}\left[g_{0}(X;v,t)\right]$ $=\int\mu(x+tv)\exp(x^{\top}{\bm{\theta}})\,\mathrm{d}x$
$=\int\mu(x)\exp((x-tv)^{\top}{\bm{\theta}})\,\mathrm{d}x$
$=\mathbb{E}_{\mu}[\exp((X-tv)^{\top}{\bm{\theta}})]\,.$
Now assume Eq. 104 holds for a natural number $r$. The assumption that the partial derivatives of $\mu$ decay super-exponentially implies that we can apply dominated convergence to obtain
$\mathbb{E}_{\mu}\left[g_{r+1}(X;v,t)\right]$ $=-\mathbb{E}_{\mu}\left[\frac{\mathrm{d}}{\mathrm{d}t}g_{r}(X;v,t)\right]$
$=-\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}_{\mu}\left[g_{r}(X;v,t)\right]$
$=-\frac{\mathrm{d}}{\mathrm{d}t}(v^{\top}{\bm{\theta}})^{r}\mathbb{E}_{\mu}[\exp((X-tv)^{\top}{\bm{\theta}})]$
$=(v^{\top}{\bm{\theta}})^{r+1}\mathbb{E}_{\mu}[\exp((X-tv)^{\top}{\bm{\theta}})]\,.$
When $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$, we recover the expressions given in Proposition 1. ∎
C.3 Proof of Proposition 2
Proof.
We first show that the bound holds for the first coordinate. By re-scaling, we may assume $B=1$. To this end, we use existing results in  Wu and Yang (2020). To begin with, we recall here Assumption 4, and fix $\Delta_{1}$ and $\underline{\alpha}$. Define
$\epsilon:={\Delta_{1}\underline{\alpha}\over 4}$
and write $\rho_{1}^{*}=\sum_{k}\alpha_{k}^{*}\delta_{\theta_{1k}^{*}}$ and $\widetilde{\rho}_{1}=\sum_{k}\bar{\alpha}_{k}\delta_{\bar{\theta}_{1k}}.$ By Proposition 1 in Wu and Yang (2020), there exists $c^{\prime}=c^{\prime}(K)$ such that, if $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}\leq c^{\prime}$, then $W(\rho_{1}^{*},\widetilde{\rho}_{1})\leq\epsilon$. We will show that our result holds by considering, separately, $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}\leq c^{\prime}$ and $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}>c^{\prime}$. We begin with the former.
We show that $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}\leq c^{\prime}$, for $c^{\prime}$ above, implies that:
  • (i)
    There exists a permutation $\varrho$ of $K$ integers such that
    $|\theta_{1k}-\bar{\theta}_{1\varrho(k)}|\leq W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})/\underline{\alpha},\ \ \mbox{for each}\ k\in[K].$ (105)
  • (ii)
    There exists a constant $C$ depending on $K$ and $\Delta_{1}$ such that
    $W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})\leq C\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}.$ (106)
The claimed result, in this case, will then follow by combining (105) and (106).
By the definition of the Wasserstein distance, and for $\Pi$ denoting a distribution with marginals $\widetilde{\rho}_{1}$ and $\rho_{1}^{*}$, we have
$W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})$ $=\inf_{\Pi}\sum_{k,k^{\prime}}\Pi_{kk^{\prime}}|\theta_{1k}^{*}-\bar{\theta}_{1k^{\prime}}|$
$\geq\sum_{k=1}^{K}\alpha_{k}^{*}\min_{k^{\prime}\in[K]}|\theta_{1k}^{*}-\bar{\theta}_{1k^{\prime}}|$
$\geq\underline{\alpha}\max_{k\in[K]}\min_{k^{\prime}\in[K]}|\theta_{1k}^{*}-\bar{\theta}_{1k^{\prime}}|,$ (107)
and so
$\min_{k^{\prime}\in[K]}|\theta_{1k}^{*}-\bar{\theta}_{1k^{\prime}}|\leq{W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})\over\underline{\alpha}},\quad\mbox{for each}\ k\in[K].$
Then, there must exist a permutation $\varrho$ such that (105) holds. Otherwise, suppose there exists some $\varrho(k)=\varrho(k^{\prime})$ for some $k\neq k^{\prime}$ such that
$|\theta_{1k}^{*}-\bar{\theta}_{1\varrho(k)}|\leq W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})/\underline{\alpha},\qquad|\theta_{1k^{\prime}}^{*}-\bar{\theta}_{1\varrho(k^{\prime})}|\leq W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})/\underline{\alpha}.$
This however leads to the contradiction
$\Delta_{1}\leq|\theta_{1k}^{*}-\theta_{1k^{\prime}}^{*}|\leq|\theta_{1k}^{*}-\bar{\theta}_{1\varrho(k)}|+|\theta_{1k^{\prime}}^{*}-\bar{\theta}_{1\varrho(k^{\prime})}|\leq{2W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})\over\underline{\alpha}}\leq{2\epsilon\over\underline{\alpha}}\leq{\Delta_{1}\over 2},$
where the penultimate inequality uses $W(\rho_{1}^{*},\widetilde{\rho}_{1})\leq\epsilon$ from $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}\leq c^{\prime}$ and the last inequality follows by the definition of $\epsilon$. This proves (105).
To show (106), without loss of generality, we assume $\varrho$ is the identity permutation. We first notice that
$\min_{k\neq k^{\prime}}|\bar{\theta}_{1k}-\bar{\theta}_{1k^{\prime}}|\geq\min_{k\neq k^{\prime}}|\theta_{1k}^{*}-\theta_{1k^{\prime}}^{*}|-2\max_{k}|\theta_{1k}-\bar{\theta}_{1k}|\geq\Delta_{1}-{2\epsilon\over\underline{\alpha}}\geq{\Delta_{1}\over 2}$ (108)
and, similarly,
$\min_{k\neq k^{\prime}}|\theta_{1k}^{*}-\bar{\theta}_{1k^{\prime}}|\geq\min_{k\neq k^{\prime}}|\theta_{1k}-\theta_{1k^{\prime}}|-\max_{k}|\theta_{1k}-\bar{\theta}_{1k}|\geq\Delta_{1}-{2\epsilon\over\underline{\alpha}}\geq{\Delta_{1}\over 2}$
Thus, the atoms of $\widetilde{\rho}_{1}$ and $\rho_{1}^{*}$ are all separated by at least $\Delta_{1}/2$. This places us in the setting of Proposition 4 in Wu and Yang (2020), which we apply (relative to their notation) with $\gamma=\Delta_{1}/2$, $l=2K$ and $\ell^{\prime}=1$ yielding the bound
$W_{1}(\widetilde{\rho}_{1},\rho_{1}^{*})\leq{4K^{2}4^{2K-1}\over\Delta_{1}^{2K-2}}\|\widetilde{\bm{m}}-{\bm{m}}\|_{2},$ (109)
which completes the proof of (106) and thus, for all $k\in[K]$
$|\bar{\theta}_{1k}-\theta_{1\varrho(k)}^{*}|\leq C^{\prime}\|\widetilde{\bm{m}}-{\bm{m}}\|_{2},$
by taking
$C^{\prime}:={4K^{2}4^{2K-1}\over\underline{\alpha}\Delta_{1}^{2K-2}}.$ (110)
On the other hand, when $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}>c^{\prime}$, we have
$|\bar{\theta}_{1k}-\theta_{1k}^{*}|\leq 2<\frac{2}{c^{\prime}}\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}\leq D_{1}\|\widetilde{\bm{m}}-{\bm{m}}\|_{2},$
for $D_{1}:=\max\{2/c^{\prime},C^{\prime}\}$, and where the first inequality holds since, for each $k\in[K]$, we have $\bar{\theta}_{1k},\theta_{1k}\in[-1,1]$.
Therefore, for all $k\in[K]$, there exist a constant $D_{1}$ as above such that
$|\bar{\theta}_{1k}-\theta_{1k}^{*}|\leq D_{1}\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}$ (111)
We next fix $i\in\{2,\ldots,L\}$ and $k\in[K]$ and show that estimation error of the remaining coordinates $\theta_{ik}^{*}$ has upper bound similar to (111), for a different constant $D_{2}$. Recall that
$\bm{m}_{1;i}=\left(m_{01;i},\ldots,m_{(K-1)1;i}\right)^{\top}\qquad\bar{\bm{m}}_{1;i}=\left(\bar{m}_{01;i},\ldots,\bar{m}_{(K-1)1;i}\right)^{\top}$
and let
$\xi:=\left(1,\theta_{1k},\ldots,\theta_{1k}^{K-1}\right)^{\top},\qquad\bar{\xi}:=\left(1,\bar{\theta}_{1k},\ldots,\bar{\theta}_{1k}^{K-1}\right)^{\top}\,.$
Finally, define the operator $\mathsf{clip}_{B}$ by
$\mathsf{clip}_{B}(x)=\begin{cases}-B&\text{if $x<-B$}\\ x&\text{if $|x|\leq B$}\\ B&\text{if $x>B$.}\end{cases}$
Using the definition of $\bar{\theta}_{ik}$ and (33), we can therefore write
$\bar{\theta}_{ik}-\theta_{ik}^{*}$ $=\mathsf{clip}_{B}(\bar{\bm{m}}_{1;i}^{\top}\widetilde{M}^{\dagger}\bar{\xi})-\bm{m}_{1;i}^{\top}M^{-1}\xi\,,$
where the $K\times K$ matrix $\widetilde{M}$ is obtained from
$\widetilde{M}:=\begin{pmatrix}1&\widetilde{m}_{1}&\dots&\widetilde{m}_{K-1}\\ \widetilde{m}_{1}&\widetilde{m}_{2}&\dots&\widetilde{m}_{K}\\ \vdots&\vdots&&\vdots\\ \widetilde{m}_{K-1}&\widetilde{m}_{K}&\dots&\widetilde{m}_{2K-2}\end{pmatrix}.$
We consider two cases. As in the preceding argument, there exists a constant $c^{\prime}$ depending on $K$, $\Delta_{1}$, $B$, and $\underline{\alpha}$ such that if $\|\widetilde{\bm{m}}-\bm{m}\|_{2}\leq c^{\prime}$, then the measure $\widetilde{\rho}_{1}$ corresponding to $\widetilde{\bm{m}}$ has $K$ atoms, each separated by at least $\Delta_{1}/4$.
Under this scenario, Lindsay (1989, Theorem 2A) implies that $\widetilde{M}$ is invertible, and we obtain
$|\bar{\theta}_{ik}-\theta_{ik}^{*}|$ $=|\mathsf{clip}_{B}(\bar{\bm{m}}_{1;i}^{\top}\widetilde{M}^{-1}\bar{\xi})-\bm{m}_{1;i}^{\top}M^{-1}\xi|$
$\leq|\bar{\bm{m}}_{1;i}^{\top}\widetilde{M}^{-1}\bar{\xi}-\bm{m}_{1;i}^{\top}M^{-1}\xi|$
$\leq|\bar{\bm{m}}_{1;i}^{\top}(\widetilde{M}^{-1}-M^{-1})\bar{\xi}|+|(\bar{\bm{m}}_{1;i}-\bm{m}_{1;i})^{\top}M^{-1}\xi|+|\bm{m}_{1;i}^{\top}M^{-1}(\bar{\xi}-\xi)|$ (112)
By Lindsay (1989, Theorem 2A), $M$ is invertible, and $\|M^{-1}\|_{\rm op}$ is bounded by a constant depending on $K$, $\Delta_{1}$, $B$, and $\underline{\alpha}$. Next, recall that we work under the assumption that $\|{\bm{\theta}}_{k}^{*}\|_{2}\leq B$, for all $k$, for some constant $B$ and thus, as $K$ is fixed, both $\|\bm{m}_{1;i}\|_{2}$ and $\|\xi\|_{2}$ are of order $\mathcal{O}(1)$.
Then, an application of the Cauchy-Schwarz inequality and (111) shows that the last two terms in (C.3) are bounded by a constant multiple of $\|\bar{\bm{m}}_{1;i}-\bm{m}_{1;i}\|_{2}+\|\widetilde{\bm{m}}-\bm{m}\|_{2}$.
For the first term in (C.3), we note that
$|\bar{\bm{m}}_{1;i}^{\top}(\widetilde{M}^{-1}-M^{-1})\bar{\xi}|\leq\|\bar{\bm{m}}_{1;i}\|_{2}\|\bar{\xi}\|_{2}\|\widetilde{M}^{-1}-M^{-1}\|_{\rm op}\,.$ (113)
The norms $\|\bar{\bm{m}}_{1;i}\|_{2}$ and $\|\bar{\xi}\|_{2}$ are both bounded by a constant depending on $B$ and $K$. Furthermore,
$\|\widetilde{M}^{-1}-M^{-1}\|_{\rm op}$ $=\|M^{-1}(M-\widetilde{M})\widetilde{M}^{-1}\|_{\rm op}\leq\|M^{-1}\|_{\rm op}\|\widetilde{M}^{-1}\|_{\rm op}\|M-\widetilde{M}\|_{\rm op}\,.$
As noted above, the assumption $\|\widetilde{\bm{m}}-\bm{m}\|_{2}\leq c^{\prime}$ implies that $\widetilde{\rho}_{1}$ has $K$ atoms separated by at least $\Delta_{1}/4$; this implies that $\|\widetilde{M}^{-1}\|_{\rm op}$ is also bounded by a constant depending on $K$, $\Delta_{1}$, $B$, and $\underline{\alpha}$. We obtain, for a constant $C^{\prime}$ different than above
$\|\widetilde{M}^{-1}-M^{-1}\|_{\rm op}\leq C^{\prime}\|\widetilde{\bm{m}}-\bm{m}\|_{2}\,.$
All together, we obtain that when $\|\widetilde{\bm{m}}-\bm{m}\|_{2}\leq c^{\prime}$, we have the bound
$|\bar{\theta}_{ik}-\theta_{ik}^{*}|\leq C^{\prime}(\|\bar{\bm{m}}_{1;i}-\bm{m}_{1;i}\|_{2}+\|\widetilde{\bm{m}}-\bm{m}\|_{2})\,.$ (114)
On the other hand, if $\|\widetilde{\bm{m}}-\bm{m}\|_{2}>c^{\prime}$, then the same argument as was given above shows that
$|\bar{\theta}_{ik}-\theta_{ik}^{*}|\leq 2B<D_{2}(\|\bar{\bm{m}}_{1;i}-\bm{m}_{1;i}\|_{2}+\|\widetilde{\bm{m}}-\bm{m}\|_{2})\,,$ (115)
where $D_{2}:=\max\{C^{\prime},2B/c^{\prime}\}$.
Finally, to establish the desired bound on $\bar{\bm{\alpha}}$, we use a very similar argument. Let $T$ be the Vandermonde matrix appearing on the right side of (34), and $\bar{T}$ its empirical counterpart in (45). If $\|\widetilde{\bm{m}}-\bm{m}\|_{2}\leq c^{\prime}$, then $T$ and $\bar{T}$ are both invertible, with smallest singular value bounded away from zero. We obtain, for some other constant $C^{\prime}$
$\|\bar{\bm{\alpha}}-\bm{\alpha}^{*}\|_{2}\lesssim\|T-\bar{T}\|_{\rm op}+\|\widetilde{\bm{m}}-\bm{m}\|_{2}\leq C^{\prime}\|\widetilde{\bm{m}}-\bm{m}\|_{2}\,.$ (116)
When $\|\widetilde{\bm{m}}-\bm{m}\|_{2}>c^{\prime}$, we use the trivial bound
$\|\bar{\bm{\alpha}}-\bm{\alpha}^{*}\|_{2}\leq 2<D_{3}\|\widetilde{\bm{m}}-\bm{m}\|_{2}$ (117)
for $D_{3}=\max\{C^{\prime},2/c^{\prime}\}$. Since $\|\widetilde{\bm{m}}-{\bm{m}}\|_{2}\leq\|\bar{\bm{m}}-{\bm{m}}\|_{2}$ from (43), taking $D=D_{1}\vee D_{2}\vee D_{3}$ completes the argument. ∎
C.4 Exponential dependence of $D$ on $K$ in Proposition 2
Remark 10.
Although we have stated Proposition 2 without explicit constants, the dependence on $K$, $B$, $\Delta_{1}$, and $\underline{\alpha}$ can be extracted from the proof. In particular, see, for instance, (110), the dependence on $\Delta_{1}$ is poor: the constant $D$ can be shown to scale as $\Delta_{1}^{-cK}$ for some absolute constant $c$. While it is possible that the exponent can be improved, the exponential dependence of this constant on $K$ cannot be entirely avoided, even for univariate mixtures ($L=1$). This follows from the fact that, when $L=1$, for any $K\geq 2$ and sufficiently small $\Delta_{1}>0$, there exist a pair of $K$-atomic probability measures $\rho$ and $\rho^{\prime}$ on [1,1][-1,1][ - 1 , 1 ] with support $\{{\bm{\theta}}_{1},\dots,{\bm{\theta}}_{K}\}$ and $\{{\bm{\theta}}^{\prime}_{1},\dots,{\bm{\theta}}^{\prime}_{K}\}$, all of which are separated by at least $\Delta_{1}$, and such that
$\|{\bm{\theta}}^{\prime}_{k}-{\bm{\theta}}_{k}\|_{2}\geq C_{K}\Delta_{1}^{-2K+2}\|\bm{m}-\bm{m}^{\prime}\|_{2},\quad\quad\text{for all $k\in[K]$}$ (118)
[…]
Proof of Remark 10.
[…]
$\|\bm{m}(\nu)-\bm{m}(\nu^{\prime})\|_{2}\leq 2\,.$
[…]
$\|\bm{m}-\bm{m}^{\prime}\|_{2}\leq 2\left(\tfrac{\Delta}{c_{K}}\right)^{2K-1}\,.$
Letting $\{{\bm{\theta}}_{1},\dots,{\bm{\theta}}_{K}\}$ and $\{{\bm{\theta}}_{1}^{\prime},\dots,{\bm{\theta}}_{K}^{\prime}\}$ denote the support of $\rho$ and $\rho^{\prime}$ respectively, we obtain
$\min_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2}\geq\Delta\geq\tfrac{1}{2}c_{K}^{2K-1}\Delta^{-2K+2}\|\bm{m}-\bm{m}^{\prime}\|_{2}\,,$
as desired. ∎
C.5 Proof of Remark 8
Proof.
To see why the remark holds, it is enough to consider $v=\mathbf{e}_{1}$, and suppose that there existed such a function $s_{1}$. To lighten notation in this argument, we let ${\bm{\theta}}_{k}:={\bm{\theta}}_{k}^{*}$, for all $k$. Using (2), and the definition of $m_{1}$ in (29), if equality held throughout in $m_{1}({\bm{\omega}}^{*})=\bar{m}_{1}({\bm{\omega}}^{*})$, then with $A_{{{\bm{\theta}}}_{k}}(x_{j})$ denoting $A_{{{\bm{\theta}}}_{k}}(x_{j}\mid x_{1},\ldots,x_{p})$, since $x_{1},\ldots,x_{p}$ are non-random, we have
$\sum_{k=1}^{K}\alpha_{k}\left[\sum_{j=1}^{p}A_{{{\bm{\theta}}}_{k}}(x_{j})s_{1}(x_{j})\right]=\sum_{k=1}^{K}\alpha_{k}\theta_{1k}.$
Let us write $\beta_{j}:=s_{1}(x_{j})$ for $j\in[p]$. Under the softmax parametrization (1), these quantities therefore satisfy
$\frac{\sum_{j=1}^{p}\exp({x_{j}^{\top}{{\bm{\theta}}_{k}}})\beta_{j}}{\sum_{\ell=1}^{p}\exp({x_{\ell}^{\top}{{\bm{\theta}}_{k}}})}$ $=\theta_{1k},\quad\quad\forall{{\bm{\theta}}_{k}}\in\mathbb{R}^{L}\,,$
or, differentiating in ${{\bm{\theta}}_{k}}$,
$\frac{\sum_{j=1}^{p}\exp({x_{j}^{\top}{{\bm{\theta}}_{k}}})x_{j}\beta_{j}}{\sum_{\ell=1}^{p}\exp({x_{\ell}^{\top}{{\bm{\theta}}_{k}}})}-\frac{\sum_{j=1}^{p}\exp({x_{j}^{\top}{{\bm{\theta}}_{k}}})\beta_{j}}{\sum_{\ell=1}^{p}\exp({x_{\ell}^{\top}{{\bm{\theta}}_{k}}})}\frac{\sum_{j=1}^{p}\exp({x_{j}^{\top}{{\bm{\theta}}_{k}}})x_{j}}{\sum_{\ell=1}^{p}\exp({x_{\ell}^{\top}{{\bm{\theta}}_{k}}})}=\mathbf{e}_{1},\quad\forall{{\bm{\theta}}_{k}}\in\mathbb{R}^{L}.$ (119)
Now, let $\mathcal{C}$ be the convex hull of ${x}_{1},\dots,{x}_{p}$. This is a nonempty polytope in $\mathbb{R}^{L}$. Assume without loss of generality that $x_{1}$ is an extreme point of $\mathcal{C}$, and let $\bm{a}\in\mathbb{R}^{L}$ be any vector in the interior of the normal cone of $\mathcal{C}$ at $x_{1}$. For any real numbers $\lambda_{1},\dots,\lambda_{p}$, it holds that
$\lim_{t\to\infty}\frac{\sum_{j=1}^{p}\exp\left({x_{j}^{\top}{(t\bm{a})}}\right)\lambda_{j}}{\sum_{\ell=1}^{p}\exp\left({x_{\ell}^{\top}{(t\bm{a})}}\right)}=\frac{\sum_{j:x_{j}=x_{1}}\lambda_{j}}{|{j:x_{j}=x_{1}}|}\,.$
Therefore, choosing ${{\bm{\theta}}_{k}}=tx_{1}$ in (119) and taking the limit $t\to\infty$ on both sides yields
$\bm{0}=x_{1}\frac{\sum_{j:x_{j}=x_{1}}\beta_{j}}{|{j:x_{j}=x_{1}}|}-x_{1}\frac{\sum_{j:x_{j}=x_{1}}\beta_{j}}{|{j:x_{j}=x_{1}}|}=\mathbf{e}_{1}\,,$
a contradiction. ∎
C.6 Proof of Theorem 4
Proof.
The claim regarding (46) follows from the following theorem in conjunction with the union bounds argument over $r\leq 2K$ and $2\leq i\leq L$.
The claim for 1 follows from Lemma 17.
Finally, throughout the proofs for random features $X$ satisfying 7, we use the fact that the event
$\mathcal{E}_{2}=\left\{\max_{j\in[p]}\|X_{j}\|_{2}\leq\bar{\sigma}\left(\sqrt{L}+\sqrt{2(s+1)\log(p)}\right)\right\}$ (120)
holds with probability at least $1-p^{-s}$ for all $s\geq 2$. See, for instance, Lemma 30. This means that $\mathcal{E}_{2}$ holds $\mu$-almost surely by the Borel-Cantelli lemma. ∎
Theorem 6.
Grant $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$ and 3. Fix any $r\leq 2K$ and $2\leq i\leq L$. For any $\delta>0$ and any $s\geq 1$, the following holds for all $p\geq p_{0}(B,s,\delta)$.
  • (1)
    For any fixed $v\in\mathbb{S}^{L-1}$, with probability at least $1-p^{-s}$,
    $\left|\bar{m}_{r}(v)-m_{r}(v)\right|$ $~\lesssim~r^{r/2}\sqrt{\log(p)\over p}+(r\log p)^{r/2}{\log(p)\over p^{1-\delta/2}},$
    $\left|\bar{m}_{r1;i}(v)-m_{r1;i}(v)\right|$ $~\lesssim~(r+1)^{(r+1)/2}\sqrt{\log(p)\over p}+[(r+1)\log p)]^{(r+1)/2}{\log(p)\over p^{1-\delta/2}}.$
  • (2)
    With probability at least $1-p^{-s}$, the following holds uniformly for all $v\in\mathbb{S}^{L-1}$:
    $\left|\bar{m}_{r}(v)-m_{r}(v)\right|$ $~\lesssim~r^{r/2}\sqrt{L\log(p)\over p}+[r(L+\log p)]^{r/2}{L\log(p)\over p^{1-\delta/2}},$
    $\left|\bar{m}_{r1;i}(v)-m_{r1;i}(v)\right|$ $~\lesssim~(r+1)^{(r+1)/2}\sqrt{L\log(p)\over p}+[(r+1)(L+\log p)]^{(r+1)/2}{L\log(p)\over p^{1-\delta/2}}.$
Proof.
We only prove the uniform convergence result in part (2) as the result for fixed $v$ in part (1) follows immediately by setting $L=1$.
Fix any $r\in[2K]$. We bound $\sup_{v\in\mathbb{S}^{L-1}}|\bar{m}_{r}(v)-m_{r}(v)|$. Recall from (140) that
$\bar{g}_{r,v}(X;{\bm{\theta}}_{k}^{*})=H_{r}(X^{\top}v)\exp(X^{\top}{\bm{\theta}}_{k}^{*}).$
Note that
$\mathbb{E}[H_{r}(X^{\top}v)\exp(X^{\top}{\bm{\theta}})]=({\bm{\theta}}^{\top}v)^{r}~\mathbb{E}[\exp(X^{\top}{\bm{\theta}})]$ (121)
which together with (62) ensures that
$({\bm{\theta}}_{k}^{*\top}v)^{r}={\mathbb{E}[\bar{g}_{r,v}(X;{\bm{\theta}}_{k}^{*})]\over\mathbb{E}[\exp(X^{\top}{\bm{\theta}}_{k}^{*})]}={p\mathbb{E}[\bar{g}_{r,v}(X;{\bm{\theta}}_{k}^{*})]\over\mathbb{E}[N_{{\bm{\theta}}_{k}^{*}}]}.$ (122)
We find that $\bar{m}_{r}(v)-m_{r}(v)$ equals to
$\sum_{k=1}^{K}\alpha_{k}^{*}\left[{\sum_{j=1}^{p}H_{r}(X_{j}^{\top}v)\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})\over N_{{\bm{\theta}}_{k}^{*}}}-({\bm{\theta}}_{k}^{*\top}v)^{r}\right]$
$=\sum_{k=1}^{K}\alpha_{k}^{*}\left[{\sum_{j=1}^{p}\bar{g}_{r,v}(X_{j};{\bm{\theta}}_{k}^{*})\over N_{{\bm{\theta}}_{k}^{*}}}-{p\mathbb{E}[\bar{g}_{r,v}(X_{j};{\bm{\theta}}_{k}^{*})]\over\mathbb{E}[N_{{\bm{\theta}}_{k}^{*}}]}\right]$ by (122)
[…]
so that it remains to bound from above
$\max_{k\in[K]}{|\sum_{j=1}^{p}\bar{g}_{r,v}(X_{j};{\bm{\theta}}_{k}^{*})-p\mathbb{E}[\bar{g}_{r,v}(X_{j};{\bm{\theta}}_{k}^{*})]|\over N_{{\bm{\theta}}_{k}^{*}}}+\max_{k\in[K]}\|{\bm{\theta}}_{k}^{*}\|_{2}^{r}{|\mathbb{E}[N_{{\bm{\theta}}_{k}^{*}}]-N_{{\bm{\theta}}_{k}^{*}}|\over N_{{\bm{\theta}}_{k}^{*}}}$
Invoking Lemma 21, Lemma 16 and Lemma 14 and taking union bounds over $k\in[K]$ give that
$\sup_{v\in\mathbb{S}^{L-1}}|\bar{m}_{r}(v)-m_{r}(v)|\lesssim r^{r/2}\sqrt{L\log(p)\over p}+[r(L+\log p)]^{r/2}{L\log(p)\over p^{1-\delta/2}}$ (123)
with probability at least $1-p^{-s}$.
Since the same argument applies to prove the bounds for the errors of the mixed-moments, we omit the proof. ∎
C.7 Extension to $\mu=\mathcal{N}_{L}(0,\Sigma)$
When $X\sim\mathcal{N}_{L}(0,\Sigma)$, one can still use
$h_{r}(X)$ $=H_{r}(X^{\top}v),$ (124)
$h_{r1;i}(X)$ $=H_{r}(X^{\top}v)(X^{\top}w_{i})\,.$ (125)
Let $U:=\Sigma^{-1/2}X\sim\mu_{0}=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. Then, for any generic ${\bm{\theta}}\in\mathbb{R}^{L}$, and given $v\in\mathbb{R}^{L}$
$\mathbb{E}_{\mu}\left[H_{r}(X^{\top}v)\exp(X^{\top}{{\bm{\theta}}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{{\bm{\theta}}})\right]$ $=$ ${\mathbb{E}_{\mu_{0}}\left[H_{r}(U^{\top}u)\exp(U^{\top}{\bar{{\bm{\theta}}}})\right]\over\mathbb{E}_{\mu_{0}}\left[\exp(U^{\top}{\bar{{\bm{\theta}}}})\right]}\ =\ (u^{\top}\bar{{{\bm{\theta}}}})^{r}\ =\ (v^{\top}\Sigma{{\bm{\theta}}})^{r}$ (126)
with $u:=\Sigma^{1/2}v$ and $\bar{{{\bm{\theta}}}}:=\Sigma^{1/2}{{\bm{\theta}}}$, where the second equality holds by Lemma 11, by the construction of $h_{r}$, since $U$ is a standard Gaussian on $\mathbb{R}^{L}$. Thus, if the procedure of Section 3.3 is applied relative to functions given by (124), but $X\sim\mathcal{N}_{L}(0,\Sigma)$, then $\bar{{\bm{\theta}}}_{k}$ approximates $\Sigma{{\bm{\theta}}}_{k}^{*}$, for each $k\in[K]$. One immediately has
$\|\Sigma^{-1}\bar{{\bm{\theta}}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq{1\over\underline{\sigma}^{2}}\|\bar{{\bm{\theta}}}_{k}-\Sigma{\bm{\theta}}_{k}^{*}\|_{2}.$
C.8 Proof of Theorem 5
Proof.
From Proposition 2, it suffices to show
$\max_{r<2K}|\widehat{m}_{r}-\bar{m}_{r}|+\max_{r<K,2\leq i\leq L}|\widehat{m}_{r1;i}-\bar{m}_{r1:i}|=\mathcal{O}_{\mathbb{P}}(\sqrt{\log(L)/N}).$
For every fixed integer $r<2K$, by an application of Chebyshev’s inequality, for any $t>0$,
$\mathbb{P}\left(|\widehat{m}_{r}-\bar{m}_{r}|\geq\frac{t}{\sqrt{N}}\mid\mathbf{X}\right)$$\leq$ $\frac{\mathbb{E}[h_{r}^{2}(Y)\mid\mathbf{X}]}{t^{2}}.$
Since the quantity
$\mathbb{E}\left[h_{r}^{2}(Y)\mid\mathbf{X}\right]$ $=$ $\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}h_{r}^{2}(X_{j})\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})\over{1\over p}\sum_{i=1}^{p}\exp(X_{i}^{\top}{\bm{\theta}}_{k}^{*})}$
has $\mu$ a.s. limit, by taking the union bounds over $r<2K$ with $K=\mathcal{O}(1)$, we conclude that for large $p$,
$\max_{r<2K}|\widehat{m}_{r}-\bar{m}_{r}|=\mathcal{O}_{\mathbb{P}}(\sqrt{\log(K)/N}).$
Similar arguments can be used to bound $\max_{r<K,2\leq i\leq L}|\widehat{m}_{r1;i}-\bar{m}_{r1:i}|$. ∎
Remark 11.
[…]
In particular, the proof of Proposition 2 reveals that its conclusion holds when each atom of $\widehat{\rho}_{1}$ and $\rho_{1}^{*}$ is at least $\Delta_{1}/2$ away from all but $\ell^{\prime}=1$ other atom (itself). However, if $\ell^{\prime}>1$, Proposition 4 in Wu and Yang (2020) shows that we cannot expect a parametric rate in the estimation of ${\bm{\theta}}_{k}$, even in one dimension, as display (109) in the proof then becomes
$W_{1}(\widehat{\rho}_{1},\rho_{1}^{*})\leq 2K\left({2K4^{2K-1}2^{2K-\ell^{\prime}-1}\over\Delta_{1}^{2K-\ell^{\prime}-1}}\right)^{1\over\ell^{\prime}}\|\widehat{\bm{m}}-{\bm{m}}\|_{2}^{1\over\ell^{\prime}},$ (127)
a rate that will be inherited by $|\theta_{1k}^{*}-\widehat{\theta}_{1k}|$, for each $k$, via (105). In the worst case, when $\ell^{\prime}=2K-1$, we obtain $W_{1}(\widehat{\rho}_{1},\rho_{1})\lesssim K\|\widehat{\bm{m}}-\bm{m}\|_{2}^{1/(2K-1)},$ by Proposition 1 in Wu and Yang (2020). Thus, although consistent estimation of the softmax mixture parameters will continue to hold when the atoms are distinct, but not well separated, neither the estimation of ${\bm{\theta}}_{k}^{*}$ nor that of ${\bm{\alpha}}^{*}$ can be expected to follow a parametric decay rate. This is confirmed by our simulation results in Section 4.
C.9 Proof of Proposition 3
Proof.
By definition, it suffices to bound
$\left\|\widehat{\Gamma}-\bar{\Gamma}\right\|_{\mathrm{op}}=\left\|{1\over N}\sum_{\ell=1}^{N}W_{\ell}\right\|_{\mathrm{op}},$
where we write
$W_{\ell}:={\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}-\mathbb{E}\left[{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\mid\mathbf{X}\right]\in\mathbb{R}^{L\times L}.$
To invoke the matrix-valued Bernstein’s inequality in Lemma 31, note that, by using (55),
$\max_{\ell\in[N]}\left\|W_{\ell}\right\|_{\mathrm{op}}\leq 2\max_{\ell\in[N]}\left\|{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\right\|_{\mathrm{op}}\leq 2C\|\mathbf{X}\|_{\infty,2}^{2}.$ (128)
while
$\left\|\sum_{\ell=1}^{N}\mathbb{E}[W_{\ell}^{2}]\right\|_{\mathrm{op}}$ $\leq N\left\|\mathbb{E}\left[{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\mid\mathbf{X}\right]\right\|_{\mathrm{op}}$
$\leq N\left\|\mathbb{E}\left[{\nabla^{2}\mu(Y_{\ell})\over\mu(Y_{\ell})}\mid\mathbf{X}\right]\right\|_{\mathrm{op}}C\|\mathbf{X}\|_{\infty,2}^{2}$ by (128)
$\leq N\left\|\bar{\Gamma}\right\|_{\mathrm{op}}C\|\mathbf{X}\|_{\infty,2}^{2}$ by (53)
$\leq N\left(\|\bar{\Gamma}-\Gamma\|_{\mathrm{op}}+\|\Gamma\|_{\mathrm{op}}\right)C\|\mathbf{X}\|_{\infty,2}^{2}$
[…]
Invoking Lemma 31 with $\sigma^{2}=C^{\prime}N\|\mathbf{X}\|_{\infty,2}^{2}$, $U=2C\|\mathbf{X}\|_{\infty,2}^{2}$ and $t=C^{\prime\prime}\sqrt{N\log(N)}\|\mathbf{X}\|_{\infty,2}$ yields that, on the event $\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})$,
$\|\widehat{\Gamma}-\bar{\Gamma}\|_{\mathrm{op}}\lesssim\|\mathbf{X}\|_{\infty,2}\sqrt{\log N\over N},$
with probability at least $1-14\exp\left(-C^{\prime\prime}\log(N)+\log(L)\right).$ The proof is complete. ∎
C.10 Proof of 3
Proof.
For the case $\Sigma=\mathbf{I}_{L}$, we have
$\|\bar{\Gamma}-\Gamma\|_{\mathrm{op}}=\sup_{v\in\mathbb{S}^{L-1}}|\bar{m}_{r}(v)-m_{r}(v)|$
so that invoking part (2) of Theorem 6 with $r=2$ gives that for any $\delta>0$, $s\geq 1$ and $p\geq p_{0}(B,\delta,s)$, $\mathbb{P}\{\mathcal{E}_{\Gamma}(\epsilon_{p}^{\prime})\}\geq 1-p^{-s}$ with
$\epsilon_{p}^{\prime}\lesssim\sqrt{L\log(p)\over p}+{(L+\log p)L\log(p)\over p^{1-\delta/2}}.$
The claim thus follows by recalling $\mathcal{E}_{2}$ in (120).
For the general case $\mathcal{N}_{L}(0,\Sigma)$, it is easy to see that $\widehat{\Gamma}$ and $\Gamma$ are rescaled version of their counterparts for $\Sigma=\mathbf{I}_{L}$ (written as $\widehat{\Gamma}_{0}$ and $\Gamma_{0}$) in the sense that
$\widehat{\Gamma}-\Gamma=\Sigma^{-1/2}(\widehat{\Gamma}_{0}-\Gamma_{0})\Sigma^{-1/2}.$
The claim thus follows immediately. ∎
C.11 Proof of Lemma 2
Proof.
Pick any $k\neq k^{\prime}$. We first bound from below
$|v^{\top}{\bm{\theta}}_{k}^{*}-v^{\top}{\bm{\theta}}_{k^{\prime}}^{*}|={|({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}\widehat{V}\widehat{V}^{\top}u|\over\|\widehat{V}\widehat{V}^{\top}u\|_{2}}={|({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}\widehat{V}(\widehat{V}^{\top}u)|\over\|\widehat{V}^{\top}u\|_{2}}.$
Since, conditioning on $\widehat{V}$, $\widehat{V}^{\top}u\sim\mathcal{N}_{K}(0,\mathbf{I}_{K})$ so that $\widehat{V}^{\top}u/\|\widehat{V}^{\top}u\|_{2}$ is uniformly distributed over $\mathbb{S}^{K-1}$, invoking Lemma 32 gives that for all $t\geq 0$,
$\mathbb{P}\left\{|v^{\top}{\bm{\theta}}_{k}^{*}-v^{\top}{\bm{\theta}}_{k^{\prime}}^{*}|<\|\widehat{V}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})\|_{2}~t\right\}<t\sqrt{K}.$
To bound from above $\|\widehat{V}^{\top}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})\|_{2}$, recall that $\widehat{V}\in\mathbb{O}_{L\times K}$ denotes the left leading eigenvectors of $\widehat{\Gamma}$. It then follows that
$\|\widehat{V}^{\top}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})\|_{2}^{2}$ $=({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}\widehat{V}\widehat{V}^{\top}({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})$
$=\|{\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*}\|_{2}^{2}-({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})({\bm{\theta}}_{k}^{*}-{\bm{\theta}}_{k^{\prime}}^{*})$
$\geq\Delta^{2}-2{\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}-2{\bm{\theta}}_{k^{\prime}}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k^{\prime}}^{*}.$
Notice that
${\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}$ $=\sup_{u\in\mathbb{S}^{L-1}}u^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})u$
$\leq{1\over\underline{\alpha}}\sup_{u\in\mathbb{S}^{L-1}}u^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})\sum_{a=1}^{K}\alpha_{a}^{*}{\bm{\theta}}_{a}^{*}{\bm{\theta}}_{a}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})u$
$\leq{1\over\underline{\alpha}}\sup_{u\in\mathbb{S}^{L-1}}u^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})~\widehat{\Gamma}~(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})u$
$\qquad+{1\over\underline{\alpha}}\sup_{u\in\mathbb{S}^{L-1}}u^{\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})(\widehat{\Gamma}-\Gamma)(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top})u$
$\leq{1\over\underline{\alpha}}\left(\lambda_{K+1}(\widehat{\Gamma})+\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}\right)$
$\leq{2\over\underline{\alpha}}\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}.$
The last step uses Weyl’s inequality and $\lambda_{K+1}(\Gamma)=0$. We write $\lambda_{1}(Q)\geq\lambda_{2}(Q)\geq\cdots\geq\lambda_{d}(Q)$ as the non-increasing eigenvalues of any $d\times d$ symmetric matrix $Q$.
Therefore, since the event $C^{\prime}\|\mathbf{X}\|_{\infty,2}\sqrt{\log N/N}+\epsilon_{p}^{\prime}\leq\underline{\alpha}\Delta^{2}$ and Proposition 3 imply
$\Delta^{2}\underline{\alpha}\geq 8\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}},$ (129)
we obtain that for all $t\geq 0$,
$\mathbb{P}\left\{|v^{\top}{\bm{\theta}}_{k}^{*}-v^{\top}{\bm{\theta}}_{k^{\prime}}^{*}|<{t\Delta\over 2\sqrt{K}}\right\}<t.$
By taking the union bounds over $k,k^{\prime}\in[K]$ with $k\neq k^{\prime}$, the proof is complete. ∎
C.12 Proof of Lemma 3
We first give a proof of Eq. 61 for completeness.
Proof of Eq. 61.
Fix ${\bm{\theta}}^{*}\in\mathbb{S}^{L-1}$. For arbitrary $\delta\in[0,1)$, Tkocz (2012) gives that for any $v$ uniformly drawn from $\mathbb{S}^{L-1}$,
$\mathbb{P}\left\{v^{\top}{\bm{\theta}}^{*}\geq\delta\right\}\leq e^{-L\delta^{2}/2}.$
It thus implies that, for any $\varepsilon\in(0,1)$,
$\mathbb{P}\left\{v_{i}^{\top}{\bm{\theta}}^{*}<\delta,~\text{for all }i\in[m]\right\}$ $~\leq~\left(1-e^{-L\delta^{2}/2}\right)^{m}$
$~\leq~e^{-me^{-L\delta^{2}/2}}$ $\text{by }1-x\leq e^{-x}$
$~\leq~\varepsilon$
provided that
$m\geq\exp(L\delta^{2}/2)\log(1/\varepsilon).$
Recall that for any $v,{\bm{\theta}}^{*}\in\mathbb{S}^{L-1}$
$\|v-{\bm{\theta}}^{*}\|_{2}\leq\delta_{0}\quad\iff\quad v^{\top}{\bm{\theta}}^{*}\geq 1-{\delta_{0}^{2}\over 2}.$
The proof is complete by taking $\delta=1-\delta_{0}^{2}/2$ above and using $(1-\delta_{0}^{2}/2)^{2}\leq 2-\delta_{0}^{2}$. ∎
Proof of Lemma 3.
Recall that ${\bm{\theta}}_{k}^{*}\in\mathbb{S}^{L-1}$. For any $i\in[m]$ with $v_{i}$ defined in (60), one has
$\|v_{i}-{\bm{\theta}}_{k}^{*}\|_{2}^{2}$ $=2-2v_{i}^{\top}{\bm{\theta}}_{k}^{*}$
$=2-{2u_{i}^{\top}\widehat{V}\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}u_{i}\|_{2}}$
$=2-{2u_{i}^{\top}\widehat{V}\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}u_{i}\|_{2}\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}}+{2u_{i}^{\top}\widehat{V}\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}u_{i}\|_{2}\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}}(1-\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2})$
$\leq\left\|{\widehat{V}^{\top}u_{i}\over\|\widehat{V}^{\top}u_{i}\|_{2}}-{\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\over\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}}\right\|_{2}^{2}+2\left(1-\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}\right).$
Using the arguments in the proof of Proposition 3, one has
$1-\|\widehat{V}^{\top}{\bm{\theta}}_{k}^{*}\|_{2}^{2}={\bm{\theta}}_{k}^{*\top}(\mathbf{I}_{L}-\widehat{V}\widehat{V}^{\top}){\bm{\theta}}_{k}^{*}\leq{2\over\underline{\alpha}}\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}:=\delta_{\Gamma}.$
The proof follows by applying (61) to the first term with $\delta^{2}=\delta_{0}^{2}-2\delta_{\Gamma}$. ∎
Appendix D Concentration inequalities for quantities related with the random embedding matrix
The following subsections contain deviation inequalities between $N_{\bm{\theta}}$, $\mathrm{I}_{\bm{\theta}}$, $\mathrm{II}_{\bm{\theta}}$, $H_{\bm{\theta}}$ in Eqs. 62 and 63 and their corresponding expectations, derived under the following distributional assumption on the rows of $\mathbf{X}$.
Assumption 7.
The rows of $\mathbf{X}$ are i.i.d. sub-Gaussian random vectors with zero mean and sub-Gaussian constant $\overline{\sigma}<\infty$. [^2] A random vector $Z\in\mathbb{R}^{d}$ is said to be $\gamma$ sub-Gaussian if for any $v\in\mathbb{R}^{d}$, $\mathbb{E}[\exp(v^{\top}Z)]\leq\exp(\|v\|_{2}^{2}\gamma^{2}/2)$.
D.1 Concentration inequalities related with $\mathrm{I}_{\theta}$, $\mathrm{II}_{\theta}$ and $N_{\theta}$
Lemma 12.
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $s\geq 2$ be arbitrary and assume $p$ is large enough such that
$c_{\mathsf{Bern}}\cdot p>L\log(7)+s\log(p)$
where $c_{\mathsf{Bern}}$ is the universal constant appearing in Bernstein’s exponential inequality for sums of independent sub-exponential random variables. We have, with probability $1-4p^{1-s}$,
$\left\|\mathrm{I}_{\bm{\theta}}-\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right\|_{\mathrm{op}}$ $\lesssim\overline{\sigma}^{2}\exp(2\overline{\sigma}^{2}B)p^{1-s/2}+\overline{\sigma}^{2}p^{{1\over 2}+\overline{\sigma}B\sqrt{2s/\log p}}\sqrt{L\log(7)+s\log(p)}$
In particular, for arbitrary $\delta>0$ and for $p$ large enough such that
$\delta^{2}\log p\geq 2s\overline{\sigma}^{2}B^{2},$
we have
$\mathbb{P}\left\{\left\|\mathrm{I}_{\bm{\theta}}-\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}p^{{1\over 2}+\delta}\sqrt{L\log(7)+s\log(p)}\right\}\geq 1-4p^{1-s}.$
Proof.
By definition and a standard discretization argument (see, for instance, Vershynin (2018))
$\left\|\mathrm{I}_{\bm{\theta}}-\mathbb{E}\left[\mathrm{I}_{\bm{\theta}}\right]\right\|_{\mathrm{op}}$ $=\sup_{v\in\mathbb{S}^{L-1}}v^{\top}\left(\mathrm{I}_{\bm{\theta}}-\mathbb{E}\left[\mathrm{I}_{\bm{\theta}}\right]\right)v$
$=\sup_{v\in\mathbb{S}^{L-1}}\sum_{j=1}^{p}\left\{(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}-\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}\right]\right\}$
$\leq 3\max_{v\in\mathcal{N}_{L}(1/3)}\sum_{j=1}^{p}\left\{(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}-\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}\right]\right\}$
Here, $\mathcal{N}_{L}(1/3)$ is a (1/3)(1/3)( 1 / 3 )-net of $\mathbb{S}^{L-1}$ and satisfies $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$ (see, for instance, Vershynin (2018)). Next, we use a truncation device. For fixed ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, the random variables $X_{j}^{\top}{\bm{\theta}}$ are zero mean sub-Gaussian random variables with sub-Gaussian constant no greater than $\|{\bm{\theta}}\|_{2}^{2}\,\overline{\sigma}^{2}\leq B^{2}\overline{\sigma}^{2}$. Consequently, the events
$\mathcal{X}_{j}(s,{\bm{\theta}})=\left\{|X_{j}^{\top}{\bm{\theta}}|\leq\overline{\sigma}B\sqrt{2s\log(p)})\right\}$ (130)
have probabilities
$\mathbb{P}\left(\mathcal{X}_{j}(s,{\bm{\theta}})\right)\geq 1-2p^{-s}.$ (131)
Clearly,
$\left|\sum_{j=1}^{p}\left\{(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}-\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}\right]\right\}\right|$
$\leq\left|\sum_{j=1}^{p}\left\{(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}-\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}\right]\right\}\right|$
$+\left|\sum_{j=1}^{p}\left\{(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}-\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}\right]\right\}\right|$
On the event $\cap_{j\in[p]}\mathcal{X}_{j}(s,{\bm{\theta}})$, which holds with probability at least $1-2p^{1-s}$, we have
$\sum_{j=1}^{p}(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}=0$
by definition, while
$\sum_{j=1}^{p}\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}\right]$ $\leq\sum_{j=1}^{p}\sqrt{\mathbb{E}\left[(X_{j}^{\top}v)^{4}e^{2X_{j}^{\top}{\bm{\theta}}}\right]}\sqrt{\mathbb{P}(\mathcal{X}_{j}^{c}(s,{\bm{\theta}}))}$
$\lesssim p\overline{\sigma}^{2}e^{2\overline{\sigma}^{2}B^{2}}2p^{-s/2}$ by Lemma 25 and (131)
Since $X_{j}^{\top}v$ is $\overline{\sigma}$ sub-Gaussian, the distribution of $(X_{j}^{\top}v)^{2}$ is sub-exponential (with parameter $\leq\overline{\sigma}^{2}$). This implies that the distribution of
$W_{j}(s,{\bm{\theta}},v):=(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j(s,{\bm{\theta}})}}$
is sub-exponential, with parameter
$\|W_{j}(s,{\bm{\theta}},v)\|_{\psi_{1}}\leq\overline{\sigma}^{2}p^{\overline{\sigma}B\sqrt{2s/\log p}}:=\kappa$
Bernstein’s inequality for sums of independent sub-exponential random variables, see (Vershynin, 2018, Section 2.8.1), states that, for some numerical constant $c>0$ and any $t\geq 0$,
$\mathbb{P}\left\{\sum_{j=1}^{p}\left(W_{j}(s,{\bm{\theta}},v)-\mathbb{E}\left[W_{j}(s,{\bm{\theta}},v)\right]\right)\geq p\kappa t\right\}\leq 2\exp\left\{-cp\min\left(t,t^{2}\right)\right\}$
We choose $t^{2}=c^{-1}(L\log(7)+s\log(p))/p<1$, and we conclude, using the union bound over $v\in\mathcal{N}_{L}(1/3)$,
$\max_{v\in\mathcal{N}_{L}(1/3)}\sum_{j=1}^{p}\left(W_{j}(s,{\bm{\theta}},v)-\mathbb{E}\left[W_{j}(s,{\bm{\theta}},v)\right]\right)\lesssim\overline{\sigma}^{2}p^{\overline{\sigma}B\sqrt{2s/\log p}}\sqrt{pL\log(7)+sp\log(p)}$ (132)
with probability at least
$1-2\cdot 7^{L}\exp\left\{-cp\min\left(t,t^{2}\right)\right\}\geq 1-2p^{-s}.$
The proof is complete. ∎
Lemma 13.
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $\delta>0$ and $s\geq 2$. For $p\geq p_{0}=p_{0}(B,\delta,s,\overline{\sigma})$, we have
$\mathbb{P}\left\{\left\|\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\right\|_{2}\lesssim\overline{\sigma}\sqrt{L+\log(p)}~p^{\frac{1}{2}+\delta}\right\}\geq 1-4p^{1-s}.$
Proof.
We use the same arguments to prove Lemma 12. Again, the standard discretization argument ensures that
$\left\|\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\right\|_{2}\leq 2\max_{v\in\mathcal{N}_{L}(1/3)}v^{\top}\left(\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\right).$
Fix any $v\in\mathcal{N}_{L}(1/3)$ and observe that
$v^{\top}\left(\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\right)$ $\leq\sum_{j=1}^{p}\left\{e^{X_{j}^{\top}{\bm{\theta}}}X_{j}^{\top}v1_{\mathcal{X}_{j}}-\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}X_{j}^{\top}v1_{\mathcal{X}_{j}}\right]\right\}+\sum_{j=1}^{p}\left|\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}X_{j}^{\top}v1_{\mathcal{X}_{j}^{c}}\right]\right|$
on the event $\cap_{j\in[p]}\mathcal{X}_{j}$ with the events $\mathcal{X}_{j}:=\mathcal{X}_{j}(s,{\bm{\theta}})$ defined in (130). The second term is no greater than
$p^{-s/2}\sum_{j=1}^{p}\sqrt{\mathbb{E}\left[e^{2X_{j}^{\top}{\bm{\theta}}}(X_{j}^{\top}v)^{2}\right]}\leq p^{1-s/2}\overline{\sigma}\left(1+\overline{\sigma}\|{\bm{\theta}}\|_{2}\right)e^{\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2}^{2}}$
by the Cauchy-Schwarz inequality, Lemma 24 and (131).
For the first term, we notice that $W_{j}(s,{\bm{\theta}},v)=e^{X_{j}^{\top}{\bm{\theta}}}X_{j}^{\top}v1_{\mathcal{X}_{j}(s,{\bm{\theta}},v)}$ is sub-Gaussian with sub-Gaussian parameter
$\|W_{j}(s,{\bm{\theta}},v)\|_{\psi_{2}}\leq\kappa^{\prime}=\overline{\sigma}\exp\left(\overline{\sigma}B\sqrt{2s\log p}\right)\leq p^{\delta}.$
Moreover, $\sum_{j=1}^{p}(W_{j}(s,{\bm{\theta}},v)-\mathbb{E}[W_{j}(s,{\bm{\theta}},v)])$ is sub-Gaussian with sub-Gaussian parameter $2\kappa^{\prime}\sqrt{p}$, whence, for all $t\geq 0$,
$\mathbb{P}\left\{\sum_{j=1}^{p}\left\{e^{X_{j}^{\top}{\bm{\theta}}}X_{j}^{\top}v1_{\mathcal{X}_{j}(s,{\bm{\theta}},v)}-\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}X_{j}^{\top}v1_{\mathcal{X}_{j}(s,{\bm{\theta}},v)}\right]\right\}\geq 2\kappa^{\prime}\sqrt{p}~t\right\}\leq 2\exp\left(-{t^{2}\over 2}\right).$ (133)
We take $t^{2}=C(L\log(7)+s\log(p))$ and take the union bound over $v\in\mathcal{N}_{L}(1/3)$ to complete the proof. ∎
Lemma 14.
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and $\delta>0$ and $s\geq 2$. For $p\geq p_{0}=p_{0}(B,\delta,s,\overline{\sigma})$, we have
$\mathbb{P}\left\{\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|\lesssim\sqrt{p^{1+\delta}\log(p)}\right\}\geq 1-4p^{1-s}.$
Proof.
Again, we follow the same arguments that we used to prove Lemmas 12 and 13. Recall the events $\{\mathcal{X}_{j}(s,{\bm{\theta}})\}_{j\in[p]}$ from (130). On the intersection of the events, we have
$|\displaystyle\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|$$\leq\left|\sum_{j=1}^{p}\left(e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}-\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}\right]\right)\right|+\sum_{j=1}^{p}\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}^{c}(s,{\bm{\theta}})}\right]$
$\leq\left|\sum_{j=1}^{p}\left(e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}-\mathbb{E}\left[e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}\right]\right)\right|+p^{1-s/2}e^{2\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2}^{2}}$
using Cauchy-Schwarz and (131). The first term on the right is a sum of independent bounded random variables, with
$\left|e^{X_{j}^{\top}{\bm{\theta}}}1_{\mathcal{X}_{j}(s,{\bm{\theta}})}\right|\leq\exp\left(\overline{\sigma}B\sqrt{2s\log p}\right)\lesssim p^{\delta}$
almost surely and the result follows easily from Hoeffding’s inequality. ∎
D.2 Concentration inequalities related with $H_{\theta}$ under Gaussianity
For any ${\bm{\theta}}\in\mathbb{R}^{L}$, the following lemma contains results on the moments related with $\mathrm{I}_{\bm{\theta}}$, $\mathrm{II}_{\bm{\theta}}$ and $N_{\bm{\theta}}$ under the condition
Assumption 8.
All the eigenvalues of $\Sigma$ belong to the fixed interval $[\underline{\sigma}^{2},\overline{\sigma}^{2}]\subset(0,\infty)$.
Lemma 15.
Grant 8. Let ${\bm{\theta}}\in\mathbb{R}^{L}$. Let $N_{\bm{\theta}}$, $\mathrm{I}_{\bm{\theta}}$ and $\mathrm{II}_{\bm{\theta}}$ be defined in (62). For any $v\in\mathbb{R}^{L-1}$, we have
$v^{\top}\mathbb{E}[\mathrm{II}_{\bm{\theta}}]=p(v^{\top}\Sigma{\bm{\theta}})e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2};$
$v^{\top}\mathbb{E}[\mathrm{I}_{\bm{\theta}}]v=p\left(v^{\top}\Sigma v+(v^{\top}\Sigma{\bm{\theta}})^{2}\right)e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2}.$
Moreover,
$\mathbb{E}[N_{\bm{\theta}}]=pe^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2},$
$\|\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}=p\|\Sigma{\bm{\theta}}\|_{2}e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2},$
$\lambda_{1}\left(\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right)=p\left(\lambda_{1}(\Sigma)+\|\Sigma{\bm{\theta}}\|_{2}^{2}\right)e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2},$
$\lambda_{L}\left(\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right)=p\lambda_{L}(\Sigma)e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2}.$
Proof.
Fix any $v\in\mathbb{R}^{L}$. By Lemma 24 with $\sigma^{2}=1$, $u=\Sigma^{1/2}v$ and ${\bm{\theta}}=\Sigma^{1/2}{\bm{\theta}}$, we have
$v^{\top}\mathbb{E}[\mathrm{II}_{\bm{\theta}}]=p\mathbb{E}\left[(v^{\top}X_{j})e^{X_{j}^{\top}{\bm{\theta}}}\right]=p(v^{\top}\Sigma{\bm{\theta}})e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2}$
and
$v^{\top}\mathbb{E}[\mathrm{I}_{\bm{\theta}}]v$ $=p\mathbb{E}\left[(X_{j}^{\top}v)^{2}e^{X_{j}^{\top}{\bm{\theta}}}\right]=p\left(v^{\top}\Sigma v+(v^{\top}\Sigma{\bm{\theta}})^{2}\right)e^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2}.$
Since
$\mathbb{E}[N_{\bm{\theta}}]=p\mathbb{E}[e^{X_{j}^{\top}{\bm{\theta}}}]=pe^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2},$
the other claims follow immediately from
$\|\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}=\sup_{v\in\mathbb{S}^{L-1}}v^{\top}\mathbb{E}[\mathrm{II}_{\bm{\theta}}]$
and
$\lambda_{1}\left(\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right)=\sup_{v\in\mathbb{S}^{L-1}}v^{\top}\mathbb{E}[\mathrm{I}_{\bm{\theta}}]v,\quad\lambda_{L}\left(\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right)=\inf_{v\in\mathbb{S}^{L-1}}v^{\top}\mathbb{E}[\mathrm{I}_{\bm{\theta}}]v.$
The proof is complete. ∎
The following lemma follows immediately from Lemmas 12, 13 & 14.
Lemma 16.
Grant 8. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, $s\geq 2$, $\delta>0$ and $\epsilon>0$. For $p\geq p_{0}(B,s,\overline{\sigma},\delta,\epsilon)$, the following holds with probability at least $1-4p^{1-s}$:
  1. (a)
    $(1-\epsilon)\mathbb{E}[N_{\bm{\theta}}]\leq N_{\bm{\theta}}\leq(1+\epsilon)\mathbb{E}[N_{\bm{\theta}}].$
  2. (b)
    $\|\mathrm{II}_{\bm{\theta}}\|_{2}\leq(1+\epsilon)p\overline{\sigma}^{2}Be^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2}$
  3. (c)
    $(1-\epsilon)\lambda_{L}\left(\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right)\leq\lambda_{L}\left(\mathrm{I}_{\bm{\theta}}\right)\leq\lambda_{1}\left(\mathrm{I}_{\bm{\theta}}\right)\leq(1+\epsilon)\lambda_{1}\left(\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\right).$
Proof.
Since $\mathbb{E}[N_{\bm{\theta}}]\geq p$ from Lemma 15, by invoking Lemma 14, the first result follows as
$N_{\bm{\theta}}\geq\mathbb{E}[N_{\bm{\theta}}]-\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|\geq\left(1-C\sqrt{\log(p)\over p^{1-\delta}}\right)\mathbb{E}[N_{\bm{\theta}}]$
with probability $1-4p^{1-s}$.
Part (b) follows by invoking Lemma 13 and Lemma 15.
Finally, by Weyl’s inequality, we have
$\left|\lambda_{k}(\mathrm{I}_{\bm{\theta}})-\lambda_{k}(\mathbb{E}[\mathrm{I}_{\bm{\theta}}])\right|\leq\|\mathrm{I}_{\bm{\theta}}-\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}.$
The last statement follows by invoking Lemma 12 and noting that $\lambda_{L}(\mathbb{E}[\mathrm{I}_{\bm{\theta}}])\geq p\lambda_{L}(\Sigma)$ from Lemma 15. ∎
The following lemma is a key result that provides deviation inequality of $\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}$ where
$\bar{H}_{\bm{\theta}}={\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\over\mathbb{E}[N_{\bm{\theta}}]}-{\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\mathbb{E}[\mathrm{II}_{\bm{\theta}}]^{\top}\over(\mathbb{E}[N_{\bm{\theta}}])^{2}}.$ (134)
Lemma 17.
Grant 8. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$ and fix any $s\geq 2$, $\delta>0$ and $\epsilon>0$. For $p\geq p_{0}(B,s,\overline{\sigma},\delta,\epsilon)$,
$\mathbb{P}\left\{\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}\right\}\geq 1-4p^{1-s}.$ (135)
Moreover, with the same probability as above, we have
$(1-\epsilon)\underline{\sigma}^{2}\leq\lambda_{L}(H_{\bm{\theta}})\leq\lambda_{1}(H_{\bm{\theta}})\leq(1+\epsilon)\overline{\sigma}^{2}.$
Proof.
By adding and subtracting terms, we first have
$\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}$ $\leq{\|\mathrm{I}_{\bm{\theta}}-\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}\over N_{\bm{\theta}}}+{\|\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}\over N_{\bm{\theta}}}{\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|\over\mathbb{E}[N_{\bm{\theta}}]}$
$\quad+{\|\mathrm{II}_{\bm{\theta}}(\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}])^{\top}\|_{\mathrm{op}}\over N_{\bm{\theta}}^{2}}+{\|(\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}])(\mathbb{E}[\mathrm{II}_{\bm{\theta}}])^{\top}\|_{\mathrm{op}}\over N_{\bm{\theta}}^{2}}$
$\quad+{\|\mathbb{E}[\mathrm{II}_{\bm{\theta}}](\mathbb{E}[\mathrm{II}_{\bm{\theta}}])^{\top}\|_{\mathrm{op}}\over N_{\bm{\theta}}^{2}}{(N_{\bm{\theta}}+\mathbb{E}[N_{\bm{\theta}}])\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|\over(\mathbb{E}[N_{\bm{\theta}}])^{2}}$
$\leq{\|\mathrm{I}_{\bm{\theta}}-\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}\over N_{\bm{\theta}}}+{\|\mathrm{II}_{\bm{\theta}}\|_{2}+\|\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}\over N_{\bm{\theta}}}{\|\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}\over N_{\bm{\theta}}}$
$\quad+\left[{\|\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}\over N_{\bm{\theta}}}+{\|\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}^{2}\over N_{\bm{\theta}}^{2}}{(N_{\bm{\theta}}+\mathbb{E}[N_{\bm{\theta}}])\over\mathbb{E}[N_{\bm{\theta}}]}\right]{\left|N_{\bm{\theta}}-\mathbb{E}[N_{\bm{\theta}}]\right|\over\mathbb{E}[N_{\bm{\theta}}]}$
Note that Lemma 16 ensures that
$N_{\bm{\theta}}\gtrsim\mathbb{E}[N_{\bm{\theta}}]\geq p,\qquad\|\mathrm{II}_{\bm{\theta}}\|_{2}\lesssim p\overline{\sigma}^{2}Be^{{\bm{\theta}}^{\top}\Sigma{\bm{\theta}}/2}$ (136)
with probability at least $1-4p^{1-s}$. In conjunction with Lemma 15, by invoking Lemma 12, Lemma 13 and Lemma 14 with $\sigma^{2}=\lambda_{1}(\Sigma)$, we conclude that
$\mathbb{P}\left\{\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}\right\}\geq 1-4p^{1-s},$
completing the proof of (135).
Regarding the second claim, observe that, for any $v\in\mathbb{S}^{L-1}$,
$v^{\top}\bar{H}_{\bm{\theta}}v$ $={\mathbb{E}[v^{\top}\mathrm{I}_{\bm{\theta}}v]\over\mathbb{E}[N_{\bm{\theta}}]}-{\mathbb{E}[v^{\top}\mathrm{II}_{\bm{\theta}}]\mathbb{E}[\mathrm{II}_{\bm{\theta}}^{\top}v]\over(\mathbb{E}[N_{\bm{\theta}}])^{2}}$
$=v^{\top}\Sigma v+(v^{\top}\Sigma{\bm{\theta}})^{2}-(v^{\top}\Sigma{\bm{\theta}})^{2}$ by Lemma 15
$=v^{\top}\Sigma v.$
The second result then follows by the definition of eigenvalues, (135) and Weyl’s inequality. ∎
D.3 Concentration inequalities related with $H_{\theta}$ under sub-Gaussianity
The following lemma bounds the moments of $\mathbb{E}[N_{\bm{\theta}}]$, $\mathbb{E}[\mathrm{II}_{\bm{\theta}}]$ and $\mathbb{E}[\mathrm{I}_{\bm{\theta}}]$ under 7.
Lemma 18.
Grant 7. For any ${\bm{\theta}}\in\mathbb{R}^{L}$, we have
$p\leq\mathbb{E}[N_{\bm{\theta}}]\leq pe^{\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2}/2}$ (137)
and
${\|\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}\over\mathbb{E}[N_{\bm{\theta}}]}\lesssim\overline{\sigma}+\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2},\qquad{\|\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}\over\mathbb{E}[N_{\bm{\theta}}]}\lesssim\overline{\sigma}^{2}+\overline{\sigma}^{4}\|{\bm{\theta}}\|_{2}^{2}.$
Proof.
The upper bound of $\mathbb{E}[N_{\bm{\theta}}]$ is easy to see and the lower bounds follows by Jensen’s inequality
$\mathbb{E}[N_{\bm{\theta}}]=p\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]\geq pe^{\mathbb{E}[X^{\top}{\bm{\theta}}]}=p.$
Regarding the other two results, fix any $v\in\mathbb{S}^{L-1}$. For arbitrary $t>0$, by using the sub-Gaussianity under 7, we have
$\mathbb{E}[v^{\top}\mathrm{I}_{\bm{\theta}}v]$ $=\mathbb{E}\left[(v^{\top}X)^{2}e^{X^{\top}{\bm{\theta}}}1\{|X^{\top}v|>t\}\right]+\mathbb{E}\left[(v^{\top}X)^{2}e^{X^{\top}{\bm{\theta}}}1\{|X^{\top}v|\leq t\}\right]$
$\leq\sqrt{\mathbb{E}[(v^{\top}X)^{4}e^{2X^{\top}{\bm{\theta}}}}~\sqrt{\mathbb{P}(|X^{\top}v|>t)}+t^{2}\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]$
$\leq\sqrt{\mathbb{E}[(v^{\top}X)^{4}e^{2X^{\top}{\bm{\theta}}}}~e^{-{t^{2}\over 4\overline{\sigma}^{2}}}+t^{2}\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]$
so that, by choosing $t^{2}=8\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2}^{2}$,
${\mathbb{E}[v^{\top}\mathrm{I}_{\bm{\theta}}v]\over\mathbb{E}[N_{\bm{\theta}}]}$ $\leq{\sqrt{\mathbb{E}[(v^{\top}X)^{4}e^{2X^{\top}{\bm{\theta}}}}\over\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]}~e^{-{t^{2}\over 4\overline{\sigma}^{2}}}+t^{2}$
$\lesssim{\overline{\sigma}^{2}e^{2\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2}^{2}-{t^{2}\over 4\overline{\sigma}^{2}}}\over\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]}+t^{2}$ by Lemma 25
[…]
Furthermore, we have
${v^{\top}\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\over\mathbb{E}[N_{\bm{\theta}}]}={\mathbb{E}[(v^{\top}X)e^{X^{\top}{\bm{\theta}}}]\over\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]}\leq\sqrt{\mathbb{E}[(v^{\top}X)^{2}e^{X^{\top}{\bm{\theta}}}]\over\mathbb{E}[e^{X^{\top}{\bm{\theta}}}]}\lesssim\overline{\sigma}+\overline{\sigma}^{2}\|{\bm{\theta}}\|_{2}.$
Since the above bounds hold for all $v$, the proof is complete. ∎
Similar as Lemma 17, we have the following result under 7.
Lemma 19.
Grant 7. Let ${\bm{\theta}}\in\mathbb{R}^{L}$ with $\|{\bm{\theta}}\|_{2}\leq B$, $s\geq 2$, $\delta>0$ and $\epsilon>0$. Assume $\lambda_{L}(\bar{H}_{\bm{\theta}})\geq\underline{\sigma}^{2}.$ Then for $p\geq p_{0}(B,s,\overline{\sigma},\underline{\sigma},\delta,\epsilon)$,
$\mathbb{P}\left\{\|H_{\bm{\theta}}-\bar{H}_{\bm{\theta}}\|_{\mathrm{op}}\lesssim\overline{\sigma}^{2}\sqrt{L+\log(p)\over p^{1-\delta}}\right\}\geq 1-4p^{1-s}.$ (138)
Moreover, for any $\epsilon>0$, with the same probability as above and some constant $C>1$, we have
$(1-\epsilon)\underline{\sigma}^{2}\leq\lambda_{L}(H_{\bm{\theta}})\leq\lambda_{1}(H_{\bm{\theta}})\leq(1+\epsilon)C\overline{\sigma}^{2}.$
Proof.
The proof of (138) is the same as that of Lemma 17 except that (136) is replaced by
$p\mathbb{E}[N_{\bm{\theta}}]\lesssim N_{\bm{\theta}}\lesssim p\mathbb{E}[N_{\bm{\theta}}]$
and
${\|\mathrm{II}_{\bm{\theta}}\|_{2}\over\mathbb{E}[N_{\bm{\theta}}]}\leq{\|\mathrm{II}_{\bm{\theta}}-\mathbb{E}[\mathrm{II}_{\bm{\theta}}]\|_{2}\over\mathbb{E}[N_{\bm{\theta}}]}+{\mathbb{E}[\|\mathrm{II}_{\bm{\theta}}\|_{2}]\over\mathbb{E}[N_{\bm{\theta}}]}\leq(1+\epsilon)(\overline{\sigma}+\overline{\sigma}^{2}B)$
by using Lemma 18. The second statement follows from Weyl’s inequality and noting that Lemma 18 implies
$\lambda_{1}(\bar{H}_{\bm{\theta}})\leq{\|\mathbb{E}[\mathrm{I}_{\bm{\theta}}]\|_{\mathrm{op}}\over\mathbb{E}[N_{\bm{\theta}}]}\lesssim\overline{\sigma}^{2}+\overline{\sigma}^{4}B^{2}.$
Appendix E Concentration inequalities related with Hermite polynomials
The following contains some concentration results related with Hermite polynomials. For any given ${\bm{\theta}}\in\mathbb{R}^{L}$, $r\in\mathbb{N}$ and $v\in\mathbb{S}^{L-1}$, define
$g_{r,v}(X_{j})$ $:=g_{r,v}(X_{j};{\bm{\theta}}):=H_{r}^{2}(X_{j}^{\top}v)\exp(X_{j}^{\top}{\bm{\theta}}),$ (139)
$\bar{g}_{r,v}(X_{j})$ $:=\bar{g}_{r,v}(X_{j};{\bm{\theta}}):=H_{r}(X_{j}^{\top}v)\exp(X_{j}^{\top}{\bm{\theta}})$ (140)
Lemma 20.
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ be any given $\|{\bm{\theta}}\|_{2}\leq B$ for some absolute constant $B$. Let $v\in\mathbb{S}^{L-1}$ be fixed. For any $r\in\mathbb{N}$, $\delta>0$, $r\geq 2$, we have for $p\geq p_{0}(s,\delta,\overline{\sigma},B)$
${1\over p}\sum_{j=1}^{p}\Bigl{(}g_{r,v}(X_{j})-\mathbb{E}[g_{r,v}(X_{j})]\Bigr{)}\lesssim r^{r}\sqrt{\log(p)\over p}+(r\log(p))^{r}{\log(p)\over p^{1-\delta}};$
${1\over p}\sum_{j=1}^{p}\Bigl{(}\bar{g}_{r,v}(X_{j})-\mathbb{E}[\bar{g}_{r,v}(X_{j})]\Bigr{)}\lesssim r^{r/2}\sqrt{\log(p)\over p}+(r\log(p))^{r/2}{\log(p)\over p^{1-\delta}}$
with probability at least $1-6p^{-s}$.
Proof.
We consider the event
$\mathcal{E}=\bigcap_{j=1}^{p}\left\{|X_{j}^{\top}{\bm{\theta}}|\leq\overline{\sigma}B\sqrt{2s\log p}\right\}\bigcap\left\{|X_{j}^{\top}v|\leq\sqrt{2s\log(p)}\right\}:=\bigcap_{j=1}^{p}\mathcal{E}_{j}$
Note that $\mathbb{P}(\mathcal{E})\geq 1-4p^{1-s}$ and on this event $\mathcal{E}$ we have
$g_{r,v}(X_{j})$ $\leq H_{r}^{2}(X_{j}^{\top}v)\exp(\overline{\sigma}B\sqrt{2s\log p})$ (141)
$\leq\left(C\sqrt{r\log(p)}\right)^{2r}p^{\delta}$
We use in the second step that the inequality $\overline{\sigma}B\sqrt{2s\log p}\leq\delta\log p$ holds for $p$ large enough, while we invoke Lemma 29 in the last step. Since
$\mathbb{E}[g_{r,v}(X_{j})1\{\mathcal{E}^{c}\}]$ $\leq\sqrt{\mathbb{E}[g_{r,v}^{2}(X_{j})]}\sqrt{\mathbb{P}(\mathcal{E}^{c})}$ by Cauchy-Schwarz
$\lesssim(B\sqrt{r})^{2r}\exp(\|{\bm{\theta}}\|_{2}^{2})~p^{-s/2}$ by Lemma 28. (142)
Next, we observe that, after invoking again Lemma 28,
$\mathbb{E}[g^{2}_{r,v}(X_{j})1\{\mathcal{E}_{j}\}]\leq\mathbb{E}[g^{2}_{r,v}(X_{j})]\lesssim(B\sqrt{r})^{4r}\exp(2B^{2}).$
Display (141) and an application of Bernstein’s inequality gives that, for any $t>0$, with probability at least $1-2e^{-t}$,
$\left|{1\over p}\sum_{j=1}^{p}\Bigl{(}g_{r,v}(X_{j})1\{\mathcal{E}_{j}\}-\mathbb{E}[g_{r,v}(X_{j})1\{\mathcal{E}_{j}\}]\Bigr{)}\right|\lesssim r^{r}\sqrt{t\over p}+(r\log(p))^{r}{t\over p^{1-\delta}}.$
Taking $t=s\log(p)$ and combining with the bound in (E) complete the proof of the first result.
The second result can be proved by the same arguments, and for this reason we omit its proof. ∎
The following lemma extends the results in Lemma 20 to uniform bounds over $v\in\mathbb{S}^{L-1}$.
Lemma 21.
Let ${\bm{\theta}}\in\mathbb{R}^{L}$ be any given $\|{\bm{\theta}}\|_{2}\leq B$ for some absolute constant $B$. For any $r\in\mathbb{N}$, $\delta>0$, $r\geq 2$, we have for $p\geq p_{0}(s,\delta,\overline{\sigma},B)$
$\sup_{v\in\mathbb{S}^{L-1}}{1\over p}\sum_{j=1}^{p}\Bigl{(}g_{r,v}(X_{j})-\mathbb{E}[g_{r,v}(X_{j})]\Bigr{)}\lesssim r^{r}\sqrt{L\log(p)\over p}+(L+\log(p))^{r}r^{r}{L\log(p)\over p^{1-\delta/2}}$
$\sup_{v\in\mathbb{S}^{L-1}}{1\over p}\sum_{j=1}^{p}\Bigl{(}\bar{g}_{r,v}(X_{j})-\mathbb{E}[\bar{g}_{r,v}(X_{j})]\Bigr{)}\lesssim r^{r/2}\sqrt{L\log(p)\over p}+(L+\log(p))^{r/2}r^{r/2}{L\log(p)\over p^{1-\delta/2}}$
with probability at least $1-6p^{-s}$.
Proof.
Again, we only prove the first claim. Using similar arguments in the above proof of Lemma 20, we have
$\mathbb{E}[g_{r,v}(X_{j})1\{\mathcal{X}_{j}^{c}\}]\lesssim p^{-s/2}(\sqrt{r})^{2r}.$
We aim to invoke Lemma 22 to bound
$\sup_{v\in\mathbb{S}^{L-1}}{1\over p}\sum_{j=1}^{p}\Bigl{(}g_{r,v}(X_{j})1\{\mathcal{X}_{j}\}-\mathbb{E}[g_{r,v}(X_{j})1\{\mathcal{X}_{j}\}]\Bigr{)}.$ (143)
To this end, we establish the order of $R_{2}$, $R_{1}$ and $L_{f}$ in (146) and (147). Let
$\bar{X}_{j}=X_{j}1\{\|X_{j}\|_{2}\leq B_{x}\},\qquad\text{with }B_{x}=2\overline{\sigma}\sqrt{L+s\log(p)}.$
Regarding $R_{2}$, we have
$\mathbb{E}\left[g^{2}_{r,v}(X_{j})1\{\mathcal{X}_{j}\}\right]$ $\leq\mathbb{E}\left[H_{r}^{4}(X_{j}^{\top}v)\exp(2X_{j}^{\top}{\bm{\theta}})\right]$
$\leq\left[(C\|{\bm{\theta}}\|_{2})^{4r}+(C\sqrt{r})^{4r}\right]\exp(2\|{\bm{\theta}}\|_{2}^{2}).$ by Lemma 28
$:=R_{2}.$
Regarding $R_{1}$, using Lemma 29 and $\bar{X}_{j}^{\top}v\leq B_{x}$, we find
$g_{r,v}(\bar{X}_{j})1\{\mathcal{X}_{j}\}\leq H_{r}^{2}(\bar{X}_{j}^{\top}v)\exp(X_{j}^{\top}{\bm{\theta}})1\{\mathcal{X}_{j}\}\leq p^{\delta}(C\sqrt{r})^{2r}B_{x}^{2r}:=R_{1}$
for $p\geq p_{0}(s,B,\delta,\overline{\sigma})$ large enough. Finally, for any $v,v^{\prime}\in\mathbb{S}^{L-1}$, and $p\geq p_{0}$,
$\left|g_{r,v}(\bar{X}_{j})1\{\mathcal{X}_{j}\}-g_{r,v^{\prime}}(\bar{X}_{j})1\{\mathcal{X}_{j}\}\right|$
$\leq\left|H_{r}^{2}(\bar{X}_{j}^{\top}v)-H_{r}^{2}(\bar{X}_{j}^{\top}v^{\prime})\right|p^{\delta}$
$\leq p^{\delta/2}\left(|H_{r}(\bar{X}_{j}^{\top}v)|+|H_{r}(\bar{X}_{j}^{\top}v^{\prime})|\right)\left|H_{r}(\bar{X}_{j}^{\top}v)-H_{r}(\bar{X}_{j}^{\top}v^{\prime})\right|$
$\leq p^{\delta}(C\sqrt{r})^{r}B^{r}\left|H_{r}(\bar{X}_{j}^{\top}v)-H_{r}(\bar{X}_{j}^{\top}v^{\prime})\right|.$
By definition, we have
$\left|H_{r}(\bar{X}_{j}^{\top}v)-H_{r}(\bar{X}_{j}^{\top}v^{\prime})\right|$ $\leq r!\sum_{j=0}^{\lfloor r/2\rfloor}{1\over 2^{j}j!(r-2j)!}\left|(\bar{X}_{j}^{\top}v)^{r-2j}-(\bar{X}_{j}^{\top}v^{\prime})^{r-2j}\right|$
$\leq r!\sum_{j=0}^{\lfloor r/2\rfloor}{r\choose 2j}{(2j)!\over 2^{j}j!}(r-2j)B_{x}^{r-2j}\|v-v^{\prime}\|_{2}$
$\leq B_{x}^{r}(C\sqrt{r})^{r}\|v-v^{\prime}\|_{2}$ (144)
The penultimate step uses the fact (see, the proof of Lemma A.3 of Doss et al. (2023))
$\left|(\bar{X}_{j}^{\top}v)^{\ell}-(\bar{X}_{j}^{\top}v^{\prime})^{\ell}\right|\leq\ell\|\bar{X}_{j}\|_{2}^{\ell}\|v-v^{\prime}\|_{2}.$
We can thus take
$L_{f}=(CB_{x}\sqrt{r})^{2r}p^{\delta}=R_{1}$
After we collect all pieces, and invoke Lemma 22 with $\epsilon=L/p$, $n=p$ and $d=L$, we find that, for any $\delta>0$, Eq. 143 is bounded from above by (in order)
$r^{r}\sqrt{L\log(p)\over p}+\left({L\log(p)\over p}\right)(L+\log(p))^{r}r^{r}p^{\delta}$
with probability at least $1-\mathcal{O}(p^{s})$. This concludes the proof of the first claim.
Regarding the second claim, we can essentially use the same arguments except for
$R_{2}$ $=\left[(C\|{\bm{\theta}}\|_{2})^{2r}+(C\sqrt{r})^{2r}\right]\exp(\|{\bm{\theta}}\|_{2}^{2}),$
$R_{1}$ $=p^{\delta/2}(C\sqrt{r})^{r}B_{x}^{r}=L_{f}.$
and we omit further details. ∎
The following technical lemma establishes a uniform rate of convergence for Lipschitz functions evaluated on sub-Gaussian random vectors.
Lemma 22.
Let $Z_{1},\ldots,Z_{n}$ be i.i.d. subGaussian random vectors in $\mathbb{R}^{d}$ with subGaussian parameter $\sigma^{2}>0$. For $i\in[n]$, we define the truncated version of $Z_{i}$ as
$\bar{Z}_{i}=Z_{i}1{\{\|Z_{i}\|_{2}\leq B_{z}\}}$
with $B_{z}=2\sigma\sqrt{d+(s+1)\log(n)}$. Let $f_{u}:\mathbb{R}^{d}\to\mathbb{R}$ be any function that satisfies
$\mathbb{E}[f_{u}^{2}(Z_{i})]$ $\leq R_{2}$ (145)
$\left|f_{u}(\bar{Z}_{i})\right|$ $\leq R_{1}$ (146)
$\left|f_{u}(\bar{Z}_{i})-f_{u^{\prime}}(\bar{Z}_{i})\right|$ $\leq L_{f}\|u-u^{\prime}\|_{2},\qquad\text{for any $u,u^{\prime}\in\mathbb{S}^{p-1}$.}$ (147)
For any $\epsilon\in(0,1)$, with probability at least $1-4p^{-s}$, we have
$\sup_{u\in\mathbb{S}^{p-1}}\left|{1\over n}\sum_{i=1}^{n}f_{u}(Z_{i})-\mathbb{E}[f_{u}(Z_{i})]\right|\lesssim\sqrt{R_{2}\{\log(n)+p\log(3/\epsilon)\}\over n}+{R_{1}\{\log(n)+p\log(3/\epsilon)\}\over n}+2\epsilon L_{f}.$
Proof.
Define the event
$\mathcal{E}=\bigcap_{i=1}^{n}\mathcal{X}_{j}:=\bigcap_{i=1}^{n}\left\{\|Z_{i}\|_{2}\leq B_{z}\right\}$
with $B_{z}=2\overline{\sigma}\sqrt{d+(1+s)\log(n)}$. Using Lemma 30, we find that
$\mathbb{P}(\mathcal{E})\geq 1-2n^{-s}$ (148)
and we proceed to work on this event $\mathcal{E}$. Since $Z_{i}=\bar{Z}_{i}$ on $\mathcal{E}$, we bound from above
$\sup_{u\in\mathbb{S}^{p-1}}\left\{\left|{1\over n}\sum_{i=1}^{n}f_{u}(\bar{Z}_{i})-\mathbb{E}[f_{u}(\bar{Z}_{i})]\right|+\left|\mathbb{E}[f_{u}(Z_{i})]-\mathbb{E}[f_{u}(\bar{Z}_{i})]\right|\right\}.$
For the second term, note that, for any $u\in\mathbb{S}^{p-1}$,
$\left|\mathbb{E}[f_{u}(Z_{i})]-\mathbb{E}[f_{u}(\bar{Z}_{i})]\right|$ $=\left|\mathbb{E}\left[\left(f_{u}(Z_{i})-f_{u}(\bar{Z}_{i})\right)1{\mathcal{E}^{c}}\right]\right|$
$\leq\sqrt{\mathbb{E}\left[\left(f_{u}(Z_{i})-f_{u}(\bar{Z}_{i})\right)^{2}\right]}\sqrt{1-\mathbb{P}(\mathcal{E})}$ by Cauchy-Schwarz
$\leq\sqrt{\mathbb{E}\left[f_{u}^{2}(Z_{i})\right]+\mathbb{E}\left[f_{u}^{2}(\bar{Z}_{i})\right]}\sqrt{2n^{-s}}$ by (148)
$\leq 2\sqrt{R_{2}n^{-s}}.$ (149)
In the last step, we used
$\mathbb{E}[f_{u}^{2}(\bar{Z}_{i})]\leq\mathbb{E}[f_{u}^{2}(Z_{i})]\leq R_{2}$
from (146). It remains to bound from above
$\sup_{u\in\mathbb{S}^{p-1}}\Delta_{u}:=\sup_{u\in\mathbb{S}^{p-1}}\left|{1\over n}\sum_{i=1}^{n}f_{u}(\bar{Z}_{i})-\mathbb{E}[f_{u}(\bar{Z}_{i})]\right|.$
We use a standard discretization argument. Let $\mathcal{N}_{\epsilon}$ be an $\epsilon$-net of $\mathbb{S}^{p-1}$ such that, for any $u\in\mathbb{S}^{p-1}$, there exists $u^{\prime}\in\mathcal{N}_{\epsilon}$ with $\|u-u^{\prime}\|_{2}\leq\epsilon$ and $|\mathcal{N}_{\epsilon}|\leq(3/\epsilon)^{p-1}$. For any $\delta>0$, let $\bar{u}\in\mathbb{S}^{p-1}$ be such that
$\sup_{u\in\mathbb{S}^{p-1}}\Delta_{u}\leq\Delta_{\bar{u}}-\delta.$
It then follows that
$\sup_{u\in\mathbb{S}^{p-1}}\Delta_{u}$ $=\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u}+\sup_{u\in\mathbb{S}^{p-1}}\Delta_{u}-\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u}$
$\leq\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u}+\Delta_{\bar{u}}-\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u}-\delta$
$\leq\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u}+\Delta_{\bar{u}}-\Delta_{\bar{u}^{\prime}}-\delta$
for some $\bar{u}^{\prime}\in\mathcal{N}_{\epsilon}$ with $\|\bar{u}-\bar{u}^{\prime}\|_{2}\leq\epsilon$. Since
$\Delta_{\bar{u}}-\Delta_{\bar{u}^{\prime}}$ $\leq 2\max_{1\leq i\leq n}\left|f_{\bar{u}}(\bar{Z}_{i})-f_{\bar{u}^{\prime}}(\bar{Z}_{i})\right|\leq 2\epsilon L_{f}$ by (147)
and $\delta$ is arbitrary, we have
$\sup_{u\in\mathbb{S}^{p-1}}\Delta_{u}$ $\leq\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u}+2\epsilon L_{f}.$ (150)
We apply Bernstein’s inequality for bounded random variables and take the union bound over $u\in\mathcal{N}_{\epsilon}$ to find that, for any $t>0$,
$\max_{u\in\mathcal{N}_{\epsilon}}\Delta_{u,v}\lesssim\sqrt{R_{2}t\over n}+{R_{1}t\over n}$
with probability at least
$1-2\left(|\mathcal{N}_{\epsilon}|\right)^{2}\exp(-t)=1-2\exp\left\{-t+2(p-1)\log\left(3\over\epsilon\right)\right\}.$
The result follows after we choose $t=2(p-1)\log(3/\epsilon)+s\log n$ and combine (149) and (150). The proof is complete. ∎
Appendix F Auxiliary lemmas
The following lemmas contains some basic results on moments related with (sub-)Gaussian random variables.
Lemma 23.
Let $Z\sim N(0,\sigma^{2})$. Then for any $t\in\mathbb{R}$,
$\mathbb{E}\left[Ze^{Zt}\right]=\sigma^{2}te^{\sigma^{2}t^{2}/2},\qquad\mathbb{E}\left[Z^{2}e^{Zt}\right]=\sigma^{2}\left(1+\sigma^{2}t^{2}\right)e^{\sigma^{2}t^{2}/2}$
Proof.
The proof follows from the Gaussian density and integration by parts. ∎
Lemma 24.
Let $Z\sim N_{L}(0,\sigma^{2}\mathbf{I}_{L})$. For any vectors $u,{\bm{\theta}}\in\mathbb{R}^{L}$, we have
$\mathbb{E}\left[(Z^{\top}u)e^{Z^{\top}{\bm{\theta}}}\right]=\sigma^{2}(u^{\top}{\bm{\theta}})e^{\sigma^{2}\|{\bm{\theta}}\|_{2}^{2}/2}$
$\mathbb{E}\left[(Z^{\top}u)^{2}e^{Z^{\top}{\bm{\theta}}}\right]=\sigma^{2}\left(\|u\|_{2}^{2}+\sigma^{2}(u^{\top}{\bm{\theta}})^{2}\right)e^{\sigma^{2}\|{\bm{\theta}}\|_{2}^{2}/2}.$
Proof.
To prove the first claim, let $Q$ be an $L\times L$ orthogonal matrix such that
$Qu=\|u\|_{2}~\mathbf{e}_{1}.$ (151)
Write $\bar{\bm{\theta}}=Q{\bm{\theta}}$ with $\bar{\bm{\theta}}=(\bar{\bm{\theta}}_{1},\bar{\bm{\theta}}_{-1}^{\top})^{\top}$, and similarly $Z=(Z_{1},Z_{-1}^{\top})^{\top}$. By the rotational invariance of spherical Gaussian, we have
$\mathbb{E}\left[(Z^{\top}u)e^{Z^{\top}{\bm{\theta}}}\right]$ $=\|u\|_{2}\mathbb{E}\left[(Z^{\top}\mathbf{e}_{1})e^{Z^{\top}\bar{\bm{\theta}}}\right]$
$=\|u\|_{2}\mathbb{E}\left[Z_{1}e^{Z_{1}\bar{{\bm{\theta}}}_{1}}\right]\mathbb{E}\left[e^{Z_{-1}^{\top}\bar{\bm{\theta}}_{-1}}\right]$ by independence between $Z_{1}$ and $Z_{-1}$
$=\|u\|_{2}\sigma^{2}\bar{\bm{\theta}}_{1}e^{\sigma^{2}\bar{\bm{\theta}}_{1}^{2}/2}e^{\sigma^{2}\|\bar{\bm{\theta}}_{-1}\|_{2}^{2}/2}$ by Lemma 23
$=\|u\|_{2}\sigma^{2}\bar{\bm{\theta}}_{1}e^{\sigma^{2}\|{\bm{\theta}}\|_{2}^{2}/2}.$
The claim follows by noting that $\bar{\bm{\theta}}_{1}={\bm{\theta}}^{\top}Q^{\top}\mathbf{e}_{1}={\bm{\theta}}^{\top}u/\|u\|_{2}$ from (151).
Regarding the second claim, by similar arguments, we have
$\mathbb{E}\left[(Z^{\top}u)^{2}e^{Z^{\top}{\bm{\theta}}}\right]$ $=\|u\|_{2}^{2}~\mathbb{E}\left[(Z^{\top}\mathbf{e}_{1})^{2}e^{Z^{\top}\bar{\bm{\theta}}}\right]$
$=\|u\|_{2}^{2}~\mathbb{E}\left[Z_{1}^{2}e^{Z_{1}\bar{\bm{\theta}}_{1}}\right]\mathbb{E}\left[e^{Z_{-1}^{\top}\bar{\bm{\theta}}_{-1}}\right]$
[…]
completing the proof. ∎
Lemma 25.
Let $Z\in\mathbb{R}^{L}$ be a zero-mean, sub-Gaussian random vector with sub-Gaussian constant $\sigma^{2}$. Then for any $u\in\mathbb{S}^{L-1}$ and ${\bm{\theta}}\in\mathbb{R}^{L}$, one has
$\mathbb{E}\left[(Z^{\top}u)^{4}e^{2Z^{\top}{\bm{\theta}}}\right]\lesssim\sigma^{4}e^{4\sigma^{2}\|{\bm{\theta}}\|_{2}^{2}}.$
Proof.
The proof follows by the Cauchy-Schwarz inequality and the sub-Gaussianity of $Z$. ∎
F.1 Lemmas related on moments of Hermite polynomials
[…]
$H_{r}(t)=r!\sum_{j=0}^{\lfloor r/2\rfloor}{(-1/2)^{j}\over j!(r-2j)!}t^{r-2j}.$ (152)
Lemma 26.
For any $r\in\mathbb{N}$,
$e(r/e)^{r}\leq r!\leq er(r/e)^{r},$ (153)
$\sum_{j=0}^{r}{r\choose j}=(r/2)2^{r},$ (154)
$(a+b)^{r}\leq(r/2)2^{r}(|a|^{r}+|b|^{r}),\qquad\forall a,b\in\mathbb{R}.$ (155)
Proof.
[…]
$\mathbb{E}[X]={r\over 2}=\sum_{j=0}^{r}j{r\choose j}2^{-r},$
the claim follows from
$(r/2)2^{r}=\sum_{j=0}^{r}j{r\choose j}=\sum_{j=1}^{r}j{r\choose j}\geq\sum_{j=1}^{r}{r\choose j}.$
Finally, regarding the last one, we have
$(a+b)^{r}=\sum_{j=0}^{r}{r\choose j}a^{j}b^{n-j}\leq\sum_{j=0}^{r}{r\choose j}(|a|^{r}+|b|^{r}).$
The result follows from (154). ∎
The following lemma bounds from above the 444th moment of $H_{r}(Z)$.
Lemma 27.
Let $Z\sim N(\mu,1)$. Then for any $r\in\mathbb{N}$,
$\mathbb{E}[H_{r}^{4}(Z)]\leq(C\mu)^{4r}+(C\sqrt{r})^{4r}.$ (156)
for some absolute constant $C>0$. Consequently, we have
$\mathbb{E}[H_{r}^{2}(Z)]\leq(C\mu)^{2r}+(C\sqrt{r})^{2r}.$ (157)
Furthermore, for any $Z\sim N(0,1)$ and ${\bm{\theta}}\in\mathbb{R}$, we have
$\mathbb{E}[H_{r}^{4}(Z)\exp(Z{\bm{\theta}})]\leq\left[(C{\bm{\theta}})^{4r}+(C\sqrt{r})^{4r}\right]\exp({\bm{\theta}}^{2}/2).$ (158)
Proof.
By (152), we have
$\mathbb{E}[H_{r}^{4}(Z)]$ $\leq\mathbb{E}\left[r!\sum_{j=0}^{\lfloor r/2\rfloor}{1\over 2^{j}j!(r-2j)!}Z^{r-2j}\right]^{4}$
$\leq(r/2+1)^{3}\sum_{j=0}^{\lfloor r/2\rfloor}\left({r!\over 2^{j}j!(r-2j)!}\right)^{4}\mathbb{E}\left[Z^{4(r-2j)}\right]$ by Holder’s inequality
$=(r/2+1)^{3}\sum_{j=0}^{\lfloor r/2\rfloor}\left[{r\choose 2j}{(2j)!\over 2^{j}j!}\right]^{4}\mathbb{E}\left[Z^{4(r-2j)}\right]$
Note that, by using (155) and upper bounds of moments of standard gaussian,
$\mathbb{E}\left[Z^{4(r-2j)}\right]\leq 2(r-2j)2^{4(r-2j)}\left(|\mu|^{4(r-2j)}+\mathbb{E}[(Z-\mu)^{4(r-2j)}]\right)\leq C^{r}(|\mu|^{4r}+(\sqrt{2r})^{4r}).$
By also using
$\sum_{j=0}^{\lfloor r/2\rfloor}{r\choose 2j}\leq\sum_{j=0}^{r}{r\choose j}\leq C^{r}$
from (154) and
${(2j)!\over 2^{j}j!}\leq{2^{j}j^{j}(2j)\over e^{j}}\leq 2jj^{j},$
from (153), we obtain
$\mathbb{E}[H_{r}^{4}(Z)]$ $\lesssim r^{3}C^{5r}\max_{0\leq j\leq\lfloor r/2\rfloor}\left[{(2j)!\over 2^{j}j!}\right]^{4}(|\mu|^{4r}+(\sqrt{2r})^{4r})$
$\lesssim(Cr)^{2r}\left(|\mu|^{4r}+(2r)^{2r}\right)$
completing the proof of (156). The second claim in (157) follows trivially.
Finally, to prove (158), we have
[…]
$=\exp({\bm{\theta}}^{2}/2)\mathbb{E}[H_{r}^{4}(Z_{\bm{\theta}})]$
with $Z_{\bm{\theta}}\sim N({\bm{\theta}},1)$. The proof is completed by invoking (156). ∎
As an application of Lemma 29, we have the following bound on $\mathbb{E}[H_{r}^{4}(Z^{\top}v)\exp(Z^{\top}{\bm{\theta}})]$ for any ${\bm{\theta}}\in\mathbb{R}^{L}$, $v\in\mathbb{S}^{L-1}$ and $Z\sim N_{L}(0,\mathbf{I}_{L})$.
Lemma 28.
Let $Z\sim N_{L}(0,\mathbf{I}_{L})$. For any ${\bm{\theta}}\in\mathbb{R}^{L}$ and $v\in\mathbb{S}^{L-1}$, we have
$\mathbb{E}[H_{r}^{4}(Z^{\top}v)\exp(Z^{\top}{\bm{\theta}})]\leq\left[(C\|{\bm{\theta}}\|_{2})^{4r}+(C\sqrt{r})^{4r}\right]\exp(\|{\bm{\theta}}\|_{2}^{2}/2)$
for some absolute constant $C>0$.
Proof.
We first argue as the proof of Lemma 24 that there exists $Q\in\mathbb{O}_{L\times L}$ such that $Qv=\mathbf{e}_{1}$. Write $\bar{\bm{\theta}}=Q{\bm{\theta}}$ with $\bar{\bm{\theta}}=(\bar{\bm{\theta}}_{1},\bar{\bm{\theta}}_{-1}^{\top})^{\top}$, and similarly $Z=(Z_{1},Z_{-1}^{\top})^{\top}$. Then
$\mathbb{E}[H_{r}^{4}(Z^{\top}v)\exp(Z^{\top}{\bm{\theta}})]$ $=\mathbb{E}[H_{r}^{4}(Z_{1})\exp(Z_{1}\bar{\bm{\theta}}_{1})]\mathbb{E}[\exp(Z_{-1}^{\top}\bar{\bm{\theta}}_{-1})]$
$=\mathbb{E}[H_{r}^{4}(Z_{1})\exp(Z_{1}\bar{\bm{\theta}}_{1})]\exp(\|\bar{\bm{\theta}}_{-1}\|_{2}^{2}/2)$
$\leq\left[(C\bar{\bm{\theta}}_{1})^{4r}+(C\sqrt{r})^{4r}\right]\exp(\bar{\bm{\theta}}_{1}^{2}/2)\exp(\|\bar{\bm{\theta}}_{-1}\|_{2}^{2}/2).$
The last step invokes (158) in Lemma 27. The result follows by noting that $\|\bar{\bm{\theta}}_{1}\|_{2}=\|{\bm{\theta}}\|_{2}$ and $\bar{\bm{\theta}}_{1}\leq\|{\bm{\theta}}\|_{2}$. ∎
The following lemma bounds from above $|H_{r}(x)|$.
Lemma 29.
For any $r\in\mathbb{N}$,
$|H_{r}(x)|\leq(C\sqrt{r})^{r}(|x|^{r}+1)$
for some absolute constant $C>0$.
Proof.
Using the same arguments of proving (156), we have, for any $x\geq 0$,
$|H_{r}(x)|$ $\leq r!\sum_{j=0}^{\lfloor r/2\rfloor}{1\over 2^{j}j!(r-2j)!}x^{r-2j}=\sum_{j=0}^{\lfloor r/2\rfloor}{r\choose 2j}{(2j)!\over 2^{j}j!}x^{(r-2j)}\leq(C\sqrt{r})^{r}\max_{0\leq j\leq\lfloor r/2\rfloor}x^{r-2j}.$
The result follows immediately. ∎
The following lemma states upper bounds of the quadratic form of a sub-Gaussian random vector (Hsu et al., 2012).
Lemma 30.
Let $\xi\in\mathbb{R}^{d}$ be a subGaussian random vector with parameter $\gamma_{\xi}$. Then, for all symmetric positive semi-definite matrices $H$, and all $t\geq 0$,
$\mathbb{P}\left\{\xi^{\top}H\xi>\gamma_{\xi}^{2}\left(\sqrt{{\rm tr}(H)}+\sqrt{2t\|H\|_{\rm op}}\right)^{2}\right\}\leq e^{-t}.$
The following lemma states the well-known matrix-valued Bernstein inequalities. See, for instance, Minsker (2017, Theorem 3.1, Corollary 3.1 and Corollary 4.1).
Lemma 31 (Matrix-valued Bernstein inequality).
Let $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}\in\mathbb{R}^{d\times d}$ be independent, symmetric random matrices with zero mean and $\max_{i\in[n]}\|\mathbf{X}_{i}\|_{\mathrm{op}}\leq U$ almost surely. Denote $\sigma^{2}:=\|\sum_{i=1}^{n}\mathbb{E}[\mathbf{X}_{i}^{2}]\|_{\mathrm{op}}$. Then for all $t\geq{1\over 6}(U+\sqrt{U^{2}+36\sigma^{2}})$,
$\mathbb{P}\left(\left\|\sum_{i=1}^{n}\mathbf{X}_{i}\right\|_{\mathrm{op}}>t\right)\leq 14\exp\left(-{t^{2}/2\over\sigma^{2}+Ut/3}+\log(d)\right).$
The next lemma states an anti-concentration inequality of $v^{\top}{\bm{\theta}}$ for any $v$ uniformly drawn from $\mathbb{S}^{d-1}$. See, for instance, the proof of Lemma 3.1 in Doss et al. (2023).
Lemma 32.
Let ${\bm{\theta}}\in\mathbb{R}^{d}$ be any fixed vector. For any $v$ uniformly drawn from $\mathbb{S}^{d-1}$, one has that, for all $t\geq 0$,
$\mathbb{P}\left\{|v^{\top}{\bm{\theta}}|<t\|{\bm{\theta}}\|_{2}\right\}<t\sqrt{d}.$